### v2

# Learning Source Biases: Multi-sourced Misspecifications and Consequences

Abstract
We study how a decision maker (DM) learns about the biases of unfamiliar information sources. Absent any friction, a rational DM uses familiar sources as yardsticks to discern the true bias of a source. If the DM has misspecified beliefs, this process fails. We derive long-run beliefs, behavior, welfare, and corresponding comparative statics when the DM holds dogmatic, incorrect beliefs about the biases of several familiar sources. The distortion due to misspecified learning aggregates misspecifications associated with familiar sources independently of sources the DM learns about. This has implications for labor market discrimination, media bias, and project finance and oversight.
1 Introduction
Good decision making in today’s information landscape involves the utilization of often biased information from a variety of sources. Some sources, such as traditional data or trusted advisors, are familiar to the decision maker (DM), and so he believes that he understands their biases well. Still others, such as algorithm recommendations and robo-advisors, are novel and unfamiliar, and the DM is much less certain about their biases. In an ideal setting, an astute decision maker might be able to learn about the bias inherent in each information source over time, using familiar sources to gauge unfamiliar ones, and his beliefs would eventually converge to the truth. However, a serious problem arises when the DM holds dogmatic, wrong, beliefs about the biases inherent in familiar sources, due to factors that elude detection and corrections. The long-run implications of this error are the subject of our paper.
Following the paradigm of misspecified Bayesian learning pioneered by Berk (1966), we model a DM who consults several information sources to predict an evolving state of the world. Initially, the DM holds different prior beliefs about the biases of familiar and unfamiliar sources. For familiar sources, the prior is precise but can be inaccurate (hereafter misspecified), in that it concentrates on an event in which the perceived bias of the source differs from the truth. For unfamiliar sources, the prior is uncertain and, crucially, contains truth in its support. Over time, the DM observes signal realizations and updates his belief about source biases in a Bayesian manner. We examine how learning in the presence of multiple, misspecified sources breeds persistent misperceptions about unfamiliar sources and fuels distortions of the DM’s long-run behavior.
Our framework can potentially be applied to a broad set of important economic phenomena. To illustrate, consider an upper-level manager who inquires about the productivity of minority employees from a mix of familiar sources — his personal source, senior raters, standard HR algorithms — and unfamiliar sources — novel employee data and newly installed committee members. Although these sources are prone to gender and racial biases, in a frictionless world, biases may be inconsequential, at least in the long run. For example, if a novel source gives consistently lower employee ratings than a known, impartial source, then it is regarded as adversarial to minorities, and its bias is accounted for when information from diverse sources is aggregated to predict minority employees’ productivity.
Reality, however, is not frictionless, particularly when the manager’s beliefs about familiar sources are misspecified. As someone who exhibits an implicit bias against minorities, the manager may give unduly low ratings to minority employees automatically and unconsciously. [^4] Implicit bias is an unconscious negative attitude towards a particular social group, making it challenging to eliminate due to its automatic nature. The renowned Implicit Association Test (IAT) developed by Greenwald et al. (1998) quickly assesses individuals’ associations between social groups and stereotypes, often in milliseconds. The opinions he gleans from other evaluators can be tainted by prejudice, stereotypes, and intergroup preferences, while his success in detecting these biases is limited, particularly in large organizations or when transparency and accountability are lacking (Roberson et al., 2007). Finally, algorithmic biases may arise due to poor designs or incomplete training, but they remained unnoticed until they garnered recent attention from academia and the popular press (Dastin, 2022). Since all these errors influence the manager’s learning about unfamiliar sources and, ultimately, the perceived productivity of minority employees and their compensation, their theoretical and practical consequences warrant a comprehensive investigation.
Multi-sourced misspecified learning could be challenging to study for several reasons. In our leading example, suppose that the manager perceives two sources as impartial, although in reality one is biased in favor of minorities and the other is biased against them. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. One potential approach is to find a middle ground and use this as a neutral point to form beliefs about other sources. Still, the question remains: How exactly can this be achieved and how does the answer depend on the correlations between the signals? Correlation affects how different signals are weighed in the learning process. With multi-sourced misspecifications, the need to weigh different sources promotes the role of correlation in shaping learning outcomes, and the compound effect can escape even the most sophisticated analysis.
We develop an approach that works for arbitrary numbers of misspecified sources. Our analysis is conducted in a Gaussian environment in which the signal generated by a source in a given period equals the bias of the source, plus the true state of the world in that period and a Gaussian error. Errors are independent of the true state and of each other in the baseline model, but can take more general forms in later sections. Our main result, presented in Section 3.2, captures the distortions of the DM’s long-run perceptions and behavior by a single, intuitive, metric. In the baseline model, this metric equals the total misspecification associated with familiar sources, weighted by their relative precision to each other. As we exacerbate the misspecification associated with a source or increase its precision relative to others, its impact on the metric increases. In the leading example, if a majority of familiar sources, especially those exhibiting high signal-to-noise ratios, are biased against minorities but are nevertheless perceived as impartial, then the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities and is mirrored in his beliefs about unfamiliar sources through misspecified learning. In the long run, the manager underestimates the bias of any adversarial source toward minorities while overestimating that of any source favorable to them. He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by our metric.
Our approach utilizes existing mathematical results on low-rank updates of invertible matrices. In the baseline model, signals (net of their biases) equal the true state plus idiosyncratic errors, so their covariance matrix is a rank-one update of a diagonal matrix. When signals are correlated through a single common channel, the impact of correlation on the outcomes of misspecified learning is limited. Consequently, the metric that captures the distortions of the DM’s long-run belief and behavior exhibits a striking feature that we call DOOM, as it depends only on sources with misspecified biases but not those that the DM learns about. DOOM continues to hold in variants of the baseline model in which the signals depend on a common latent factor, or when the true state and the signals can take high-dimensional values (see Sections 4.1 and 4.2). When correlations between signals become arbitrary, the DM’s misperceptions about source biases satisfy a weaker form of DOOM, as they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else. Further, the distortion in the DM’s long-run behavior always satisfies DOOM, due to the simple fact that the DM’s limiting belief concentrates on the minimizer of the Kullback-Leibler divergence from truth.
The comparative statics of our metric have important implications for reality. In the context of labor market discrimination, DOOM implies that the damage caused by a few rotten apples cannot be rectified by installing objectively neutral but lesser-known members of the minority group to the evaluation panel, or by training junior raters to reduce errors in employee assessments. The same cannot be said about sources that the DM holds dogmatic, wrong, beliefs about, as our metric is in general nonmonotonic in their signal-to-noise ratios. In the presence of multiple misspecified sources, improving the precision of one source renders the others relatively imprecise. The overall impact on our metric is in general ambiguous, and this has implications for the use of rater training to curtail discrimination. Details are in Section 3.5.
While it is tempting to conjecture that correcting the misspecifications associated with familiar sources may improve learning outcomes, reality is more subtle when the covariance matrix of the signals is also subject to misspecified learning. In Section 4.3, we show that learning the covariance matrix of the signals, in addition to their biases, makes the signals appear noisier to the DM than they truly are. Consequently, the DM relies less on signal realizations when predicting the true state of the world, and this countervailing force is shown to partially correct the distortions due to misspecified learning. Policies that operate by correcting misspecifications block this novel channel and so could entail an ambiguous consequence on learning outcomes. This finding speaks to the debiasing programs used by real-world organizations to address discrimination, such as implicit bias training, priming shared identities, and promoting algorithmic fairness. It suggests that even if these programs are effective, their behavioral and welfare consequences may remain ambiguous.
In addition to employee evaluation and compensation, the aforementioned phenomenon occurs when voters form perceptions of the biases inherent in new social media sites based on their erroneous beliefs about traditional media. There is evidence that partisans tend to view party-congenial sources as neutral, while displaying excessive hostility toward out-party sources (Pennycook and Rand, 2021; Hassell et al., 2020). Likewise, a manager seeking project advice may allow his hardwired misperceptions about familiar sources — such as overoptimism, confirmation bias, and overtrust in one’s inner circle (Malmendier, 2018) — to shape his views about unfamiliar sources such as external consultants and robo-advisors. In Section 5, we demonstrate how these ingrained biases can spill into new domains, generating the hostile media effect and false polarization, [^5] The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is. False polarization measures the difference between the perceived distance between two groups and the actual distance between them. Both phenomena have attracted considerable scholarly attention within the domains of political science, communication, and psychology. Section 5.1 offers a synthesis of the literature. in addition to polarizing voters’ behavior and distorting managers’ portfolio choices. We further contend, based on our comparative statics, that corrective measures that target political misinformation or managerial biases may give rise to unintended consequences. For example, while fact-checking new media always improves voters’ welfare because of DOOM, fact-checking tradition media generally has an ambiguous impact due to the distortions that arise from misspecified learning. The latter are nonmonotonic in the accuracy of familiar sources and the treatment effects of misperception recalibration.
1.1 Related literature
The study of misspecified Bayesian learning models has a long tradition in statistical decision theory. The canonical paradigm endows the DM with a prior belief over the models that describe the underlying environment. If the prior does not contain the true model in its support, it is misspecified. The seminal work of Berk (1966) provides conditions under which the DM’s limiting posterior belief concentrates on the set of models that minimize the Kullback-Leibler (KL) divergence from truth.
Economic models of misspecified Bayesian learning can be broadly categorized into passive learning models and active learning models. The former, including ours, assume that the actions taken by the DM have no impact on the signals generated by truth. The work most related to ours is carried out by Heidhues et al. (2019) (HKS). These authors consider environments where the signals generated by unknown fundamentals follow multivariate Guassian distributions. Such environments are shown to satisfy Berk’s conditions for the convergence of the limiting belief to KL minimizers. Our proof of the convergence of the limiting belief builds on this result.
The main differences between our results and those of HKS lie in the characterization of the limiting belief. In particular, HKS focus on situations where a single dimension of the fundamentals is misspecified, which is the natural setting to understand the implications of overconfidence. We entertain the possibility of multidimensional misspecifications, which is essential to understand how a DM aggregates biased sources in the settings discussed previously. Our characterization extends beyond that of HKS and works for arbitrary numbers of misspecified sources with general covariance matrices. The structural insights it uncovers, such as DOOM, have escaped the attention of previous work. Our comparative statics are sharpest when the covariance matrix of the signals constitutes low-rank updates of diagonal matrices. HKS consider alternative forms of covariance matrices that are suitable for their exploration of the consequences of overconfidence on intergroup prejudice.
In contrast, models of misspecified active learning allow the DM’s actions to affect the signals that he subsequently uses to form beliefs about the truth. The literature on misspecified active learning has recently advanced significantly. Among others, Esponda and Pouzo (2016) propose Berk-Nash equilibrium as a steady-state notion for misspecified learning dynamics. Esponda et al. (2021), Fudenberg et al. (2021), Bohren and Hauser (2021), and Frick et al. (2023) examine the convergence of Berk-Nash equilibria in single and multi-agent environments. Ba (2021) entertains the possibility of regime shifts. He and Libgober (2020) examine the evolutionary foundations of perpetual model misspecification. None of these papers shares our focus on the characterization of steady states with multi-sourced misspecifications.
There is a broader conceptual connection between our work and the Bayesian network approach of Spiegler (2020) to model misspecified causality, in that both allow for misspecifications of high-dimensional objects. Yet Spiegler’s use of directed acyclic graphs to model causality has no obvious counterpart in our setting, and his solution concept requires that the DM take endogenous actions to fully explain the observed data. Our steady state does not involve endogenous actions and has a misspecified-Bayesian-learning foundation.
Developments on the theoretical frontier have spurred novel research in various areas, including labor market discrimination, media bias, and finance. Within the realm of labor market discrimination, Bohren et al. (2019) and Bohren et al. (2023) propose methods to distinguish discrimination due to erroneous beliefs from those rooted in tastes or rational beliefs. The current paper adds to this agenda by exploring a novel channel: misspecified learning of unfamiliar information sources. The main takeaways from our analysis — namely dogmatic, wrong, beliefs about familiar sources can breed persistent misperceptions about unfamiliar sources and generate nuanced comparative statics when subject to informational interventions — have no counterparts in existing theories of discrimination. To the extent that interventions such as improved source accuracy or misperception recalibration are also proposed by Bohren et al. (2023) to identify and address discrimination, our analysis could impact how policies are designed to combat discrimination in the real world.
The economic literature on media bias is vast, but is devoted to understanding actual media biases (Anderson et al., eds, 2016).666(Mis)perceived media bias is an important topic in political science, communication, and related fields (Feldman, 2014). These strands of literature are discussed in Section 5.1. A notable exception is made by Gentzkow et al. (2018) (GWZ), whose model of misperceived media bias features a payoff-irrelevant confounder that is independent of the payoff-relevant state. A source is biased if it exhibits a nonzero correlation with the confounder. The DM misperceives his personal source as unbiased and uses this belief to guide his consumption of other media sources.
The model of GWZ differs from ours in its construct and predictions, in addition to requiring different methodologies to analyze. First, GWZ examine an active learning model with a single misspecified source. Second, their way of modeling biases (as correlations with a random confounder) implies that the distortion in the DM’s long-run behavior has a zero unconditional mean, which is not the case here. A central focus of GWZ is on the endogenous trust in news sources, defined as the latter’s (mis)perceived correlations with the true state. Trust is correctly specified throughout this paper and is known to the DM for the most part. When endogenized as a subject of misspecified learning in Section 4.3, we find no effect on perceived news biases and only a moderation of the DM’s behavior.
The literature on behavioral finance is vast (see Malmendier 2018 for a survey), although research that explores misspecified Bayesian learning as a source of friction remains sparse. Molavi et al. (2023) examine a passive learning model in which asset returns follow a high-dimensional hidden factor model, although the models used by market participants to explain observed data have a lower dimensionality. The authors examine the consequences of this departure from the rational expectations paradigm for asset pricing and the dynamics of forecast errors. The Sherman-Morrison formula — a tool to invert rank-one perturbations of invertible matrices — is present in their mathematical derivations (and those of Levy et al. 2022) but plays a less significant role than in our analysis.
2 Baseline model
Time is discrete and infinite. In each period $t=1,2,\cdots$, a random state $\omega_{t}$ is drawn independently from the standard normal distribution. A decision maker (DM) — informed by a finite set $\mathcal{S}\coloneqq\{1,\cdots,N\}$ of sources — predicts the mean of $\omega_{t}$. The signal generated by source $i\in\mathcal{S}$ in period $t$ is $X_{i,t}=\omega_{t}+b_{i}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}$ is the bias of the source, and $\varepsilon_{i,t}$ is a Gaussian error with mean zero and variance $v_{i}>0$. In the baseline model, errors are assumed to be independent of the true state, across each other, and over time.
Initially, the DM holds a prior belief about source biases. In every subsequent period, he updates his belief in a Bayesian manner, based on the prior and signal realizations. It is well known that if true biases, which we denote by $b_{i}^{*}$s, lie in the support of the prior, then the DM’s posterior belief will eventually converge to the truth. His long-run behavior is the best linear predictor of the state given signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$, under the (correct) belief that source biases have value $b^{*}\coloneqq[b_{1}^{*},\text{ }\cdots\text{ },b_{N}^{*}]^{\top}$.
We examine the case where the DM’s prior is partly misspecified. In the baseline model, we focus on a particular kind of misspecification, whereby the DM holds dogmatic, wrong, beliefs about the biases of some sources and uses them to guide his learning about other sources. We examine how such misspecifications distort the DM’s learning and, ultimately, his long-run belief and behavior.
Formally, we say that the DM misspecifies the bias of source $i$ if he initially assigns probability one to an event in which $b_{i}=\tilde{b}_{i}$, for some $\tilde{b}_{i}\neq b_{i}^{*}$. Let $\mathcal{M}\coloneqq\{1,\cdots,M\}$ denote the set of sources with misspecified biases. Suppose that $1\leq M<N$, so that the DM misspecifies the biases of some but not all sources. For those sources in $\mathcal{S}-\mathcal{M}\coloneqq\{M+1,\cdots,N\}$, we assume that the DM’s prior is nondegenerate on $\mathbb{R}^{N-M}$. The support of the prior is $\Theta\coloneqq\{b\in\mathbb{R}^{N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. The outcomes of Bayesian learning based on this prior are the subject of the current study.
Three remarks are in order. For now, we assume that signals equal biases plus the true state and independent errors, and that their covariance matrix is known to the DM from the outset. In Sections 3.4 and 4.3, we entertain more general forms of signals whose covariance matrix may also be a subject of misspecified learning. Second, we treat misspecification and degenerate, wrong, prior as synonyms for most of the paper; alternative forms of misspecifications are discussed in Section 4.1. Finally, it will become clear that adding sources with correctly specified biases to the analysis would not affect our predictions in any meaningful way.
The example presented below serves to illustrate our framework and results throughout the paper.
Example 1.
[…]
3 Analysis
To proceed with our analysis, we first present some preliminary concepts and results. These are followed by a formal statement of the main results, then a discussion of their intuitions, proofs, and policy implications.
3.1 Preliminaries
We first formalize the problem faced by the DM. Let $\mathcal{N}(b,\Sigma)$ denote an arbitrary $N$-dimensional normal distribution with mean $b\in\mathbb{R}^{N}$ and covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$, and let $\mathcal{N}(b^{*},\Sigma^{*})$ denote the true distribution of the signals. The Kullback-Leibler divergence from $\mathcal{N}(b,\Sigma)$ to $\mathcal{N}(b^{*},\Sigma^{*})$ is
$D_{KL}(b^{*},\Sigma^{*}\|b,\Sigma)=\frac{1}{2}\left(\text{tr}(\Sigma^{-1}\Sigma^{*})-N+(b-b^{*})^{\top}\Sigma^{-1}(b-b^{*})+\log\frac{\text{det}\Sigma}{\text{det}\Sigma^{*}}\right).$ (1)
The next lemma — which adapts the seminal result of Berk (1966) and that of HKS to the current setting — shows that the DM’s long-run belief concentrates on the set of distributions that minimize the Kullback-Leibler divergence from the true distribution (hereafter KL minimizers). That is, for every open cover of the set of KL minimizers, the DM assigns probability one to it in the limit, as $t\rightarrow+\infty$.
Lemma 1.
[…]
Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ be a typical KL minimizer. Since $\Sigma^{*}$ is known to the DM from the outset, $\widehat{\Sigma}=\Sigma^{*}$. Simplifying the KL divergence accordingly turns the DM’s problem into:
$\min_{b\in\Theta}\frac{1}{2}(b-b^{*})^{\top}\Sigma^{*-1}(b-b^{*}).$ (2)
In Section 3.3, we show that the solution to (2) is unique and fully captured by $\Delta\coloneqq\hat{b}-b^{*}$. Call $\Delta$ the DM’s misperception about source biases in the long run, and note that $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}$ $\forall i\in\mathcal{M}$. Let $\text{BLP}^{*}(X)$ and $\widehat{\text{BLP}}(X)$ denote the best linear predictors of the state given signals, under the belief that source biases have values $b^{*}$ and $\hat{b}$, respectively. These are the long-run behaviors of the DM under the correctly specified model and the misspecified model, respectively, and their difference captures the distortionary effect of misspecified learning on the DM’s long-run behavior.
3.2 Results
Our main results are twofold. The first characterizes the distortionary effects of misspecified learning on the DM’s long-run belief and behavior. The second investigates comparative statics of distortionary effects.
For starters, write $\nu_{i}$ for the signal-to-noise ratio $1/v_{i}$ of source $i$. For each $i\in\mathcal{M}$, define
$\gamma_{i}\coloneqq\frac{\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$
as the relative precision of the signal generated by source $i$ compared to other misspecified sources. Define
$\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\gamma_{i}\Delta_{i}$
as the total misspecification across the sources in $\mathcal{M}$, weighted by their relative precision to each other. Then
Theorem 1.
  1. (i)
    $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$; $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
  2. (ii)
    $\overline{\Delta}$ depends only on $(\Delta_{i},\nu_{i})_{i\in\mathcal{M}}$. For each $i\in\mathcal{M}$, $\partial\overline{\Delta}/\partial\Delta_{i}>0$, while $\partial\overline{\Delta}/\partial\nu_{i}$ generally has an ambiguous sign.
Part (i) of Theorem 1 reduces the distortionary effects of misspecified learning to a single metric: $\overline{\Delta}$. To develop intuitions, suppose that the manager in our leading example begins with an erroneous belief that two sources are both impartial, although in reality one is biased against minorities and the other is in favor of them, that is, $b_{1}^{*}<0<b_{2}^{*}$ and $\tilde{b}_{1}=\tilde{b}_{2}=0$. Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences. One potential approach is to find a middle ground, that is, a position in $(b_{1}^{*},b_{2}^{*})$, and use it as a neutral point to form beliefs about other sources. The more precise source $i$ is relative to source $j$, the closer the neutral point is to $b_{i}^{*}$. In the case where $\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}<0$, the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities of magnitude $\overline{\Delta}=|\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}|$, and it is mirrored in his beliefs about unfamiliar sources through misspecified learning. In the long run, the DM underestimates the bias of any source that is adversarial toward minorities while overestimating that of any source that is favorable to them, assuming that $\overline{\Delta}$ is not excessively large. He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by $\overline{\Delta}$.
Part (ii) of Theorem 1 investigates the comparative statics of our metric. The result $\partial\overline{\Delta}/\partial\Delta_{i}>0$ is the most intuitive. In the leading example, this means that correcting the unconscious bias associated with any familiar source reduces its impact on the metric.
The comparative statics regarding source precision are more nuanced. As we increase the absolute precision of source $i$ (as captured by $\nu_{i}$), we also increase its relative precision (as measured by $\gamma_{i}$), while decreasing the relative precision $\gamma_{j}$ of any other source $j\in\mathcal{M}-\{i\}$. As a consequence, the DM takes source $i$ more seriously in the learning process, and misspecifications of its bias distort the DM’s long-run belief and behavior more heavily than before. The opposite happens to source $j$. Depending on the exact misspecifications associated with each individual source, the outcome is either an exacerbation of the distortionary effect, or a mitigation thereof. Such an ambiguity hinges on the multi-sourced nature of misspecifications, i.e., $M\geq 2$. If instead $M=1$ (as in HKS), then raising $\nu_{i}$ unambiguously increases the magnitude of $\overline{\Delta}$, due to the first effect described above.
Perhaps the most striking feature of $\overline{\Delta}$ is that it depends only on misspecified sources (DOOM) but not those with initially unknown biases. In general, DOOM fails because the correlations between the signals affect how the latter are weighed in the learning process. The presence of multi-sourced misspecifications can significantly complicate the analysis, as the need to weigh different sources promotes the role of correlation in shaping learning outcomes. The compound effect may escape even the most sophisticated analysis. When correlations are concentrated on only a few channels (as in the baseline model), their impacts on learning outcomes are limited. The proof presented in the next section formalizes these ideas.
3.3 Proof of Theorem 1
Throughout this paper, we use $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ to denote the square matrix created from a vector, and $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$. Write the covariance matrix $\Sigma^{*}$ of the signals as
$\left[\begin{array}[]{c|c}A_{M\times M}&B_{M\times(N-M)}\\ \hline\cr C_{(N-M)\times M}&D_{(N-M)\times(N-M)}\end{array}\right],$
where
$A\coloneqq\begin{bmatrix}\Sigma^{*}_{1,1}&\cdots&\Sigma^{*}_{1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{M,1}&\cdots&\Sigma^{*}_{M,M}\end{bmatrix}\text{ and }C\coloneqq\begin{bmatrix}\Sigma^{*}_{M+1,1}&\cdots&\Sigma^{*}_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{N,1}&\cdots&\Sigma^{*}_{N,M}\end{bmatrix}$
capture the correlations between familiar sources and those between familiar and unfamiliar sources, respectively.
By Lemma 1, the DM’s long-run belief is a solution to (2). Rewrite (2) as
$\min_{\Delta\in\mathbb{R}^{N}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i=1,\cdots,M}\frac{1}{2}\Delta^{\top}\Sigma^{*-1}\Delta,$
and note that its solution is fully determined by the system of first-order conditions: $\forall j=M+1,\cdots,N$,
$\sum_{i=1}^{N}\Sigma^{*-1}_{ji}\Delta_{i}=0.$ (3)
We proceed in two steps.
Step 1.
Consider first the case where $\Sigma^{*}$ is an arbitrary symmetric positive definite matrix. The next lemma solves for (3) in this general case. For ease of notation, write $\widetilde{\Delta}$ for the vector $[\Delta_{1},\cdots,\Delta_{M}]^{\top}$ of misspecifications associated with familiar sources, and $\widehat{\Delta}$ for the vector $[\Delta_{M+1},\cdots,\Delta_{N}]^{\top}$ of misperceptions about unfamiliar sources.
Lemma 2.
For any symmetric positive definite matrix $\Sigma^{*}\in\mathbb{R}^{N\times N}$, the solution to (3) is $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ and in particular satisfies $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$.
The proof of Lemma 2 exploits only the first-order conditions associated with KL minimization, together with the block matrix inversion formula. The result prescribes a two-step procedure for aggregating the misspecifications associated with familiar sources into distortions of the DM’s long-run belief and behavior. The first step weighs misspecifications using the correlations between familiar sources (as captured by $A^{-1}\widetilde{\Delta}$). The second step transforms these weighted sums into the DM’s misperceptions about unfamiliar sources, using the correlations between familiar and unfamiliar sources as captured by $C$. Interestingly, correlations between unfamiliar sources do not enter this weighting process. Furthermore, the distortion of the DM’s long-run behavior depends only on misspecified sources but not those he learns about, thus always satisfying DOOM.
Step 2.
While the DM’s misperceptions about source biases do not generally satisfy DOOM due to the complications that arise in the weighting process (as captured by $CA^{-1}$), they become relatively straightforward when the correlations between the signals are concentrated on only a few channels. In the environment considered in the baseline model, distortions of the DM’s long-run perceptions and behavior collapse into a single metric: $\overline{\Delta}$.
A seemingly innocuous assumption turns out to be key, namely signals net of their biases equal the true state plus idiocyncratic errors. As a result, their covariance matrix $\Sigma^{*}$ is the sum of a diagonal matrix: $\text{diag}(v)$, and a matrix of rank one: $\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. The next lemma, proven by Sherman and Morrison (1950), prescribes a formula to invert rank-one updates of invertible matrices.
Lemma 3 (Sherman and Morrison 1950).
Let $G\in\mathbb{R}^{N\times N}$ be an invertible matrix and $x,y\in\mathbb{R}^{N}$ be column vectors. Then $G+xy^{\top}$ is invertible if and only if $1+y^{\top}G^{-1}x\neq 0$, in which case
$(G+xy^{\top})^{-1}=G^{-1}-\frac{G^{-1}xy^{\top}G^{-1}}{1+y^{\top}G^{-1}x}.$
Applying the Sherman-Morrison formula to $A=\text{diag}(v_{1},\cdots,v_{M})+\mathbf{1}_{M}\mathbf{1}_{M}^{\top}$ yields
$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1}\\ \vdots\\ \nu_{M}\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}$
and, hence,
$\mathbf{1}_{M}^{\top}A^{-1}=\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
Substituting these results, together with $C=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}$, into Lemma 2 yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\overline{\Delta}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=-\overline{\Delta}$ as desired.
3.4 Discussions
When distortions of the DM’s long-run perceptions violate DOOM, they still satisfy a weaker form of DOOM (hereafter WDOOM), in that they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else. In the simple environment presented below, the comparative statics of the distortions are analogous to those described in Theorem 1.
Proposition 1.
In the case where $X_{i,t}=b_{i}+\alpha_{i}\omega_{t}+\varepsilon_{i,t}$, $\Delta_{i}=\alpha_{i}\overline{\Delta}$ for all $i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i\in\mathcal{M}}\alpha_{i}^{2}\nu_{i}).$
Proposition 1 shows that when signals are correlated with the true state to varying degrees, the long-run misperceptions held by the DM will vary between sources, albeit only through $\alpha_{i}$. In addition to comparative statics with respect to $\Delta_{i}$ and $\nu_{i}$, a new insight emerges from the analysis: In the case where $\alpha_{i}>0$ $\forall i\in\mathcal{S}$ (if not, let $-X_{i}$ be the subject of analysis), increasing $\alpha_{i}$, $i\in\mathcal{S}-\mathcal{M}$ makes source $i$ more informative and therefore more susceptible to the distortionary effect of misspecified learning. In the long run, this worsens the DM’s misperception about source $i$ relative to other sources that he learns about.
The aforementioned environments satisfy a common feature, namely signals depend on each other only through the true state of the world. In the presence of additional sources of correlation, the weighting formula prescribed by Lemma 2 becomes complex very quickly, as shown by the example presented below.
Example 2.
Suppose that the period-$t$ signal generated by source $i$ is $X_{i,t}=b_{i}+\omega_{t}+\beta_{i}\theta_{t}+\varepsilon_{i,t}$, where $\theta_{t}\sim\mathcal{N}(0,1)$ represents shocks to people’s attitudes toward minorities triggered by social movements such as MeToo, and is assumed to be independent of the true state and the Gaussian errors. The presence of $\theta_{t}$ creates an additional source of correlation between the signals. The resulting covariance matrix $\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}+\beta\beta^{\top}$ is a rank-two update of a diagonal matrix, so its inverse can be obtained by iteratively applying the Sherman-Morrison formula first to $\widetilde{\Sigma}\coloneqq\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ and then $\Sigma^{*}=\widetilde{\Sigma}+\beta\beta^{\top}$.
The algebra is brutal in general. In the case where $N=5$ and $M=3$, it can be shown that
$\begin{bmatrix}\Delta_{4}\\ \Delta_{5}\end{bmatrix}=\begin{bmatrix}\frac{P\beta_{4}+Q}{R}\\ \frac{P\beta_{5}+Q}{R}\end{bmatrix}$
where
$P=\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
$Q=\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
and
$R=\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right).$
$\Delta_{4}$ and $\Delta_{5}$ depend on the correlations between familiar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{M}$) and those between familiar and unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i\in\mathcal{M}$ and $j\in\mathcal{S}-\mathcal{M}$). They are independent of the correlations between unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{S}-\mathcal{M}$), thus satisfying WDOOM.
3.5 Implications
We have discussed the intuitions and mathematical underpinnings of our results in previous sections. We now examine their implications for the various policy interventions used by real-world organizations to address discrimination.
Minority representation at upper echelons.
There is a large economic literature that studies the antidiscrimination effect of increasing minority representations at upper echelons. According to Bertrand and Duflo (2017), a simple mechanical mechanism may nevertheless prove to be game changing. Historically, individuals from underrepresented groups have had limited influence in decision making. To the extent that majority decision makers are biased against minority subordinates, discrimination tends to persist. The presence of minority stakeholders at the various stages of recruitment, evaluation, and promotion can serve as a counteracting force and alter the atmosphere at the top level.
Our result offers a gloomier prospect, suggesting that the damage caused by existing members of the evaluation panel may not be easily rectified by infusing new perspectives. Indeed, DOOM stipulates that the introduction of fresh, lesser-known committee members from the minority group does not alter the metric that captures the distortionary effect of misspecified learning. Diversity efforts at the top may work through alternative channels, but not through expanding the set $\mathcal{S}-\mathcal{M}$.
Reduce decision errors.
Contemporaneous rater training programs offered by numerous real-world organizations emphasize an important lesson: When making decisions that can adversely affect others (e.g., performance reviews), stakeholders should “slow down, meditate, deliberate, and follow elaborate structured procedures” so as to base decisions on facts rather than instincts (Eberhardt, 2020). Recently, Kleinberg et al. (2018) demonstrate the potential of using properly trained AI algorithms to overcome human decision errors in the context of bail decision making.
To the extent that the aforementioned interventions may increase the signal-to-noise ratios of our sources, their impacts on the outcomes of misspecified learning can be quite nuanced. For one thing, changing the signal-to-noise ratios of unfamiliar sources is ineffectual. For another, scaling the signal-to-noise ratios of all familiar sources by the same factor $\alpha>1$ increases the DM’s reliance on them when learning about the biases of unfamiliar sources, thereby amplifying the distortions in his long-run perception and behavior:
$\left|\frac{\sum_{i\in\mathcal{M}}\alpha\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\alpha\nu_{i}}\right|=\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{\alpha^{-1}+\sum_{i\in\mathcal{M}}\nu_{i}}\right|>\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}\right|=|\overline{\Delta}|.$
When changes occur disproportionately at different sources, the outcomes of learning may be improved due to the ambiguity of the sign of $\partial\overline{\Delta}/\partial\nu_{i}$. Yet for this approach to succeed, precise information about the misspecifications associated with each source is needed, and this knowledge is unlikely to be possessed by the DM.
Debiasing programs.
Certainly, the aforementioned interventions may operate through an alternative channel, that of correcting the misspecifications associated with familiar sources (bring $\Delta_{i}$s closer to zero). Indeed, the idea of using one’s own mental resources to debias decisions is backed by years of sociopsychological research into implicit biases, and recent evidence suggests combining it with the use of structured evaluations to prohibit decision making based on automatic, unconscious instincts (Bohnet, 2016; Greenwald and Lai, 2020).888 Outside the corporate realm, the Oakland Policy Department recently adjusted its foot pursuit policy so that officers were instructed to “step back, slow down, call for backup, and think it through” rather than to follow suspects into backyards or blind alleys. Nextdoor Neighbor, a social network for residential neighbors to communicate through, recently started asking its users to provide detailed descriptions of the suspicious activities they wish to report, because “adding frictions allows users to act based on information rather than instinct” (Eberhardt, 2020). Meanwhile, Devine et al. (2012) provide the first evidence that prejudice, stereotypes, and intergroup hostility are “bad habits” that can be broken through longitudinal interventions such as counterstereotypical thinking, priming shared identities, perspective taking, individulating, and intergroup contact. Finally, the growing literature on algorithmic fairness proposes various remedies to remove biases from algorithm design and training (Mitchell et al., 2021). In addition to interventions that can help bring $b_{i}^{*}$s closer to zero, Kalev et al. (2006) and Raji et al. (2020) stress the importance of accountability and transparency. The latter raise the DM’s awareness and consciousness and can potentially bring perceived biases ($\tilde{b}_{i}$s) closer to true biases ($b_{i}^{*}$s).
The efficacy of the aforementioned programs in addressing discrimination is a subject of active research and hot debate (more on this in Section 4.3). In case they are effective in bringing $\Delta_{i}$s closer to zero, their impacts on the DM’s belief and behavior are straightforward in the context of the baseline model. Outside the baseline model, things can become muddier, a subject to which we turn in Section 4.3.
4 Extensions
This section presents several extensions of the baseline model. We first allow the signals to depend on a common latent factor and the true state and the signals to take high-dimensional values. We then examine a problem in which both the biases of the signals and their covariance matrix are the subject of misspecified learning.
4.1 Latent factor
This section examines a variant of the baseline model in which the means of the signals can depend on a common latent factor $b_{N+1}\in\mathbb{R}$, such as employees’ productivity-enhancing investments. Redefine $b\coloneqq[b_{1},\cdots,b_{N+1}]^{\top}$ and call it fundamental. The mean vector of the signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$ is $\Gamma b$, where $\Gamma\coloneqq[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$ is the transformation matrix that operates by adding the latent factor $b_{N+1}$ to the mean of each signal.
Initially, the true value of the latent factor $b_{N+1}$, together with the biases of the sources in $\mathcal{S}-\mathcal{M}$, is unknown to the DM. Everything else is as in the baseline model. The support of DM’s prior becomes $\Theta=\{b\in\mathbb{R}^{N+1}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.
[…]
$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
Proposition 2.
Let everything be as above. Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\Delta_{N+1}=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}/\sum_{i\in\mathcal{M}}\nu_{i}.$
Proposition 2 shows that the presence of a common latent factor has little qualitative impact on the metric that captures the distortionary effects of misspecifed learning. The latter still equals the total misspecification associated with the sources in $\mathcal{M}$, weighted by their relative precision. Its comparative statics are the same as in the baseline model. In the leading example where the distortion of the manager’s worldview constitutes an unconscious negative bias against minorities, $\overline{\Delta}>0$. Proposition 2 predicts that in the long run, the DM undervalues minority workers’ productivity-enhancing investment by $\overline{\Delta}$, in addition to viewing sources favorable to them as too lenient and those adversarial toward them as impartial. Such a misperception clearly has impact on minority workers’ incentive to undertake costly investments, a subject we plan to explore in depth in a follow-up paper.
The proof of Proposition 2 shows that the transformed matrix $\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma$ is a rank-one update of a diagonal matrix, and so the KL minimization problem can be solved using an analogous approach to that developed in Section 3.3. Although it is challenging to exhaust all transformations of the DM’s problem that preserve the structural insights of the current analysis, it is certainly possible to consider specific transformations, such as one where the support of the DM’s prior (equivalently, the choice set in the KL minimization problem) is of the form ×$\times_{i\in\mathcal{S}}[\underline{b}_{i},\overline{b}_{i}]$. The baseline model is a special case where $\underline{b}_{i}=\overline{b}_{i}=\tilde{b}_{i}$ for $i\in\mathcal{M}$, whereas $\underline{b}_{i}=-\infty$ and $\overline{b}_{i}=+\infty$ for $i\in\mathcal{S}-\mathcal{M}$. It is easy to show, using the Lagrangian method, that this model variation has no qualitative impact on the distortionary effect of misspecified learning. All it does is endogenizing $\mathcal{M}$ as the set of sources whose interval constraints are binding; given $\mathcal{M}$, everything said so far remains valid.
4.2 Multidimensional state and signals
This section generalizes the baseline model to encompass multidimensional states and signals. In each period $t=1,2,\cdots,$ Nature draws the true state $\omega_{t}$ from a $K$-dimensional Gaussian distribution with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{0}$. Each dimension represents a distinct skill, such as problem solving, communication, teamwork, etc. The DM consults a set $\mathcal{S}=\{1,\cdots,N\}$ of information sources before making a prediction about the true state. The signal generated by source $i$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}^{K}$ is the $K$-dimensional bias of the source, and $\varepsilon_{i,t}$ is a source-specific, idiosyncratic error with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{i}$. The state and errors are jointly independent as in the baseline model, although different dimensions of the state or an error can exhibit correlations.
The true bias vector is $b^{*}\coloneqq[b_{1}^{*},\cdots,b_{N}^{*}]^{\top}\in\mathbb{R}^{NK}$. Initially, the DM misspecifies the biases of the sources in =$\mathcal{M}=\{1,\cdots,M\}$ as $\tilde{b}_{i}\neq b_{i}^{*}$, $i\in\mathcal{M}$, while being uncertain of the biases of the remaining sources. The support of his prior is $\{b\in\mathbb{R}^{NK}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$. Everything else is as in the baseline model. In particular, the DM knows the true covariance matrix $\Sigma^{*}$ of the signals, which in the current setting equals $\text{diag}(\Omega_{1},\cdots,\Omega_{N})+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\otimes\Omega_{0}$ ($\otimes$ denotes the Kronecker product).
Key variables of our interest, including the distortion of DM’s long-run perceptions about source biases and that of his long-run behavior, are defined the same as before. The next proposition establishes the analog of Theorem 1 in the current multidimensional setting.
Proposition 3.
Redefine $\overline{\Delta}\coloneqq(\sum_{i=0}^{M}\Omega_{i}^{-1})^{-1}\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}.$ Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\emph{BLP}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
The proof of Proposition 3 replaces the Sherman-Morrison formula with the Woodbury formula for inverting block matrices (Woodbury, 1950). Given the close resemblance between $\overline{\Delta}$ and its single-dimensional counterpart, we have no doubt that its comparative statics are generally ambiguous. However, we do know that $\overline{\Delta}$ continues to satisfy DOOM and so is independent of $\Omega_{i}$, $i\in\mathcal{S}-\mathcal{M}$. We could also gain insight into simple economic problems such as the one presented below, based on the closed-form formula for $\overline{\Delta}$.
Example 3.
The DM faces a stream of female employees whose analytical skills and communication skills (hereafter denoted by $A$ and $B$, respectively) are drawn independently from the standard Normal distribution, i.e., $\Omega_{0}=\mathbb{I}_{2\times 2}$. The DM consults two sources about employees’ skills but misspecifies the biases of source one. Although the latter undervalues female employees’ analytical skills and overvalues their communication skills, the DM perceives it as impartial, i.e., $\Delta^{A}>0$ and $\Delta^{B}<0$. What the DM does know is that source one’s assessments of different skills exhibit a halo effect (Bohnet, 2016). That is, a positive impression about one’s analytical skill enhances the assessment of her communication skill, and vice versa. Let $\Omega_{1}=\small{\begin{bmatrix}v&v\rho\\ v\rho&v\end{bmatrix}}$ be the covariance matrix of the errors associated with source one’s signal. We capture the halo effect by letting $\rho>0$.
We examine how the DM’s misspecification about source one’s biases distorts his learning about source two’s biases, as well as his predictions about employees’ skills. Tedious algebra shows that
$\begin{bmatrix}\overline{\Delta}^{A}\\ \overline{\Delta}^{B}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{(1+v)\Delta^{A}-\rho v\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{+}\\ \underbrace{\frac{-\rho v\Delta^{A}+(1+v)\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{-}\end{bmatrix}\text{ and }\begin{bmatrix}\frac{\partial\overline{\Delta}^{A}}{\partial\rho}\\ \frac{\partial\overline{\Delta}^{B}}{\partial\rho}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{2\rho v^{2}(1+v)\Delta^{A}-v(1+v(2+v+\rho^{2}v))\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{+}\\ \underbrace{\frac{-v(1+v(2+v+\rho^{2}v))\Delta^{A}+2\rho v^{2}(1+v)\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{-}\end{bmatrix}.$
In words, the DM ends up perceiving source two as too lenient about women’s analytical skills and too critical about their communication skills. This misperception is exacerbated by the halo effect, whose presence requires that the DM subtract the communication score from the analytical score to tease out the fundamentals from the errors. It is through this channel that the DM’s unconscious, negative, bias about the communication score assigned by source one creeps into his learning about the analytical score assigned by source two, making the latter appear more positively biased than it actually is.
In the long run, the DM underestimates female employees’ analytical skills and overestimates their communication skills. To the extent that these estimates are key to HR decision making, such as task assignment, hiring, and retention, they can potentially reinforce occupational segregation and stereotypes that are pervasive in the workplace (see Blau and Kahn 2017 and Bertrand 2020 literature surveys). It is also natural for employees who anticipate such a misperception to overinvest in communication skills and to underinvest in analytical skills, thus fueling a vicious circle (Arrow, 1971). Our model uncovers the channels through which misspecified learning can potentially affect these labor market outcomes. A comprehensive exploration of these mechanisms is better left for future research.
While it is tempting to suggest, based on the simple model presented above, that companies focus rater training on correcting for halo effects, complications may arise in the presence of multiple misspecified sources, as correcting for the halo effect exhibited by one source makes those exhibited by other sources more salient. The total effect could be an exacerbation or a mitigation of the distortions, for the same reason that renders the comparative statics of the signal-to-noise ratio ambiguous in Theorem 1(ii).
4.3 Unknown covariance matrix
So far we have assumed that the covariance matrix of the signals is known to the DM from the outset. In environments such as large organizations with numerous divisions and hierarchies, a more natural assumption seems to be that the covariance matrix is also a subject of misspecified learning. We entertain this possibility in this section, with a focus on whether the additional source of misspecified learning leads to more biased beliefs and behavior on the part of the DM, or a mitigation thereof.
Formally, fix the covariance matrix of Gaussian errors $\varepsilon_{1,t}\cdots,\varepsilon_{N,t}$ to any $\Omega^{*}\in\mathbb{R}^{N\times N}$ that is symmetric and positive definite (SPD). Fix any $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$, [^9] Throughout the paper, we use $a\wedge b$ to denote the minimum of $a$ and $b$ and $a\vee b$ to denote the maximum of $a$ and $b$. and define $\Lambda$ as the set of $N\times N$ SPD matrices whose minimum eigenvalues are at least $\underline{\lambda}$. Initially, the DM holds a prior belief about the distribution of the signals. The prior is correct about signals equal to their biases plus the true state and state-independent errors. However, it misspecifies the biases of some sources as in Section 2, while allowing the covariance matrix of the errors to be anything in $\Lambda$. The support of the prior, again denoted by $\Theta$, is
$\begin{Bmatrix}(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\text{ for some }\Omega\in\Lambda\end{Bmatrix}.$
Importantly, $\Theta$ contains the true covariance matrix of the signals $\Sigma^{*}=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
We characterize outcomes of misspecified learning in two steps, beginning with the following lemma.
Lemma 4.
[…]
Let $(\hat{b},\widehat{\Sigma})$ denote the KL minimizer. Define the DM’s misperception of source biases, as well as the distortion of his long-run behavior, as before. The next theorem pinpoints the difference between these quantities and their analogs in the baseline model where the true covariance matrix is known to the DM from the outset.
Theorem 2.
Let everything be as above. Then
  1. (i)
    $\Delta$ is the same as in the case where the DM knows $\Sigma^{*}$ from the outset, whereas $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$;
  2. (ii)
    $\mathbb{E}_{(b^{*},\Sigma^{*})}[\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)]$ equals $\delta$ times its counterpart in the case where $\Sigma^{*}$ is known from the outset, where $\delta\coloneqq(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}\in(0,1)$.
  3. (iii)
[…]
Part (i) of Theorem 2, which characterizes the DM’s long-run belief, adapts Theorem A.1 Case (III) of HKS to the current setting where the covariance matrix of the signals takes the form of $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. The remaining parts of the theorem are new. On the one hand, misspecified learning of the covariance matrix does not interfere with that of source biases and leaves the DM’s long-run misperceptions about the latter unaffected. On the other hand, misspecified learning of source biases makes the signals appear noisier to the DM than they truly are, in that the perceived covariance matrix equals its true counterpart, plus a rank-one update $\Delta\Delta^{\top}$ that is positive definite. [^10] The perceived noisiness of the signals as measured by, e.g., the determinant of their covariance matrix, is larger with compound misspecified learning than without, i.e., $\text{det}(\widehat{\Sigma})=\text{det}(\Sigma^{*})(1+\Delta^{\top}\Sigma^{*-1}\Delta)>\text{det}(\Sigma^{*})$, where the equality follows from Ding and Zhou (2007). Consequently, when making predictions about the true state, the DM relies less on signal realizations compared to when he knows the true covariance matrix from the outset. This is shown to mitigate the impact of misspecified learning on his behavior, resulting in a less biased prediction of the state on average.
To better understand the bias mitigation effect, we decompose the distortion of the DM’s long-run behavior as follows:
$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ $=\mathbf{1}_{N}^{\top}\widehat{\Sigma}^{-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})$
$=\underbrace{-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta}_{\text{(I)}}+\underbrace{\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*})}_{\text{(II)}}\underbrace{-\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})\Delta}_{\text{(III)}}.$
Effect (I) in the above expression is due to the misspecified learning of source biases and is shown to equal $-\overline{\Delta}$ in the baseline model. Effect (II) arises from the misspecified learning of the covariance matrix and has zero mean under the true signal distribution. Effect (III) captures the compound effect of misspecified learning. It equals
[…]
Turning to comparative statics: In light of Theorem 1(ii), it is unsurprising that increasing signal-to-noise ratios could have unintended consequences on the outcomes of misspecified learning. What is new here is the ambiguity associated with correcting misspecifications, whose comparative statics in the baseline model were straightforward. The ambiguity arises from compound misspecified learning: By correcting the misspecifications associated with familiar sources, we make the environment appear less noisy to the DM than before. This reduces the bias mitigation effect discussed above, i.e., $\delta=(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ increases, and in some cases could undo the benefit of misspecification corrections.
Theorem 2 sheds light on the behavioral and welfare consequences of the debiasing programs introduced toward the end of Section 3.2. Years of economic and psychological research have yielded mixed insights into the efficacy of these programs. For example, while the success of using one’s own mental resources to intercept implicit biases has been documented in prominent case studies (recall Footnote 8), recent meta-analysis of organizations that have implemented programs of similar kinds reveals a mixed pattern, leading Greenwald and Lai (2020) to conclude that “Convincing evidence for the effectiveness of these strategies is not yet available in peer-reviewed publications.” Even worse, efforts to suppress stereotypes or to reduce intergroup animosity may backfire (Kulik et al., 2000; Kalev et al., 2006), leading scholars including Bohnet (2016) to advocate alternative strategies such as evidence-based continuing learning. We take no stand on this debate and instead caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect.
5 Further applications
In this section, we utilize our theoretical framework to analyze two further applications: misperceived media bias and project investment with misspecified learning.
5.1 Misperceived media bias
A Republican voter consults several news sources on an evolving issue. Some sources, such as the voter’s personal source and traditional news outlets, carry biases that he feels familiar with. Others, such as social media sites operated by unknown people or bots, are unfamiliar. Initially, the voter is uncertain about the horizontal biases of the unfamiliar sources, while holding dogmatic, wrong beliefs about the biases of some familiar sources. As someone who grew up watching Fox News, he believes that this source is more neutral than it actually is due to familiarity bias (Pennycook and Rand, 2021). He underestimates the bias associated with his own opinions and judgments, in part due to the bias blind spot and third-person effect (Pronin et al., 2002; Davison, 1983), and in part because he is working toward self-enhancement and would like to view himself as middle-of-the-road (Dennis, 1988). [^11] Bias blind spot refers to the tendency for people to recognize biases in human judgment except when that bias is their own. The third-person effect predicts that people perceive mass media messages to have a greater effect on others than on themselves. Finally, he believes in a liberal media bias like many other Republicans and displays excessive hostility towards CNN (Hassell et al., 2020). These cognitive biases are known to enhance with age, be resistant to corrective information, or even pass down through generations (Pronin et al., 2002; Jennings et al., 2009). They can be captured by letting ={personal source,Fox $\mathcal{M}=\{\text{personal source},\text{Fox News},\text{CNN}\}$ and $\tilde{b}_{i}<b_{i}^{*}$ $\forall i\in\mathcal{M}$ in our framework.
In each subsequent period $t=1,2,\cdots$, the voter takes a position $a_{t}\in\mathbb{R}$ and incurs a loss $-(a_{t}-\omega_{t})^{2}$ equal to the quadratic distance between this position and the true state of the world. The optimal position in the long run equals $\text{BLP}^{*}(X)$ under the correctly specified model and $\widehat{\text{BLP}}(X)$ under the misspecified model. The expected utilities generated by these positions, evaluated under the true signal distribution, equal $\text{EU}^{*}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}\left(\omega-\text{BLP}^{*}(X)\right)^{2}$ and $\widehat{\text{EU}}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}(\omega-\widehat{\text{BLP}}(X))^{2}$, respectively. The difference between them captures the welfare loss from misspecified learning.
Proposition 4.
$\widehat{\text{EU}}-\text{EU}^{*}$ equals $-\overline{\Delta}^{2}$ in the baseline model and $-\delta\overline{\Delta}^{2}$ in the environment described in Section 4.3.
Proposition 4, together with Theorems 1 and 2, fully describes the distortionary effects of misspecified learning on the voter’s long-run beliefs, behavior, and welfare. Since $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}<0$ $\forall i\in\mathcal{M}$ in the current example, $\overline{\Delta}<0$. The distortion of the voter’s worldview due to misspecifications constitutes a liberal media bias of magnitude $|\overline{\Delta}|$. Our results establish a mechanism through which this distortion can spill into the online news sphere and facilitate analysis of its behavioral and welfare consequences.
Hostile media effect and false polarization.
The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is. Evidence supporting the existence of this phenomenon abounds in political science, communication, and psychology (Feldman, 2014), [^12] In a classical experiment conducted by Arpan and Raney (2003), student subjects read a balanced story about their hometown college football team in one of three newspapers: home town, rival town, or neutral town outside the state. Subjects viewed the coverage as favoring the rival university, even if it was believed to be from the neutral town. More recently, Hassell et al. (2020) present journalists with balanced news stories that differ only in their ideological content (e.g., the party affiliation of the candidate running for office). Journalists exhibit no gatekeeping bias when selecting news stories to cover, despite the fact that many of them are ideologically liberal, thus debunking the “liberal media bias” that is widely perceived among Republicans. although research on learning and belief formation about new online sources has remained sparse. We demonstrate how the hostile media effect in traditional media can spill into the digital news sphere, as the voter in our model overperceives the bias of any liberal source while underperceiving the bias of any conservative source in the long run.
False polarization is defined as the difference between the perceived distance between the positions of two groups and the actual distance between them. Its presence in the American public was rigorously established by Levendusky and Malhotra (2016b). [^13] These authors used the survey data gleaned from a nationally representative sample that asked about the subjects’ self-placements on an issue scale, as well as their placements of a typical Democratic and Republican voter on that scale. The data revealed that Americans significantly misperceive the public to be more divided along partisan lines than they are in reality. To the extent that individual human beings can now express their opinions online at almost no cost, our model predicts a false polarization in the digital news sphere, as well as a causal relationship between misspecified news bias and false polarization. The latter is consistent with the literature’s common attribution of false polarization to partisan media exposure (Levendusky and Malhotra, 2016a).
The existing literature is divided on the impact of (mis)perceived polarization and media bias on voters’ actual positions, with some authors such as Levendusky and Malhotra (2016a) documenting a moderation effect caused by false polarization. Our results suggest that this finding may be reversed as time goes by, as the voter in our model behaves more conservatively than under the true model in the long run.
Fact-checking.
DOOM stipulates that sources with initially unknown biases are irrelevant to the distortionary effect of misspecified learning. As we increase the signal-to-noise ratios of these sources, the voter’s expected utility under the true model and, hence, the misspecified model, increases unambiguously ($\text{EU}^{*}=-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $\widehat{\text{EU}}=\text{EU}^{*}-\overline{\Delta}^{2}$). The same cannot be said about sources with misspecified biases, as increasing their signal-to-noise ratios generally has an ambiguous effect on the welfare distortion $\overline{\Delta}^{2}$ from misspecified learning.
Our finding suggests that policy interventions that target misinformation and fake news through improving source accuracy (e.g., fact-checking) may entail unintended welfare consequences, especially when they operate by modulating sources that people already hold dogmatic misperceptions about. When it comes to lesser-known sources, fact-checking always improves welfare. To the extent that professional fact-checking is costly and does not scale easily, the allocation of time and efforts should prioritize lesser-known sources over established ones. This finding is, to our best knowledge, new to the literature on (political) misinformation (see Nyhan 2020 for a survey).
Misperception recalibration.
A large empirical literature studies the treatment effects of misperception recalibration (Bursztyn and Yang, 2022). While early research indicated the possibility of a “backfire effect” whereby recalibration efforts (e.g., provision of corrective information, cross-cutting exposure) end up reinforcing rather than reducing subjects’ misperceptions, an emerging research consensus finds that the backfire effect could be more tenuous than previously thought and that more attention should be paid to alternative issues, such as the durability and scalability of the interventions (Nyhan, 2021). We caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect studied in Section 4.3.
5.2 Project investment with misspecified learning
A manager invests an initial wealth of $W>0$ in $K>1$ projects whose returns $\omega_{k}$, $k=1,\cdots,K$ are independently drawn from the standard normal distribution. Prior to decision making, he inquires about the return of each project from a mix of familiar sources — his personal source, internal consultants — and unfamiliar sources — external consultants, robo-advisors. The signal generated by each source is the same as in the baseline model, with its bias now capturing how optimistic the source is about the return of the project.
[…]
Over time, the manager observes signal realizations and updates his belief about source biases. His long-run belief underestimates the biases of sources that are optimistic about the returns of projects, while overestimating the biases of those that are neutral or pessimistic. We take this belief as given and examine its consequences for the manager’s investment decisions and welfare. For ease of notation, we assume, in the remainder of this section, that only sources that report on project $k=1$ have their biases misspecified by the manager. In addition, the manager has a CARA Bernoulli utility function with an Arrow-Pratt measure $\lambda>0$ of absolute risk aversion.
Specifically, for each project $k=1,\cdots,K$, we use $\mu_{k}^{*}$ to denote a typical realization of the best linear predictor of its return under the correctly specified model. The counterpart of this quantity under the misspecified model is denoted by $\widehat{\mu}_{k}$. The variance $\eta_{k}^{-1}$ of the best linear predictor is known to decrease with the signal-to-noise ratios of the signals that report on project $k$. For any given profile of $\mu_{k}$, $k=1,\cdots,K$, the optimal investment decisions, denoted by $\alpha_{k}$, $k=1,\cdots,K$, solve for
$\max_{\alpha_{k},k=1,\cdots,K}\sum_{k=1}^{K}\alpha_{k}\mu_{k}-\frac{\lambda}{2}\sum_{k=1}^{K}\alpha_{k}^{2}\eta_{k}^{-1}\text{ s.t. }\sum_{k=1}^{K}\alpha_{k}\leq W.$
Let $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ denote the optimal portfolio choices under the correctly specified model and the misspecified model, respectively. The difference $\Delta_{\alpha_{k}}\coloneqq\widehat{\alpha}_{k}-\alpha_{k}^{*}$ captures the distortionary effect of misspecified learning on the investment in project $k$. The expected utilities generated by $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ are denoted by $\text{EU}^{*}$ and $\widehat{\text{EU}}$, respectively. Their difference captures the distortionary effect of misspecified learning on the manager’s welfare.
Proposition 5.
$\Delta_{\alpha_{1}}=-\frac{\eta_{1}}{\lambda}\left(1-\frac{\eta_{1}}{\sum_{k=1}^{K}\eta_{k}}\right)\overline{\Delta},\text{ }\Delta_{\alpha_{j}}=\frac{\eta_{1}\eta_{j}}{\lambda\sum_{k=1}^{K}\eta_{k}}\overline{\Delta}\text{ }\forall j\neq 1,$
[…]
and their absolute values are all decreasing in $\eta_{k}^{-1}$, $k=1,\cdots,K$.
We perform four comparative static exercises. Consider first the effect of expanding the set $\mathcal{S}-\mathcal{M}$ of unfamiliar sources, due to the advent of robo-advisors or the popular advice that projects should be evaluated by “uninvolved managers” to counteract entrenched biases in managerial decision making (Guenzel and Malmendier, 2020). This change has no impact on $\overline{\Delta}$ by DOOM while clearly reducing the errors associated with the best linear predictors (as captured by $\eta_{k}^{-1}$, $k=1,\cdots,K$). Consequently, it mitigates the distortionary effects of misspecified learning on the manager’s decision making and expected utility.
Consider next the effect of increasing the signal-to-noise ratios of the sources, which causes $\eta_{k}^{-1}$, $k=1,\cdots,K$ to decrease. Although this change unambiguously mitigates the distortions of the manager’s decisions and expected utility, holding $\overline{\Delta}$ fixed, its impact on $\overline{\Delta}$ is generally ambiguous. This suggests that technological advances that lower data collection and processing costs (e.g., the AI revolution) could have differential impacts on our manager, depending on whether they operate by increasing the signal-to-noise ratios of familiar sources or those of novel sources.
Consider next corrective measures aimed at bringing misspecifications closer to zero, which in reality range from debiasing board members who are prone to select and retain like-minded CEOs (Guenzel and Malmendier, 2020) to nudges such as premortem, i.e., imagine all twists and turns (Cohee and Barnhart, 2023). To the extent that the comparative statics of $\overline{\Delta}$ against these measures are generally ambiguous in the presence of compound misspecified learning, their impacts on the manager’s decision making and welfare can be rather nuanced.
Lastly, consider the effect of expanding the portfolio of the projects overseen by the manager. Since this change mechanically increases $\sum_{k=1}^{K}\eta_{k}$, it exacerbates the distortions of the manager’s decisions and expected utility when the distributions of the signals that report on the additional projects are correctly specified. In the presence of misspecified learning, adding projects with independent returns to the portfolio creates the illusion that the optimal portfolio is well diversified. Consequently, more funds are allocated to projects the manager is overconfident about, thereby exacerbating the distortions of his decisions and welfare.
6 Concluding remarks
This paper takes the first step toward understanding the outcomes of Bayesian learning in the presence of multiple misspecified information sources. We propose a framework that encompasses a broad set of economic phenomena and develop a methodology that delivers tractable results and new insights. Several questions remain open.
The first question concerns the utilization of current findings as pivotal steps in approaching new inquiries. For example, how do managerial misperceptions interact with minority employees’ decisions to undertake productivity-enhancing investment? How does the polarization of voters’ behavior influence politicians’ choices of platforms and policies? How are equilibrium asset prices affected by distortions in individual investors’ portfolio choices? Successfully addressing these questions requires that we incorporate misspecified active learning into the current framework and tackle additional theoretical challenges such as the convergence of equilibria.
As discussed in Section 1.1, our results on labor market discrimination are distinctive from those due to tastes or rational beliefs, prompting the need for experimental validation. Some of the experimental designs proposed in Section 1.1, such as testing the comparative statics regarding informational interventions, have already been implemented by Bohren et al. (2023), albeit in a simpler decision environment devoid of misspecified learning about unfamiliar sources. Testing models of misspecified learning can be challenging, although successful endeavors have been undertaken previously; see, e.g., Bohren et al. (2019) for using the dynamics of beliefs to identify the source of discrimination. Considering the state of the literature, an empirical evaluation of our theory by means of experiments seems both interesting and feasible.
Outside the realm of misspecified Bayesian learning, misperceptions are often explained by learning frictions such as non-Bayesian updating or errors in reasoning. Ortoleva (2022) provides an up-to-date survey of these models and their economic applications. In contrast, our DM has a misspecified prior but is otherwise a standard Bayesian. We view the two approaches as complementary. Their connection is explored by Bohren and Hauser (2023), who provide conditions when a non-Bayesian updating rule has a misspecified-learning representation.
Appendix A Proofs
A.1 Proof outline
We begin by describing the problem studied by HKS and discuss the relations between our analyses. HKS examine a problem of misspecified Bayesian learning in which the DM observes in each period $t=1,2,\cdots$ a $D$-dimensional Gaussian signal with mean $\Gamma f\in\mathbb{R}^{D}$ and covariance matrix $\Sigma\in\mathbb{R}^{D\times D}$. Initially, the DM knows the transformation matrix $\Gamma\in\mathbb{R}^{D\times N}$ while misspecifying the first entry of the fundamental $f\in\mathbb{R}^{N}$. He may or may not know the true covariance matrix $\Sigma^{*}$, depending on whether he is in Case (I) or Case (III). [^14] Case (II) of HKS captures situations in which the DM holds a degenerate prior belief about the fundamental value. The support of his prior belief is
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma=\Sigma^{*}\}$
in Case (I). In Case (III), it becomes
$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}}\},$
where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$. Over time, the DM updates his belief about $(f,\Sigma)$ in a Bayesian manner based on the prior and signal realizations. For both cases, HKS demonstrate that the DM’s limiting posterior belief concentrates on the KL minimizer and provide characterizations for the latter.
The current analysis differs from that of HKS in the following aspects.
  1. D1
    We allow the DM to misspecify multiple entries of the fundamental, while HKS focus on the case of a single misspecified entry.
  2. D2
    We are interested in the exercise of predicting the true state of the world based on signal realizations. To this end, we assume that signals equal biases plus the true state, plus state-independent errors with a covariance matrix $\Omega$.
  3. D3
    $\Omega$ is taken to be arbitrary when we solve the KL minimizer and prove that it is the limit point of the DM’s posterior belief. When applying our results to derive economic insights in more specific applications, we restrict $\Omega$ to be low-rank updates of diagonal matrices.
How do these differences impact the analysis? For establishing the convergence of the limiting posterior belief to the KL minimizer, it turns out that D1 does not matter, whereas the differences due to D2 are quite nuanced. We highlight them in Appendix A.2 when we replicate the convergence proof of HKS line by line.
Our solution to the KL minimization problem nests that of HKS due to D1. Our method exploits the first-order conditions associated with KL minimization, together with the block matrix inversion formula. Its difference from HKS’ guess-and-verify approach is underscored in Appendix A.3 when we present the proof of Lemma 2.
The restriction of $\Omega$ to low-rank updates of diagonal matrices is suitable for the economic applications of our interest. The proofs presented in Appendix A.4 exploit existing mathematical results on low-rank updates of invertible matrices.
We follow the notational conventions developed in the main text. Specifically, $x$ denotes the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ denotes the diagonal matrix created from a vector, $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$, and $\mathbb{I}_{k\times k}$ denotes the $k\times k$ identity matrix. $\otimes$ denotes the Kronecker product. $\max\{a,b\}$ is written as $a\vee b$; $\min\{a,b\}$ as $a\wedge b$.
A.2 Proofs of the concentration of limiting posterior belief on KL minimizer
This appendix gathers proofs for the concentration of the DM’s limiting belief on the set of KL minimizers. We begin by presenting a useful lemma.
Lemma 5.
Let $G\in\mathbb{R}^{N\times N}$ be any symmetric matrix and $\lambda_{k}(G)$ denote the $k^{th}$ largest eigenvalue of $G$, $k=1,\cdots,N$. The following are true for any $x\in\mathbb{R}^{N}$:
$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+x^{\top}x)$
[…]
Proof.
Weyl’s inequality (Weyl, 1912) states that $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-N}(A+B)$ for any Hermitian matrices $A$ and $B$ that belong to an inner product space of dimension $N$. Letting $A=G$ and $B=xx^{\top}$ in Weyl’s inequality yields
$\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)+\lambda_{N}(xx^{\top})=\lambda_{1}(G)+0\leq\lambda_{1}(G+xx^{\top})$ (4)
[…]
where the equalities use the fact that $xx^{\top}$ has rank one and so has two eigenvalues: 00 and $\text{tr}(xx^{\top})$, together with $\text{tr}(xx^{\top})=\text{tr}(x^{\top}x)=x^{\top}x>0$. Interchanging the roles of $G+xx^{\top}$ and $G$ in the above derivation gives us the following:
$\lambda_{N}(G)\leq\lambda_{1}(G+xx^{\top})+\lambda_{N}(-xx^{\top})=\lambda_{1}(G+xx^{\top})-x^{\top}x\leq\lambda_{1}(G)$ (6)
[…]
Together, (4), (5), and (7) imply that
$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+xx^{\top}),$
that is the first inequality in the statement of the lemma. (4), (5), and (6) imply that
$\lambda_{1}(G)\vee(\lambda_{N}(G)+xx^{\top})\leq\lambda_{1}(G+xx^{\top})\leq\lambda_{1}(G)+x^{\top}x,$
that is the second inequality in the statement of the lemma. ∎
Proof of Lemma 4.
In the environment considered in 4.3, the support of the DM’s prior belief is
$\Theta=\begin{Bmatrix}(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\text{ for some }\Omega\in\Lambda\end{Bmatrix},$
where $\Lambda$ denotes the set of $N\times N$ SPD matrices whose minimum eigenvalues are bounded below by $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$. The proof presented below verifies that this environment satisfies the assumptions of Berk (1966), hence the DM’s limiting posterior belief concentrates on the set of KL minimizers.
Part (i) of Berk’s assumption says that the subjective density of the signals:
$l(x\mid b,\Sigma)=\frac{1}{\sqrt{(2\pi)^{N}\text{det}\Sigma}}\exp\left(-\frac{1}{2}(x-b)^{\top}\Sigma^{-1}(x-b)\right),$
is continuous in $(b,\Sigma)$ on $\Theta$. This is clearly true, since $\text{det}\Sigma$ and $\Sigma^{-1}$ are both continuous in $\Sigma$.
Part (ii) of Berk’s assumption says that the subjective density equals zero on a set of measure zero under the true signal distribution. This is clearly true, since the subjective density is always positive.
Part (iii) of Berk’s assumption says for every $(b,\Sigma)\in\Theta$, there exists an open neighborhood $U\subset\Theta$ of $(b,\Sigma)$ such that
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in U}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]<\infty.$
Let $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ denote the minimum eigenvalue and the maximum eigenvalue of $\Sigma^{\prime}$, respectively. Then
$|\log l(x\mid b^{\prime},\Sigma^{\prime})|$ $=|\frac{1}{2}\log((2\pi)^{N}\text{det}\Sigma^{\prime})+\frac{1}{2}(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})|$
$\leq\frac{1}{2}\left(N|\log(2\pi\lambda_{\max}(\Sigma^{\prime}))|\vee|\log(2\pi\lambda_{\min}(\Sigma^{\prime}))|+\frac{1}{\lambda_{\min}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right),$
where the inequality uses the facts that the determinant of a matrix equals the product of its eigenvalues, the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded above by $\lambda_{\max}(\Sigma^{\prime-1})=1/\lambda_{\min}(\Sigma^{\prime})$, and $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$. Since $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ are continuous in the entries of $\Sigma^{\prime}$, the last expression is finite on every finite neighborhood $U$.
The proof so far is the same as that of HKS’ Theorem A.1 Case (III). The last step of the proof differs slightly due to our restriction of the covariance matrix to those of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. Part (iv) of Berk’s assumption states that for every $c\in\mathbb{R}$, there exists $D\subset\Theta$ with a compact complement $\Theta-D$ such that
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in D}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]\leq c.$
Fix any $k_{1},k_{2}>0$, and let $D$ be the set of $(b^{\prime},\Sigma^{\prime})$ such that $\|b^{\prime}-b^{*}\|>k_{1}$ and the corresponding $\Omega^{\prime}$ satisfies $\lambda_{\max}(\Omega^{\prime})>k_{2}$. By Lemma 5, $\lambda_{\max}(\Sigma^{\prime})\geq\lambda_{\max}(\Omega^{\prime})(>k_{2})$ and $\lambda_{\min}(\Sigma^{\prime})\geq\lambda_{\min}(\Omega^{\prime})(\geq\underline{\lambda})$. Substituting these bounds into HKS’ proof of their Theorem A.1 Case (III) and replicating line by line, we obtain the following:
$\log l(x\mid b^{\prime},\Sigma^{\prime})$
$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee\|x-b^{\prime}\|^{2})\right),$
where the first inequality uses the facts that $\text{det}\Sigma^{\prime}\geq\lambda_{\min}^{N-1}(\Sigma^{\prime})\lambda_{\max}(\Sigma^{\prime})$, $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$, and the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded below by $\lambda_{\min}(\Sigma^{\prime-1})=1/\lambda_{\max}(\Sigma^{\prime})$. The second inequality follows from minimizing $\log\lambda_{\max}(\Sigma^{\prime})+\|x-b^{\prime}\|^{2}/\lambda_{\max}(\Sigma^{\prime})$ over the range $\lambda_{\max}(\Sigma^{\prime})>k_{2}$. To obtain a lower bound for $\|x-b^{\prime}\|^{2}$, we note that $x-b^{\prime}=x-b^{*}+(b^{*}-b^{\prime})=(\omega+\varepsilon)+(b^{*}-b^{\prime})$ and so $\|x-b^{\prime}\|^{2}\geq\|\omega+\varepsilon\|^{2}+\|b^{*}-b^{\prime}\|^{2}-2\|\omega+\varepsilon\|\|b^{*}-b^{\prime}\|$ by the Cauchy-Schwarz inequality. Since $\|b^{*}-b^{\prime}\|>k_{1}$ by assumption, it follows that for any given value of $\|\omega+\varepsilon\|<k_{1}/2$, $\|b^{*}-b^{\prime}\|^{2}+\|\omega+\varepsilon\|^{2}-2\|b^{*}-b^{\prime}\|\|\omega+\varepsilon\|$ is minimized at $\|b^{*}-b^{\prime}\|=k_{1}$ and so is bounded below by $k_{1}^{2}/4$. Consequently, the last line of the above inequality is bounded above by
$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right).$
Meanwhile, $\log l(x\mid b^{\prime},\Sigma^{\prime})$ is bounded above by
$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
$\leq-\frac{1}{2}N\log\lambda_{\min}(\Sigma^{\prime})\leq N|\log\underline{\lambda}|/2$
when $k_{2}(<\lambda_{\max}(\Sigma^{\prime}))$ is large, where the second inequality uses the bound that $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}$ and the assumption that $\underline{\lambda}<1$. Together, we obtain the following:
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in\Theta}\log l(x\mid b^{\prime},\Sigma^{\prime})\right]\leq\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\mathbf{1}_{\|\omega+\varepsilon\|\leq k_{1}/2}\sup_{(b^{\prime},\Sigma^{\prime})\in D}l(x\mid b^{\prime},\Sigma^{\prime})\right]+N|\log\underline{\lambda}|/2$
$\leq-\frac{1}{2}\left[N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right]\mathbb{P}(\|\omega+\varepsilon\|\leq k_{1}/2)+N|\log\underline{\lambda}|/2,$
where the last line becomes arbitrarily negative as $k_{1},k_{2}\rightarrow+\infty$.
It remains to verify that $\Theta-D$ is compact. By definition, this set consists of all vectors $b^{\prime}$ such that $\|b^{\prime}-b^{*}\|\leq k_{1}$, together with all covariance matrices of the form $\Omega^{\prime}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, whereby the eigenvalues of $\Omega^{\prime}$ belong to $[\underline{\lambda},k_{2}]$. The last set is compact, and this completes the proof. ∎
Proof of Lemma 1.
In the environment considered in the baseline model, the support of the DM’s prior belief is
$\Theta=\{(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\}$
where $\Omega^{*}$ denotes the true covariance matrix of Gaussian errors. Since $\Omega^{*}\in\Lambda$, all we need is to replace every $\Sigma$ and $\Sigma^{\prime}$ in the proof of Lemma 4 with $\Sigma^{*}\coloneqq\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$. ∎
A.3 Proof of Lemma 2
Since $\Sigma^{*}$ is a positive definite matrix, it is invertible. Applying the block matrix inversion formula to $\Sigma^{*}$ yields
$\Sigma^{*-1}=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right],$
where $A$, $B$, $C$, and $D$ were defined in Section 3.3. The system (3) of first-order conditions associated with KL minimization, in its matrix form, is
$-(D-CA^{-1}B)^{-1}CA^{-1}\widetilde{\Delta}+(D-CA^{-1}B)^{-1}\widehat{\Delta}=\mathbf{0}_{N-M}.$
Multiplying both sides of this equation by $D-CA^{-1}B$ yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ as desired.
Using this result to simplify $\Sigma^{*-1}\Delta$, we obtain the following:
$\Sigma^{*-1}\Delta$ $=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right]\left[\begin{array}[]{c}\widetilde{\Delta}\\ \hline\cr\widehat{\Delta}\end{array}\right]$ (12)
$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}+A^{-1}B(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right]$ (14)
$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right].$ (17)
Simplifying $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta$ accordingly yields $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$ as desired. ∎
A.4 Proofs of results that characterize KL minimizers
Proof of Proposition 1.
The environment described in the proposition features $\Sigma^{*}=\text{diag}(v)+\alpha\alpha^{\top}$, $A=\text{diag}(v_{1},\cdots,v_{M})+[\alpha_{1},\cdots,\alpha_{M}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$ and $C=[\alpha_{M+1},\cdots,\alpha_{N}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$. Since $A$ is a rank-one update of an invertible matrix,
$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\begin{bmatrix}\alpha_{1}\nu_{1}\\ \vdots\\ \alpha_{M}\nu_{M}\end{bmatrix}\begin{bmatrix}\alpha_{1}\nu_{1}&\cdots&\alpha_{M}\nu_{M}\end{bmatrix}$
by the Sherman-Morrison formula. Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
$\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\overline{\Delta},$
where $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i})$. To calculate the distortion in the best-linear predictor, multiply (17) by $\alpha^{\top}$ to obtain
$\alpha^{\top}\Sigma^{*-1}\Delta=[\alpha_{1},\cdots,\alpha_{M}]A^{-1}\widetilde{\Delta}=\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\overline{\Delta}.$
Since $\text{BLP}^{*}(X)=\alpha^{\top}\Sigma^{*-1}(X-b^{*})$ and $\widehat{\text{BLP}}(X)=\alpha^{\top}\Sigma^{*-1}(X-\hat{b})$ in the current environment, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\alpha^{\top}\Sigma^{*-1}\Delta=-\overline{\Delta}$ as desired. ∎
Proof of Proposition 2.
With a common latent factor, the KL minimization problem becomes
$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
Since $\Sigma^{*}=\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, $\Sigma^{*-1}=\text{diag}(\nu)-\nu\nu^{\top}/(1+\mathbf{1}_{N}^{\top}\nu)$ by the Sherman-Morrison formula. Thus
$\Phi=\Gamma^{\top}\text{diag}(\nu)\Gamma-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\Gamma^{\top}\nu\nu^{\top}\Gamma,$
where the first part of the right-hand side equals (recall that $\Gamma=[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$)
$\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\text{diag}(\nu)\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right]=\left[\begin{array}[]{c|c}\text{diag}(\nu)&\nu\\ \hline\cr\nu^{\top}&\mathbf{1}_{N}^{\top}\nu\end{array}\right],$
and the second part equals
$\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\nu\nu^{\top}\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right].$
Subtracting the second part from the first yields
$\Phi=\left[\begin{array}[]{c|c}\text{diag}(\nu)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\nu^{\top}&\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\\ \hline\cr\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu^{\top}&1-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\end{array}\right].$
The system of first-order conditions associated with KL minimization is
$E\begin{bmatrix}\Delta_{M+1}\\ \vdots\\ \Delta_{N+1}\end{bmatrix}=-F\begin{bmatrix}\Delta_{1}\\ \vdots\\ \Delta_{M}\end{bmatrix},\text{ where }$
$E\coloneqq\begin{bmatrix}\Phi_{M+1,M+1}&\cdots&\Phi_{M+1,N+1}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,M+1}&\cdots&\Phi_{N+1,N+1}\end{bmatrix}\text{ and }F\coloneqq\begin{bmatrix}\Phi_{M+1,1}&\cdots&\Phi_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,1}&\cdots&\Phi_{N+1,M}\end{bmatrix}.$
From the expression for $\Phi$, we can read off those for $E$ and $F$:
$E=\text{diag}(\nu_{M+1}\cdots,\nu_{N},1)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},&\cdots,&\nu_{N},&-1\end{bmatrix}$
and
$F=-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
Since $E$ is a rank-one update of an invertible matrix, applying the Sherman-Morrison formula shows that
$E^{-1}=\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)$
$+\frac{\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)/(1+\mathbf{1}_{N}^{\top}\nu)}{1-\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}/(1+\mathbf{1}_{N}^{\top}\nu)}$
$=\text{diag}(v_{M+1},\cdots,v_{N},1)+\frac{1}{\sum_{i=1}^{M}\nu_{i}}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right]\left[\begin{array}[]{c|c}\mathbf{1}_{N-M}^{\top}&-1\end{array}\right].$
Simplifying the system of first-order conditions accordingly yields
$\left[\begin{array}[]{c}\Delta_{M+1}\\ \vdots\\ \Delta_{N}\\ \hline\cr\Delta_{N+1}\end{array}\right]=\overline{\Delta}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right],\text{ where }\overline{\Delta}\coloneqq\frac{\sum_{i=1}^{M}\nu_{i}\Delta_{i}}{\sum_{i=1}^{M}\nu_{i}}.\qed$
Proof of Proposition 3.
With multidimensional skills, the true covariance matrix of the signals becomes
$\Sigma^{*}=\text{diag}\left(\Omega_{1},\cdots,\Omega_{N}\right)+\left(\mathbf{1}_{N}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{N}^{\top}\otimes\mathbb{I}_{K\times K}\right),$
where $\otimes$ denotes the Kronecker product. Write $\Sigma^{*}$ as
$\left[\begin{array}[]{c|c}A_{KM\times KM}&B_{KM\times K(N-M}\\ \hline\cr C_{K(N-M)\times KM}&D_{K(N-M)\times K(N-M)}\end{array}\right],$
where
$A=\underbrace{\text{diag}(\Omega_{1},\cdots,\Omega_{M})}_{G}+\underbrace{\left(\mathbf{1}_{M}\otimes\mathbb{I}_{K\times K}\right)}_{x}\Omega_{0}\underbrace{\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right)}_{y}$
and
$C=\left(\mathbf{1}_{N-M}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right).$
Since $A$ is a rank-$K$ update of an invertible matrix,
$A^{-1}$ $=G^{-1}+G^{-1}x\left(\Omega_{0}^{-1}+yG^{-1}x\right)^{-1}yG^{-1}$
$=\text{diag}(\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1})-\begin{bmatrix}\Omega_{1}^{-1}\\ \vdots\\ \Omega_{M}^{-1}\end{bmatrix}\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\begin{bmatrix}\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1}\end{bmatrix}$
where the first equality uses the Woodbury formula (Woodbury, 1950). Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
$\widehat{\Delta}$ $=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\otimes\left(\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}\right)$
$=\mathbf{1}_{N-M}\otimes\overline{\Delta},$
where the second line uses the fact that
$\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=0}^{M}\Omega^{-1}_{i}=\Omega_{0}\Omega_{0}^{-1}=\mathbb{I}_{K\times K}$
and so
$\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}=\left(\sum_{i=0}^{M}\Omega_{i}\right)^{-1}.$
The distortion in the best linear predictor can be obtained likewise, as follows:
$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ $=-\left(\mathbf{1}_{M}^{\top}\otimes\Omega_{0}\right)A^{-1}\widetilde{\Delta}$
$=-\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}$
$=-\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\sum_{i=1}^{M}\Omega_{i}^{-i}\Delta_{i}=-\overline{\Delta}.\qed$
Proof of Theorem 2.
Part (i): The KL miminization problem considered in Case (III) of HKS (recall Appendix A.1) is
$\min_{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}}D_{KL}(f^{*},\Sigma^{*}\mid f,\Sigma)\text{ s.t. }f_{1}=\tilde{f}_{1}\text{ and }\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}},$
where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$. The solution $(\hat{f},\widehat{\Sigma})$ to this problem is shown to satisfy two properties: First, $\Delta f_{i}\coloneqq\hat{f}_{i}-f_{i}^{*}$ is the same as when the DM knows the true covariance matrix $\Sigma^{*}$ from the outset. Second, $\widehat{\Sigma}=\Sigma^{*}+(\Gamma\Delta f)(\Gamma\Delta f)^{\top}$. The assumption that only a single entry of $f$ is misspecified plays no role in this result. This can be verified by replicating the proof of HKS line by line.
Letting $f=b$ and $\Gamma=\mathbb{I}_{N\times N}$ in the above argument leads us to conclude that in the environment outlined in Section 4.3, the solution to the KL minimization problem would be $\Delta=\hat{b}_{i}-b_{i}^{*}$ and $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, had our DM been able to choose any SPD matrix whose minimum eigenvalue is bounded below by $\underline{\underline{\lambda}}$. The additional constraint we impose is that the covariance matrix must be of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, where $\Omega$ is PSD with its minimum eigenvalue bounded below by $\underline{\lambda}<1\wedge\lambda_{\min}(\Omega^{*})$. We are left to verify that the solution to the unconstrained problem satisfies this additional constraint, and so must also solve the constrained optimization problem.
To verify, note that the solution to the unconstrained problem: $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, is attained by $\widehat{\Omega}=\Omega^{*}+\Delta\Delta^{\top}$. $\widehat{\Omega}$ is clearly PSD, and $\lambda_{\min}(\widehat{\Omega})\geq\lambda_{\min}(\Omega^{*})>\underline{\lambda}$ by Lemma 5. This completes the proof of Part (i) of the theorem.
Part (ii) of the theorem follows from the decomposition exercise conducted in Section 4.3. Part (iii) follows from straightforward algebra. ∎
Proof of Proposition 4.
Consider first the setting examined in Section 3. There, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, and so
$\widehat{\text{EU}}-\text{EU}^{*}$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[-(\widehat{\text{BLP}}-\omega)^{2}+(\text{BLP}^{*}-\omega)^{2}]$
$=\mathbb{E}_{(b^{*},\Sigma^{*})}[-(-\overline{\Delta}+\text{BLP}^{*}-\omega)^{2}+(\text{BLP}^{*}-\omega)^{2}]$
$=-\overline{\Delta}^{2}+2\overline{\Delta}\mathbb{E}_{(b^{*},\Sigma^{*})}[\text{BLP}^{*}-\omega]$
$=-\overline{\Delta}^{2}.$
Consider next the setting examined in Section 4.3. There,
$\widehat{\text{EU}}-\text{EU}^{*}$ $=-\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\widehat{\text{BLP}})^{2}]+\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})^{2}]$
$=-\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*}-(\widehat{\text{BLP}}-\text{BLP}^{*}))^{2}]+\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})^{2}]$
$=-\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\widehat{\text{BLP}}-\text{BLP}^{*})^{2}]}_{G}+2\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]}_{H}.$
The remainder of the proof shows that $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ and $H=0$. Before proceeding, recall that $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$, and note that
$\Sigma^{*-1}\Delta=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}]^{\top}$ (18)
obtains from elaborating on (17).
Step 1.
The decomposition exercise in Section 4.3 showed that
$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}+\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*}).$
Since $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ by Theorem 2, $\widehat{\Sigma}^{-1}-\Sigma^{*-1}=-\Sigma^{*-1}\Delta\Delta^{\top}\Sigma^{*-1}/(1+\Delta^{\top}\Sigma^{*-1}\Delta)$ by the Sherman-Morrison formula. Thus
$\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*})=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\Delta^{\top}\Sigma^{*-1}(X-b^{*}),$
where the equality uses the fact that $\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta=\overline{\Delta}$. Simplifying $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ accordingly yields
$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}(1+\Delta^{\top}\Sigma^{*-1}(X-b^{*}))$
and, hence,
$G=\left(\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\right)^{2}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[(\Delta^{\top}\Sigma^{*-1}(X-b^{*}))^{2}+1+2\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right].$
The third term in the expectation operator has zero mean under the true signal distribution. The first term equals
$\mathbb{E}_{(b^{*},\Sigma^{*})}\left(\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right)^{2}$
[…]
$=\left(\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\right)^{2}\mathbb{E}[\omega^{2}]+\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})^{2}\mathbb{E}[\varepsilon_{i}^{2}]$
[…]
$=\Delta^{\top}\Sigma^{*-1}\Delta,$ ($\because\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$)
and inserting this back into the expression for $G$ yields $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$.
Step 2.
Results so far imply that
$\omega-\text{BLP}^{*}(X)$ $=\omega-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=\omega-\mathbf{1}_{N}^{\top}\left(\text{diag}(\nu)-\frac{\nu\nu^{\top}}{1+\mathbf{1}_{N}^{\top}\nu}\right)(\omega\mathbf{1}_{N}+\varepsilon)$
$=\frac{\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}}{1+\mathbf{1}_{N}^{\top}\nu},$
and that
$\Delta^{\top}\Sigma^{*-1}(X-b^{*})$ $=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}](\omega\mathbf{1}_{N}+\varepsilon)$
$=\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}.$
Simplifying $H$ accordingly yields
$H$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]$
$=\frac{-\overline{\Delta}}{(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(1+\Delta^{\top}\Sigma^{*-1}(X-b^{*}))]$
$=\frac{-\overline{\Delta}}{\left(1+\mathbf{1}_{N}^{\top}\nu\right)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\left(\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}\right)\left(1+\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}\right)\right]$
$=\frac{-\overline{\Delta}}{(1+\mathbf{1}_{N}^{\top}\nu)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\left(\overline{\Delta}\mathbb{E}[\omega^{2}]-\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})\mathbb{E}[\varepsilon_{i}^{2}]\right)$
[…]
as desired. ∎
Proof of Proposition 5.
The solution to the portfolio choice problem presented in Section 5.2 is
$\alpha^{*}_{j}=\frac{\eta_{j}}{\lambda}\left(\mu^{*}_{j}-\frac{\sum_{k=1}^{K}\mu^{*}_{k}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\right)\text{ }\forall j=1,\cdots,K$
under the correctly specified model, and
$\widehat{\alpha}_{j}=\frac{\eta_{j}}{\lambda}\left(\widehat{\mu}_{j}-\frac{\sum_{k=1}^{K}\widehat{\mu}_{k}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\right)\text{ }\forall j=1,\cdots,K$
under the misspecified model. Simplifying their difference using $\widehat{\mu}_{1}-\mu_{1}^{*}=-\overline{\Delta}$ and $\widehat{\mu}_{j}=\mu_{j}^{*}$ $\forall j\neq 1$ yields
$\Delta_{\alpha_{1}}\coloneqq\widehat{\alpha}_{1}-\alpha_{1}^{*}=-\frac{\eta_{1}}{\lambda}\left(1-\frac{\eta_{1}}{\sum_{k=1}^{K}\eta_{k}}\right)\overline{\Delta}\text{ and }\Delta_{\alpha_{j}}\coloneqq\widehat{\alpha}_{j}-\alpha_{j}^{*}=\frac{\eta_{1}\eta_{j}}{\lambda\sum_{k=1}^{K}\eta_{k}}\overline{\Delta}\text{ }\forall j\neq 1.$
The distortion in the DM’s objective expected utility thus equals
$\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})\mu_{k}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})^{2}/\eta_{k}-\sum_{k=1}^{K}\alpha^{*}_{k}\mu_{\alpha_{k}}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}\left(\alpha_{k}^{*}\right)^{2}/\eta_{k}$
$=\sum_{k=1}^{K}\Delta_{\alpha_{k}}(\mu_{k}^{*}-\lambda\alpha_{k}^{*}/\eta_{k})-\frac{\lambda}{2}\sum_{k=1}^{K}(\Delta_{\alpha_{k}})^{2}/\eta_{k}$
$=0-\frac{\lambda}{2}\sum_{k=1}^{K}(\Delta_{\alpha_{k}})^{2}/\eta_{k}=-\frac{\left(\overline{\Delta}\right)^{2}}{2\lambda}\eta_{1}\left(1-\frac{\eta_{1}}{\sum_{k=1}^{K}\eta_{k}}\right),$
where the second equality uses the facts that
$\mu_{j}^{*}-\lambda\alpha_{j}^{*}/\eta_{j}=\frac{\sum_{k=1}^{K}\mu_{k}^{*}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\text{ }\forall j=1,\cdots,K$
and $\sum_{k=1}^{K}\Delta_{\alpha_{k}}=0$. The third equality follows from straightforward algebra.∎

### v3

@@
-Learning Source Biases: Multi-sourced Misspecifications and Consequences
+Learning Source Biases: Multisource Misspecifications and Their Impact on Predictions
 Abstract
-We study how a decision maker (DM) learns about the biases of unfamiliar information sources.
-Absent any friction, a rational DM uses familiar sources as yardsticks to discern the true bias of a source.
-If the DM has misspecified beliefs, this process fails.
-We derive long-run beliefs, behavior, welfare, and corresponding comparative statics when the DM holds dogmatic, incorrect beliefs about the biases of several familiar sources.
-The distortion due to misspecified learning aggregates misspecifications associated with familiar sources independently of sources the DM learns about.
+We study how a Bayesian decision maker (DM) learns about the biases of novel information sources to predict a random state.
+Absent frictions, the DM uses familiar sources as yardsticks to accurately discern the biases of novel sources.
+We derive the distortion of the DM’s long-run prediction when the DM holds misspecified beliefs about the biases of several familiar sources.
+The distortion aggregates misspecifications across familiar sources independently of the number and nature of the novel sources the DM learns about.
 This has implications for labor market discrimination, media bias, and project finance and oversight.
 1 Introduction
-Good decision making in today’s information landscape involves the utilization of often biased information from a variety of sources.
-Some sources, such as traditional data or trusted advisors, are familiar to the decision maker (DM), and so he believes that he understands their biases well.
+Making good predictions in today’s information landscape requires a decision maker (DM) to synthesize information from a variety of biased sources.
+Some sources, such as traditional data or trusted advisors, are familiar to the DM, and so he believes that he understands their biases well.
 Still others, such as algorithm recommendations and robo-advisors, are novel and unfamiliar, and the DM is much less certain about their biases.
-In an ideal setting, an astute decision maker might be able to learn about the bias inherent in each information source over time, using familiar sources to gauge unfamiliar ones, and his beliefs would eventually converge to the truth.
+In an ideal setting, an astute DM might be able to learn about the bias inherent in each information source over time — using familiar sources to gauge unfamiliar ones — and his long-run prediction will be accurate.
 However, a serious problem arises when the DM holds dogmatic, wrong, beliefs about the biases inherent in familiar sources, due to factors that elude detection and corrections.
 The long-run implications of this error are the subject of our paper.
-Following the paradigm of misspecified Bayesian learning pioneered by Berk (1966), we model a DM who consults several information sources to predict an evolving state of the world.
+Following the paradigm of misspecified Bayesian learning, we model a DM who consults several information sources to predict an evolving state of the world.
 Initially, the DM holds different prior beliefs about the biases of familiar and unfamiliar sources.
-For familiar sources, the prior is precise but can be inaccurate (hereafter misspecified), in that it concentrates on an event in which the perceived bias of the source differs from the truth.
-For unfamiliar sources, the prior is uncertain and, crucially, contains truth in its support.
-Over time, the DM observes signal realizations and updates his belief about source biases in a Bayesian manner.
-We examine how learning in the presence of multiple, misspecified sources breeds persistent misperceptions about unfamiliar sources and fuels distortions of the DM’s long-run behavior.
-Our framework can potentially be applied to a broad set of important economic phenomena.
+For familiar sources, the prior is precise but can be inaccurate (hereinafter misspecified), in that it concentrates on an event in which the perceived bias of the source differs from the truth.
+For unfamiliar sources, the prior is uncertain and, crucially, contains the truth in its support.
+Over time, the DM observes signal realizations and updates his beliefs about source biases in a Bayesian manner.
+We examine how learning in the presence of multiple, misspecified sources breeds persistent misperceptions about unfamiliar sources and distorts the DM’s long-run prediction.
+Our framework can be applied to a broad set of important economic phenomena.
 To illustrate, consider an upper-level manager who inquires about the productivity of minority employees from a mix of familiar sources — his personal source, senior raters, standard HR algorithms — and unfamiliar sources — novel employee data and newly installed committee members.
-Although these sources are prone to gender and racial biases, in a frictionless world, biases may be inconsequential, at least in the long run.
-For example, if a novel source gives consistently lower employee ratings than a known, impartial source, then it is regarded as adversarial to minorities, and its bias is accounted for when information from diverse sources is aggregated to predict minority employees’ productivity.
-Reality, however, is not frictionless, particularly when the manager’s beliefs about familiar sources are misspecified.
-As someone who exhibits an implicit bias against minorities, the manager may give unduly low ratings to minority employees automatically and unconsciously.
-[^4] Implicit bias is an unconscious negative attitude towards a particular social group, making it challenging to eliminate due to its automatic nature.
-The renowned Implicit Association Test (IAT) developed by Greenwald et al.
-(1998) quickly assesses individuals’ associations between social groups and stereotypes, often in milliseconds.
-The opinions he gleans from other evaluators can be tainted by prejudice, stereotypes, and intergroup preferences, while his success in detecting these biases is limited, particularly in large organizations or when transparency and accountability are lacking (Roberson et al., 2007).
-Finally, algorithmic biases may arise due to poor designs or incomplete training, but they remained unnoticed until they garnered recent attention from academia and the popular press (Dastin, 2022).
-Since all these errors influence the manager’s learning about unfamiliar sources and, ultimately, the perceived productivity of minority employees and their compensation, their theoretical and practical consequences warrant a comprehensive investigation.
-Multi-sourced misspecified learning could be challenging to study for several reasons.
+Although these sources are prone to gender and racial biases, such biases may be inconsequential in the long run absent learning frictions: if a novel source consistently gives lower ratings to minorities than a known, impartial source, then its negative bias can be accounted for when information is aggregated to predict minority employees’ productivity.
+In reality, however, the manager may hold misspecified beliefs about familiar sources, due to his implicit bias against minorities (Greenwald et al., 1998), or his ignorance of the prejudice, stereotypes, and intergroup preferences inherent in the evaluations of other raters or algorithms (Roberson et al., 2007; Dastin, 2022).
+Since these unconscious biases can result in persistent underprediction and underpayment of minority employees, their theoretical and practical implications warrant a comprehensive investigation.
+Our analysis is conducted in a Gaussian environment in which the signal generated by a source in a given period equals the bias of the source, plus the true state of the world in that period and an error.
+Errors are independent of the true state but can exhibit arbitrary correlations among themselves.
+Studying multi-sourced misspecified learning in this environment could be challenging.
 In our leading example, suppose that the manager perceives two sources as impartial, although in reality one is biased in favor of minorities and the other is biased against them.
 Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences.
-One potential approach is to find a middle ground and use this as a neutral point to form beliefs about other sources.
-Still, the question remains: How exactly can this be achieved and how does the answer depend on the correlations between the signals?
-Correlation affects how different signals are weighed in the learning process.
-With multi-sourced misspecifications, the need to weigh different sources promotes the role of correlation in shaping learning outcomes, and the compound effect can escape even the most sophisticated analysis.
-We develop an approach that works for arbitrary numbers of misspecified sources.
-Our analysis is conducted in a Gaussian environment in which the signal generated by a source in a given period equals the bias of the source, plus the true state of the world in that period and a Gaussian error.
-Errors are independent of the true state and of each other in the baseline model, but can take more general forms in later sections.
-Our main result, presented in Section 3.2, captures the distortions of the DM’s long-run perceptions and behavior by a single, intuitive, metric.
-In the baseline model, this metric equals the total misspecification associated with familiar sources, weighted by their relative precision to each other.
-As we exacerbate the misspecification associated with a source or increase its precision relative to others, its impact on the metric increases.
-In the leading example, if a majority of familiar sources, especially those exhibiting high signal-to-noise ratios, are biased against minorities but are nevertheless perceived as impartial, then the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities and is mirrored in his beliefs about unfamiliar sources through misspecified learning.
-In the long run, the manager underestimates the bias of any adversarial source toward minorities while overestimating that of any source favorable to them.
-He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by our metric.
-Our approach utilizes existing mathematical results on low-rank updates of invertible matrices.
-In the baseline model, signals (net of their biases) equal the true state plus idiosyncratic errors, so their covariance matrix is a rank-one update of a diagonal matrix.
-When signals are correlated through a single common channel, the impact of correlation on the outcomes of misspecified learning is limited.
-Consequently, the metric that captures the distortions of the DM’s long-run belief and behavior exhibits a striking feature that we call DOOM, as it depends only on sources with misspecified biases but not those that the DM learns about.
-DOOM continues to hold in variants of the baseline model in which the signals depend on a common latent factor, or when the true state and the signals can take high-dimensional values (see Sections 4.1 and 4.2).
-When correlations between signals become arbitrary, the DM’s misperceptions about source biases satisfy a weaker form of DOOM, as they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else.
-Further, the distortion in the DM’s long-run behavior always satisfies DOOM, due to the simple fact that the DM’s limiting belief concentrates on the minimizer of the Kullback-Leibler divergence from truth.
-The comparative statics of our metric have important implications for reality.
-In the context of labor market discrimination, DOOM implies that the damage caused by a few rotten apples cannot be rectified by installing objectively neutral but lesser-known members of the minority group to the evaluation panel, or by training junior raters to reduce errors in employee assessments.
-The same cannot be said about sources that the DM holds dogmatic, wrong, beliefs about, as our metric is in general nonmonotonic in their signal-to-noise ratios.
+One potential approach is to aggregate the differences between these sources and use the outcome as a middle ground to form beliefs about other sources and the true state.
+Still, the question remains: How exactly can this be achieved in general environments where familiar sources, unfamiliar sources, and the true state may exhibit arbitrary correlations?
+Our main result, presented in Section 3.2, describes how the misspecifications associated with familiar sources are aggregated into distortions of the DM’s long-run beliefs about source biases and his prediction of the true state.
+Strikingly, the distortion in his prediction does not depend on unfamiliar sources at all, but instead depends only on misspecified sources, a property we call DOOM.
+The proof of DOOM exploits fundamental properties of the statistical decision problem faced by the DM, which has two stages.
+In the first stage, the DM learns the biases of unfamiliar sources using misspecified, familiar sources as yardsticks, which amounts to minimizing the Kullback-Leibler (KL) divergence from the truth.
+In linear-Gaussian environments, this is equivalent to forming the constrained least squares estimator of the biases of unfamiliar sources (at population level).
+In the second stage, the DM uses the first-stage result to form the best linear predictor of the true state.
+Since the former is the best constrained linear estimator of source biases, distortions thereof — which have already been optimized — should not further distort the second-stage predictor, regardless of the number and nature of the unfamiliar sources being learned.
+We can draw sharper results by imposing additional structures on the covariance matrix.
+In Section 3.3, we examine a simple case in which the errors associated with the signals are jointly independent.
+Leveraging matrix algebra theories on rank-one updates of matrices, we develop a single metric that fully captures the distortionary effects of misspecified learning.
+The metric equals the total misspecification associated with familiar sources, weighted by their relative precision.
+In our leading example, it takes the form an unconscious negative bias against minorities and is mirrored in the manager’s beliefs about unfamiliar sources through misspecified learning.
+In the long run, the DM underestimates the bias of any unfamiliar source that is adversarial toward minorities while overestimating the bias of any unfamiliar source that is favorable to them.
+Consequently, he underpredicts minority employees’ productivity and pays them less than under the correctly specified model.
+The comparative statics of our distortionary metric have important implications for reality.
+In the context of labor market discrimination, DOOM implies that the damage caused by a few rotten apples cannot be rectified by installing objectively neutral but lesser-known members of the minority group to the evaluation panel, or by training junior raters to reduce rating errors.
+The same cannot be said about sources that the DM holds dogmatic, wrong, beliefs about, as our metric is nonmonotonic in their signal-to-noise ratios in general.
 In the presence of multiple misspecified sources, improving the precision of one source renders the others relatively imprecise.
-The overall impact on our metric is in general ambiguous, and this has implications for the use of rater training to curtail discrimination.
-Details are in Section 3.5.
-While it is tempting to conjecture that correcting the misspecifications associated with familiar sources may improve learning outcomes, reality is more subtle when the covariance matrix of the signals is also subject to misspecified learning.
-In Section 4.3, we show that learning the covariance matrix of the signals, in addition to their biases, makes the signals appear noisier to the DM than they truly are.
-Consequently, the DM relies less on signal realizations when predicting the true state of the world, and this countervailing force is shown to partially correct the distortions due to misspecified learning.
-Policies that operate by correcting misspecifications block this novel channel and so could entail an ambiguous consequence on learning outcomes.
-This finding speaks to the debiasing programs used by real-world organizations to address discrimination, such as implicit bias training, priming shared identities, and promoting algorithmic fairness.
-It suggests that even if these programs are effective, their behavioral and welfare consequences may remain ambiguous.
-In addition to employee evaluation and compensation, the aforementioned phenomenon occurs when voters form perceptions of the biases inherent in new social media sites based on their erroneous beliefs about traditional media.
-There is evidence that partisans tend to view party-congenial sources as neutral, while displaying excessive hostility toward out-party sources (Pennycook and Rand, 2021; Hassell et al., 2020).
-Likewise, a manager seeking project advice may allow his hardwired misperceptions about familiar sources — such as overoptimism, confirmation bias, and overtrust in one’s inner circle (Malmendier, 2018) — to shape his views about unfamiliar sources such as external consultants and robo-advisors.
-In Section 5, we demonstrate how these ingrained biases can spill into new domains, generating the hostile media effect and false polarization, [^5] The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is.
-False polarization measures the difference between the perceived distance between two groups and the actual distance between them.
-Both phenomena have attracted considerable scholarly attention within the domains of political science, communication, and psychology.
-Section 5.1 offers a synthesis of the literature.
-in addition to polarizing voters’ behavior and distorting managers’ portfolio choices.
-We further contend, based on our comparative statics, that corrective measures that target political misinformation or managerial biases may give rise to unintended consequences.
-For example, while fact-checking new media always improves voters’ welfare because of DOOM, fact-checking tradition media generally has an ambiguous impact due to the distortions that arise from misspecified learning.
-The latter are nonmonotonic in the accuracy of familiar sources and the treatment effects of misperception recalibration.
+The overall impact on our metric is generally ambiguous.
+While it is tempting to conjecture that correcting the misspecifications associated with familiar sources may improve predictions, reality is more subtle when the covariance matrix of the signals is also subject to misspecified learning.
+In Section 5.1, we show that learning the covariance matrix of the signals, in addition to their biases, makes the signals appear noisier to the DM without further distorting his beliefs about unfamiliar sources compared to the baseline model.
+Consequently, the DM relies less on signal realizations than before when predicting the true state, and this countervailing force is shown to partially mitigate the distortions caused by misspecified learning.
+Policies aimed at correcting misspecifications may undermine this bias mitigation effect, resulting in an exacerbation of distortions.
+This finding speaks to the debiasing programs used by real-world organizations to curb implicit discrimination (e.g., implicit bias training, efforts to reduce intergroup animosity and to promote algorithmic fairness), suggesting that even if these programs are effective, their behavioral and welfare consequences may remain ambiguous.
+Our model can additionally be applied to study misperceived media bias and project management under misspecified learning.
+We summarize our findings — particularly the implications of DOOM for polarization, fact-checking and the efficacy of corporate repair policies — in Section 6 and defer the analysis to the Online Appendix.
+The desirability of DOOM, including its use to identify misspecified sources, its relevance to alternative decision situations, and its robustness to alternative forms of misspecifications and high-dimensional states and signals, are explored in Sections 4 and 5.2.
 1.1 Related literature
 The study of misspecified Bayesian learning models has a long tradition in statistical decision theory.
-The canonical paradigm endows the DM with a prior belief over the models that describe the underlying environment.
-If the prior does not contain the true model in its support, it is misspecified.
-The seminal work of Berk (1966) provides conditions under which the DM’s limiting posterior belief concentrates on the set of models that minimize the Kullback-Leibler (KL) divergence from truth.
-Economic models of misspecified Bayesian learning can be broadly categorized into passive learning models and active learning models.
-The former, including ours, assume that the actions taken by the DM have no impact on the signals generated by truth.
-The work most related to ours is carried out by Heidhues et al.
-(2019) (HKS).
-These authors consider environments where the signals generated by unknown fundamentals follow multivariate Guassian distributions.
-Such environments are shown to satisfy Berk’s conditions for the convergence of the limiting belief to KL minimizers.
-Our proof of the convergence of the limiting belief builds on this result.
-The main differences between our results and those of HKS lie in the characterization of the limiting belief.
-In particular, HKS focus on situations where a single dimension of the fundamentals is misspecified, which is the natural setting to understand the implications of overconfidence.
-We entertain the possibility of multidimensional misspecifications, which is essential to understand how a DM aggregates biased sources in the settings discussed previously.
-Our characterization extends beyond that of HKS and works for arbitrary numbers of misspecified sources with general covariance matrices.
-The structural insights it uncovers, such as DOOM, have escaped the attention of previous work.
-Our comparative statics are sharpest when the covariance matrix of the signals constitutes low-rank updates of diagonal matrices.
-HKS consider alternative forms of covariance matrices that are suitable for their exploration of the consequences of overconfidence on intergroup prejudice.
-In contrast, models of misspecified active learning allow the DM’s actions to affect the signals that he subsequently uses to form beliefs about the truth.
-The literature on misspecified active learning has recently advanced significantly.
-Among others, Esponda and Pouzo (2016) propose Berk-Nash equilibrium as a steady-state notion for misspecified learning dynamics.
-Esponda et al.
-(2021), Fudenberg et al.
-(2021), Bohren and Hauser (2021), and Frick et al.
+The seminal work of Berk (1966) provides conditions under which the limiting posterior beliefs concentrate on the model that minimizes the Kullback-Leibler divergence from the truth.
+Economic models of misspecified Bayesian learning can be broadly categorized into passive and active learning models.
+The former, including ours, assume that the actions taken by the DM do not affect the signals generated by the underlying truth.
+Our contributions to passive misspecified learning are twofold.
+First, we propose a model of making predictions using familiar and unfamiliar information sources applicable to broad phenomena.
+Second, we obtain general characterizations of the long-run outcomes using fundamental principles from statistical decision making.
+The most related work to ours is Heidhues et al.
+(2019) (hereinafter, HKS), which examines a passive learning model with a single source of misspecification and demonstrates that Berk’s convergence result applies in multidimensional Gaussian environments.
+Our convergence proof builds on their analysis.
+The main difference between HKS and the current work lies in the DM’s decision problem.
+HKS examine how the DM’s overconfidence in his caliber distorts his estimates about ingroups and outgroups — a problem akin to what our DM faces in the first stage.
+However, our DM uses multiple misspecified sources to guide his learning of unfamiliar sources and, crucially, uses the learned model to predict a random state.
+This second stage, absent from HKS, generates new insights such as DOOM that have eluded previous researchers.
+Applications of misspecified passive learning models to other areas of economics include Bohren et al.
+(2019) and Molavi et al.
+(2023).
+Bohren et al.
+(2019) examine, theoretically and experimentally, a model in which a DM holds erroneous prior beliefs about minorities and updates his beliefs over time based on new information.
+The authors characterize the trajectory of beliefs and distinguish their result from traditional taste-based or rational statistical discrimination.
+Our model addresses labor market discrimination in the long run and its relationship between traditional and novel information sources.
+Molavi et al.
+(2023) examine a model in which asset returns follow a high-dimensional hidden factor model, while market participants are restricted to using lower-dimensional models to explain observed data.
+Their misspecification differs from ours in nature, and they explore its consequences for asset pricing and the dynamics of forecast errors.
+Models of misspecified active learning allow the DM’s actions to affect the signals used to form beliefs about the truth.
+The theory of misspecified active learning has recently advanced significantly, with a primary focus on the convergence properties of various solution concepts.
+[^5] Among others, Esponda and Pouzo (2016) propose Berk-Nash equilibrium as a steady state notion for misspecified learning dynamics.
+Fudenberg et al.
+(2021) and Frick et al.
 (2023) examine the convergence of Berk-Nash equilibria in single and multi-agent environments.
 Ba (2021) entertains the possibility of regime shifts.
-He and Libgober (2020) examine the evolutionary foundations of perpetual model misspecification.
-None of these papers shares our focus on the characterization of steady states with multi-sourced misspecifications.
-There is a broader conceptual connection between our work and the Bayesian network approach of Spiegler (2020) to model misspecified causality, in that both allow for misspecifications of high-dimensional objects.
-Yet Spiegler’s use of directed acyclic graphs to model causality has no obvious counterpart in our setting, and his solution concept requires that the DM take endogenous actions to fully explain the observed data.
-Our steady state does not involve endogenous actions and has a misspecified-Bayesian-learning foundation.
-Developments on the theoretical frontier have spurred novel research in various areas, including labor market discrimination, media bias, and finance.
-Within the realm of labor market discrimination, Bohren et al.
-(2019) and Bohren et al.
-(2023) propose methods to distinguish discrimination due to erroneous beliefs from those rooted in tastes or rational beliefs.
-The current paper adds to this agenda by exploring a novel channel: misspecified learning of unfamiliar information sources.
-The main takeaways from our analysis — namely dogmatic, wrong, beliefs about familiar sources can breed persistent misperceptions about unfamiliar sources and generate nuanced comparative statics when subject to informational interventions — have no counterparts in existing theories of discrimination.
-To the extent that interventions such as improved source accuracy or misperception recalibration are also proposed by Bohren et al.
-(2023) to identify and address discrimination, our analysis could impact how policies are designed to combat discrimination in the real world.
-The economic literature on media bias is vast, but is devoted to understanding actual media biases (Anderson et al., eds, 2016).666(Mis)perceived media bias is an important topic in political science, communication, and related fields (Feldman, 2014).
-These strands of literature are discussed in Section 5.1.
-A notable exception is made by Gentzkow et al.
-(2018) (GWZ), whose model of misperceived media bias features a payoff-irrelevant confounder that is independent of the payoff-relevant state.
-A source is biased if it exhibits a nonzero correlation with the confounder.
-The DM misperceives his personal source as unbiased and uses this belief to guide his consumption of other media sources.
-The model of GWZ differs from ours in its construct and predictions, in addition to requiring different methodologies to analyze.
-First, GWZ examine an active learning model with a single misspecified source.
-Second, their way of modeling biases (as correlations with a random confounder) implies that the distortion in the DM’s long-run behavior has a zero unconditional mean, which is not the case here.
-A central focus of GWZ is on the endogenous trust in news sources, defined as the latter’s (mis)perceived correlations with the true state.
-Trust is correctly specified throughout this paper and is known to the DM for the most part.
-When endogenized as a subject of misspecified learning in Section 4.3, we find no effect on perceived news biases and only a moderation of the DM’s behavior.
-The literature on behavioral finance is vast (see Malmendier 2018 for a survey), although research that explores misspecified Bayesian learning as a source of friction remains sparse.
-Molavi et al.
-(2023) examine a passive learning model in which asset returns follow a high-dimensional hidden factor model, although the models used by market participants to explain observed data have a lower dimensionality.
-The authors examine the consequences of this departure from the rational expectations paradigm for asset pricing and the dynamics of forecast errors.
-The Sherman-Morrison formula — a tool to invert rank-one perturbations of invertible matrices — is present in their mathematical derivations (and those of Levy et al.
-2022) but plays a less significant role than in our analysis.
+He and Libgober (2020) examine the evolutionary foundation of perpetual model misspecification.
+There is also a broader conceptual connection between our work and the Bayesian network approach of Spiegler (2020) to model misspecified causality.
+Both allow for misspecifications of high-dimensional objects, but Spiegler’s use of directed acyclic graphs to model causality has no obvious counterpart in our setting, and his solution concept requires that the DM take endogenous actions to fully explain the observed data.
 2 Baseline model
 Time is discrete and infinite.
-In each period $t=1,2,\cdots$, a random state $\omega_{t}$ is drawn independently from the standard normal distribution.
-A decision maker (DM) — informed by a finite set $\mathcal{S}\coloneqq\{1,\cdots,N\}$ of sources — predicts the mean of $\omega_{t}$.
-The signal generated by source $i\in\mathcal{S}$ in period $t$ is $X_{i,t}=\omega_{t}+b_{i}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}$ is the bias of the source, and $\varepsilon_{i,t}$ is a Gaussian error with mean zero and variance $v_{i}>0$.
-In the baseline model, errors are assumed to be independent of the true state, across each other, and over time.
-Initially, the DM holds a prior belief about source biases.
-In every subsequent period, he updates his belief in a Bayesian manner, based on the prior and signal realizations.
-It is well known that if true biases, which we denote by $b_{i}^{*}$s, lie in the support of the prior, then the DM’s posterior belief will eventually converge to the truth.
-His long-run behavior is the best linear predictor of the state given signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$, under the (correct) belief that source biases have value $b^{*}\coloneqq[b_{1}^{*},\text{ }\cdots\text{ },b_{N}^{*}]^{\top}$.
+In each period $t\in\mathbb{N}$, a random state $\omega_{t}$ is drawn independently from the standard normal distribution.
+A DM — informed by a finite set $\mathcal{S}\coloneqq\{1,\cdots,N\}$ of sources — predicts the mean of $\omega_{t}$.
+The signal generated by source $i\in\mathcal{S}$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\epsilon_{i,t}$, where $b_{i}\in\mathbb{R}$ is the bias of the source, and $\epsilon_{i,t}$ is a state-independent Gaussian error with mean zero and variance $v_{i}>0$.
+Initially, the DM holds prior beliefs about source biases.
+In every subsequent period, he updates his beliefs in a Bayesian manner based on the prior and signal realizations.
+It is well known that if true biases, which we denote by $b_{i}^{*}$, $i\in\mathcal{S}$, lie in the support of the prior, then the DM’s posterior beliefs will eventually concentrate on the truth.
+His long-run behavior is the best linear predictor of the state given signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$, under the correct belief that source biases have value $b^{*}\coloneqq[b_{1}^{*},\text{ }\cdots\text{ },b_{N}^{*}]^{\top}$.
 We examine the case where the DM’s prior is partly misspecified.
 In the baseline model, we focus on a particular kind of misspecification, whereby the DM holds dogmatic, wrong, beliefs about the biases of some sources and uses them to guide his learning about other sources.
-We examine how such misspecifications distort the DM’s learning and, ultimately, his long-run belief and behavior.
+We examine how such misspecifications distort the DM’s learning and, ultimately, his long-run beliefs and prediction.
 Formally, we say that the DM misspecifies the bias of source $i$ if he initially assigns probability one to an event in which $b_{i}=\tilde{b}_{i}$, for some $\tilde{b}_{i}\neq b_{i}^{*}$.
@@
 The support of the prior is $\Theta\coloneqq\{b\in\mathbb{R}^{N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.
-The outcomes of Bayesian learning based on this prior are the subject of the current study.
-Three remarks are in order.
-For now, we assume that signals equal biases plus the true state and independent errors, and that their covariance matrix is known to the DM from the outset.
-In Sections 3.4 and 4.3, we entertain more general forms of signals whose covariance matrix may also be a subject of misspecified learning.
-Second, we treat misspecification and degenerate, wrong, prior as synonyms for most of the paper; alternative forms of misspecifications are discussed in Section 4.1.
-Finally, it will become clear that adding sources with correctly specified biases to the analysis would not affect our predictions in any meaningful way.
+The long-run outcomes of learning based on this prior are the subject of the current study.
+Several remarks are in order.
+First, the assumption of unit covariance between the true state and the signals is made to simplify the exposition.
+The case of general, known, covariances is discussed in Section 3.2.
+Second, we treat misspecification and degenerate, wrong, prior as synonyms for most of the paper; alternative forms of misspecifications are discussed in Section 4.
+Third, we assume, for now, that the true covariance matrix $\Omega^{*}$ of the errors — which can be any symmetric and positive definite matrix in $\mathbb{R}^{N\times N}$ — is known to the DM.
+In Section 5.1, we consider the possibility that the covariance matrix may also be a subject of misspecified learning.
+Finally, it will become evident that adding sources with correctly specified biases to the analysis would not affect our predictions in any meaningful way.
 The example presented below serves to illustrate our framework and results throughout the paper.
@@
 To proceed with our analysis, we first present some preliminary concepts and results.
-These are followed by a formal statement of the main results, then a discussion of their intuitions, proofs, and policy implications.
+These are followed by a formal statement of the main results, then a discussion of their intuitions, proofs, and implications.
 3.1 Preliminaries
-We first formalize the problem faced by the DM.
+Throughout the paper, we follow the convention of using $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$.
+For each $k\in\mathbb{N}$, we use $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, and $\mathbb{I}_{k\times k}$ to denote the $k\times k$ identity matrix.
+We first formalize the DM’s problem.
 Let $\mathcal{N}(b,\Sigma)$ denote an arbitrary $N$-dimensional normal distribution with mean $b\in\mathbb{R}^{N}$ and covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$, and let $\mathcal{N}(b^{*},\Sigma^{*})$ denote the true distribution of the signals.
@@
 $D_{KL}(b^{*},\Sigma^{*}\|b,\Sigma)=\frac{1}{2}\left(\text{tr}(\Sigma^{-1}\Sigma^{*})-N+(b-b^{*})^{\top}\Sigma^{-1}(b-b^{*})+\log\frac{\text{det}\Sigma}{\text{det}\Sigma^{*}}\right).$ (1)
-The next lemma — which adapts the seminal result of Berk (1966) and that of HKS to the current setting — shows that the DM’s long-run belief concentrates on the set of distributions that minimize the Kullback-Leibler divergence from the true distribution (hereafter KL minimizers).
+The next lemma adapts the seminal result of Berk (1966) and, more recently, Theorem A.1 Case (I) of HKS, to the current setting, showing that the DM’s long-run beliefs concentrate on the set of distributions that minimize the Kullback-Leibler divergence from the true distribution (hereinafter, KL minimizers).
 That is, for every open cover of the set of KL minimizers, the DM assigns probability one to it in the limit, as $t\rightarrow+\infty$.
 Lemma 1.
-[…]
-Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ be a typical KL minimizer.
+The DM’s long-run beliefs concentrate on the set of distributions that minimize the KL divergence (1), subject to the constraint that $b\in\Theta$ and $\Sigma=\Sigma^{*}$.
+Proof.
+Omitted proofs from the main text are relegated to the appendix. ∎
+Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ denote a typical KL minimizer.
 Since $\Sigma^{*}$ is known to the DM from the outset, $\widehat{\Sigma}=\Sigma^{*}$.
@@
 $\min_{b\in\Theta}\frac{1}{2}(b-b^{*})^{\top}\Sigma^{*-1}(b-b^{*}).$ (2)
-In Section 3.3, we show that the solution to (2) is unique and fully captured by $\Delta\coloneqq\hat{b}-b^{*}$.
-Call $\Delta$ the DM’s misperception about source biases in the long run, and note that $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}$ $\forall i\in\mathcal{M}$.
-Let $\text{BLP}^{*}(X)$ and $\widehat{\text{BLP}}(X)$ denote the best linear predictors of the state given signals, under the belief that source biases have values $b^{*}$ and $\hat{b}$, respectively.
-These are the long-run behaviors of the DM under the correctly specified model and the misspecified model, respectively, and their difference captures the distortionary effect of misspecified learning on the DM’s long-run behavior.
-3.2 Results
-Our main results are twofold.
-The first characterizes the distortionary effects of misspecified learning on the DM’s long-run belief and behavior.
-The second investigates comparative statics of distortionary effects.
-For starters, write $\nu_{i}$ for the signal-to-noise ratio $1/v_{i}$ of source $i$. For each $i\in\mathcal{M}$, define
-$\gamma_{i}\coloneqq\frac{\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$
-as the relative precision of the signal generated by source $i$ compared to other misspecified sources. Define
-$\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\gamma_{i}\Delta_{i}$
-as the total misspecification across the sources in $\mathcal{M}$, weighted by their relative precision to each other. Then
+In the next section, we show that the solution to (2) is unique and fully captured by $\Delta\coloneqq\hat{b}-b^{*}$.
+Call $\Delta$ the DM’s misperception about source biases in the long run, and note that $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\neq 0$ for all $i\in\mathcal{M}$.
+Let $\text{BLP}^{*}(X)$ and $\widehat{\text{BLP}}(X)$ denote the best linear predictors of the state given signals, under the beliefs that source biases have values $b^{*}$ and $\hat{b}$, respectively.
+These are the long-run predictions made by the DM under the correctly specified and misspecified models, and their difference captures the distortionary effect of misspecified learning on the long-run prediction.
+In the leading example, $\Delta$ represents the DM’s misperception about the biases of the various sources evaluating minority employees.
+To understand the difference between $\widehat{\text{BLP}}(X)$ and $\text{BLP}^{*}(X)$, suppose there are additional, unmodeled sources that provide information about majority employees, but the DM doesn’t hold any unconscious bias against them.
+In the limit, the DM’s model about majority employees is accurate, and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ represents the persistent wage gap between minority and majority employees, holding the profile of evaluations fixed.
+3.2 General case
+This section characterizes the distortionary effects of misspecified learning in general environments.
+To begin, write
+$A\coloneqq\begin{bmatrix}\Sigma^{*}_{1,1}&\cdots&\Sigma^{*}_{1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{M,1}&\cdots&\Sigma^{*}_{M,M}\end{bmatrix}\text{ and }C\coloneqq\begin{bmatrix}\Sigma^{*}_{M+1,1}&\cdots&\Sigma^{*}_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{N,1}&\cdots&\Sigma^{*}_{N,M}\end{bmatrix}$
+for the covariance matrix between familiar sources and its counterpart between familiar and unfamiliar sources, respectively.
+Denote the vector $[\Delta_{1},\cdots,\Delta_{M}]^{\top}$ of misspecifications associated with familiar sources by $\widetilde{\Delta}$, and the vector $[\Delta_{M+1},\cdots,\Delta_{N}]^{\top}$ of misperceptions about unfamiliar sources by $\widehat{\Delta}$.
 Theorem 1.
-  1. (i)
-    $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$; $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
-  2. (ii)
-    $\overline{\Delta}$ depends only on $(\Delta_{i},\nu_{i})_{i\in\mathcal{M}}$.
-For each $i\in\mathcal{M}$, $\partial\overline{\Delta}/\partial\Delta_{i}>0$, while $\partial\overline{\Delta}/\partial\nu_{i}$ generally has an ambiguous sign.
-Part (i) of Theorem 1 reduces the distortionary effects of misspecified learning to a single metric: $\overline{\Delta}$.
-To develop intuitions, suppose that the manager in our leading example begins with an erroneous belief that two sources are both impartial, although in reality one is biased against minorities and the other is in favor of them, that is, $b_{1}^{*}<0<b_{2}^{*}$ and $\tilde{b}_{1}=\tilde{b}_{2}=0$.
-Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences.
-One potential approach is to find a middle ground, that is, a position in $(b_{1}^{*},b_{2}^{*})$, and use it as a neutral point to form beliefs about other sources.
-The more precise source $i$ is relative to source $j$, the closer the neutral point is to $b_{i}^{*}$.
-In the case where $\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}<0$, the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities of magnitude $\overline{\Delta}=|\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}|$, and it is mirrored in his beliefs about unfamiliar sources through misspecified learning.
-In the long run, the DM underestimates the bias of any source that is adversarial toward minorities while overestimating that of any source that is favorable to them, assuming that $\overline{\Delta}$ is not excessively large.
-He also underestimates minority employees’ productivity and pays them less than under the correctly specified model by $\overline{\Delta}$.
-Part (ii) of Theorem 1 investigates the comparative statics of our metric.
-The result $\partial\overline{\Delta}/\partial\Delta_{i}>0$ is the most intuitive.
-In the leading example, this means that correcting the unconscious bias associated with any familiar source reduces its impact on the metric.
-The comparative statics regarding source precision are more nuanced.
-As we increase the absolute precision of source $i$ (as captured by $\nu_{i}$), we also increase its relative precision (as measured by $\gamma_{i}$), while decreasing the relative precision $\gamma_{j}$ of any other source $j\in\mathcal{M}-\{i\}$.
-As a consequence, the DM takes source $i$ more seriously in the learning process, and misspecifications of its bias distort the DM’s long-run belief and behavior more heavily than before.
-The opposite happens to source $j$.
-Depending on the exact misspecifications associated with each individual source, the outcome is either an exacerbation of the distortionary effect, or a mitigation thereof.
-Such an ambiguity hinges on the multi-sourced nature of misspecifications, i.e., $M\geq 2$.
-If instead $M=1$ (as in HKS), then raising $\nu_{i}$ unambiguously increases the magnitude of $\overline{\Delta}$, due to the first effect described above.
-Perhaps the most striking feature of $\overline{\Delta}$ is that it depends only on misspecified sources (DOOM) but not those with initially unknown biases.
-In general, DOOM fails because the correlations between the signals affect how the latter are weighed in the learning process.
-The presence of multi-sourced misspecifications can significantly complicate the analysis, as the need to weigh different sources promotes the role of correlation in shaping learning outcomes.
-The compound effect may escape even the most sophisticated analysis.
-When correlations are concentrated on only a few channels (as in the baseline model), their impacts on learning outcomes are limited.
-The proof presented in the next section formalizes these ideas.
-3.3 Proof of Theorem 1
-Throughout this paper, we use $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ to denote the square matrix created from a vector, and $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$.
-Write the covariance matrix $\Sigma^{*}$ of the signals as
-$\left[\begin{array}[]{c|c}A_{M\times M}&B_{M\times(N-M)}\\ \hline\cr C_{(N-M)\times M}&D_{(N-M)\times(N-M)}\end{array}\right],$
-where
-$A\coloneqq\begin{bmatrix}\Sigma^{*}_{1,1}&\cdots&\Sigma^{*}_{1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{M,1}&\cdots&\Sigma^{*}_{M,M}\end{bmatrix}\text{ and }C\coloneqq\begin{bmatrix}\Sigma^{*}_{M+1,1}&\cdots&\Sigma^{*}_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Sigma^{*}_{N,1}&\cdots&\Sigma^{*}_{N,M}\end{bmatrix}$
-capture the correlations between familiar sources and those between familiar and unfamiliar sources, respectively.
-By Lemma 1, the DM’s long-run belief is a solution to (2). Rewrite (2) as
+For any symmetric and positive definite matrix $\Sigma^{*}\in\mathbb{R}^{N\times N}$, $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ and $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$.
+Theorem 1 prescribes a two-step procedure for aggregating the misspecifications associated with familiar sources into distortions of the DM’s long-run beliefs and prediction.
+The first step weighs the misspecifications associated with familiar sources ($\widetilde{\Delta}$) by their covariances ($A$).
+The second step converts these weighted sums into the DM’s misperceptions about unfamiliar sources and the distortion in his prediction of the state, using their respective covariances between familiar sources ($C$ and $\mathbf{1}_{M}$).
+Notably, the distortion of the prediction depends only on misspecified sources, but not sources with initially unfamiliar biases or the opportunity to learn about their biases over time, a property we call DOOM.
+Misperceptions about unfamiliar sources satisfy a weaker form of DOOM, or WDOOM, in that they depend on the covariances between these sources and familiar ones, but not on the covariances between unfamiliar sources.
+We examine the implications of (W)DOOM in Section 4.
+In Section 1, we discussed the potential challenges associated with studying multi-sourced misspecified learning with general covariance structures.
+Given this, the results demonstrated by Theorem 1 are notably sharp.
+What is more remarkable is the simplicity of their proof, which exploits the fundamental connection between learning source biases and making corresponding predictions.
+We present the proof below.
+Proof of Theorem 1.
+Our DM moves in two stages.
+In the first stage, he forms beliefs about the biases of unfamiliar sources, which amounts to solving the KL minimization problem.
+In the current linear-Gaussian environment with a known covariance matrix, the KL minimizer is nothing but the constrained least squares estimator of source biases (at population level).
+[^7] To establish the connection formally, consider, for each $t$, the vector expression $X=b^{*}+(\omega\mathbf{1}_{N}+\epsilon).$ In Section 3.1 we have defined the variance matrix $\Sigma^{*}:=Var[\omega\mathbf{1}_{N}+\epsilon]$.
+Since $\Sigma^{*}$ is positive definite, define $\epsilon_{u}:={\Sigma^{*}}^{-1/2}[\omega\mathbf{1}_{N}+\epsilon]$, and note that $Var[\epsilon_{u}]\equiv\mathbb{E}[\epsilon_{u}\epsilon_{u}^{\top}]=Identity.$ So the above vector expression becomes ${\Sigma^{*}}^{-1/2}X_{t}={\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut},$ where $t$ indices the $t$th observation.
+Define the constrained least-square estimator for $b^{*}$ as $\mbox{argmin}_{b\in\Theta}\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right).$ The objective function is equivalent to $\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}X_{t}-{\Sigma^{*}}^{-1/2}b\right)$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\left({\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut}-{\Sigma^{*}}^{-1/2}b\right)^{\top}\left({\Sigma^{*}}^{-1/2}b^{*}+\epsilon_{ut}-{\Sigma^{*}}^{-1/2}b\right)$ $=$ $\frac{1}{T}\sum_{t=1}^{T}\left[\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1}\left(b^{*}-b\right)+2\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1/2}\epsilon_{ut}+\epsilon_{ut}^{\top}\epsilon_{ut}\right].$ To solve the minimization problem at the population level, either take $T\to\infty$, or take the expectation at $T=1$.
+Either way the result is the same, that the population level objective function is just $\left(b^{*}-b\right)^{\top}{\Sigma^{*}}^{-1}\left(b^{*}-b\right).$ Minimizing the population level objective function is simply (2).
+To solve, we rewrite the KL minimization problem as
 $\min_{\Delta\in\mathbb{R}^{N}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i=1,\cdots,M}\frac{1}{2}\Delta^{\top}\Sigma^{*-1}\Delta,$
 and note that its solution is fully determined by the system of first-order conditions: $\forall j=M+1,\cdots,N$,
-$\sum_{i=1}^{N}\Sigma^{*-1}_{ji}\Delta_{i}=0.$ (3)
-We proceed in two steps.
-Step 1.
-Consider first the case where $\Sigma^{*}$ is an arbitrary symmetric positive definite matrix.
-The next lemma solves for (3) in this general case.
-For ease of notation, write $\widetilde{\Delta}$ for the vector $[\Delta_{1},\cdots,\Delta_{M}]^{\top}$ of misspecifications associated with familiar sources, and $\widehat{\Delta}$ for the vector $[\Delta_{M+1},\cdots,\Delta_{N}]^{\top}$ of misperceptions about unfamiliar sources.
-Lemma 2.
-For any symmetric positive definite matrix $\Sigma^{*}\in\mathbb{R}^{N\times N}$, the solution to (3) is $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ and in particular satisfies $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$.
-The proof of Lemma 2 exploits only the first-order conditions associated with KL minimization, together with the block matrix inversion formula.
-The result prescribes a two-step procedure for aggregating the misspecifications associated with familiar sources into distortions of the DM’s long-run belief and behavior.
-The first step weighs misspecifications using the correlations between familiar sources (as captured by $A^{-1}\widetilde{\Delta}$).
-The second step transforms these weighted sums into the DM’s misperceptions about unfamiliar sources, using the correlations between familiar and unfamiliar sources as captured by $C$.
-Interestingly, correlations between unfamiliar sources do not enter this weighting process.
-Furthermore, the distortion of the DM’s long-run behavior depends only on misspecified sources but not those he learns about, thus always satisfying DOOM.
-Step 2.
-While the DM’s misperceptions about source biases do not generally satisfy DOOM due to the complications that arise in the weighting process (as captured by $CA^{-1}$), they become relatively straightforward when the correlations between the signals are concentrated on only a few channels.
-In the environment considered in the baseline model, distortions of the DM’s long-run perceptions and behavior collapse into a single metric: $\overline{\Delta}$.
-A seemingly innocuous assumption turns out to be key, namely signals net of their biases equal the true state plus idiocyncratic errors.
-As a result, their covariance matrix $\Sigma^{*}$ is the sum of a diagonal matrix: $\text{diag}(v)$, and a matrix of rank one: $\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
-The next lemma, proven by Sherman and Morrison (1950), prescribes a formula to invert rank-one updates of invertible matrices.
-Lemma 3 (Sherman and Morrison 1950).
+$\sum_{i=1}^{N}\Sigma^{*-1}_{ji}\Delta_{i}=0.$
+Write the covariance matrix $\Sigma^{*}$ as $\footnotesize\left[\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right]$, where $A$ and $C$ are defined above.
+Applying the block matrix inverse formula to $\Sigma^{*}$ yields
+$\Sigma^{*-1}=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right],$
+hence the system of first-order conditions can be rewritten as
+$-(D-CA^{-1}B)^{-1}CA^{-1}\widetilde{\Delta}+(D-CA^{-1}B)^{-1}\widehat{\Delta}=\mathbf{0}_{N-M}.$
+Multiplying both sides of the last equation by $D-CA^{-1}B$, we obtain $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ as the DM’s misperceptions about the biases of unfamiliar sources.
+In the second stage, the DM uses the result obtained in the first stage to form the best linear predictor of the true state.
+Since the former is the best constrained linear estimator of source biases, distortions thereof — which have already been optimized in the first stage — should not further distort the prediction in the second stage.
+Formally, by substituting $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ into the block covariance matrix and simplifying, we obtain
+$\Sigma^{*-1}\Delta$ $=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right]\left[\begin{array}[]{c}\widetilde{\Delta}\\ \hline\cr\widehat{\Delta}\end{array}\right]$ (7)
+$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}+A^{-1}B(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\\ \hline\cr-(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\end{array}\right]$ (9)
+$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right].$ (12)
+Consequently,
+$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta},$
+where the last term depends only on misspecified sources and so satisfies DOOM. ∎
+Two remarks are in order.
+First, we assume unit covariance between the true state and the signals to simplify exposition.
+In general, the distortion of BLP equals $\alpha_{M}^{\top}A^{-1}\widetilde{\Delta}$, where $\alpha_{M}$ is the vector of covariances between misspecified sources and the true state.
+Second, in the above derivation, the linearity of both the first-stage estimator and the second-stage predictor is key, while the normality assumption is only used to establish the equivalence between minimizing KL and constrained least squares.
+In Section 4, we explore the consequence of this observation for alternative decision situations.
+3.3 The case of independent errors
+This section delves deeper into the roles of familiar sources in the formation of distortions by examining a simple environment in which the errors associated with the signals are jointly independent.
+Many of our key economic and methodological insights can be effectively illustrated within this simple framework.
+To begin, write $\nu_{i}$ for the signal-to-noise ratio $1/v_{i}$, or absolute precision, of source $i$. For each $i\in\mathcal{M}$, define
+$\gamma_{i}\coloneqq\frac{\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$
+as the relative precision of source $i$ compared to other misspecified sources. The term
+$\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\gamma_{i}\Delta_{i}$
+captures the total misspecification across the sources in $\mathcal{M}$, weighted by their relative precision.
+The next proposition shows that $\overline{\Delta}$ fully captures the distortionary effects of misspecified learning in the current context.
+Proposition 1.
+When errors are jointly independent:
+  1. (i)
+    The DM’s misperception about the bias of any unfamiliar source $i\in\mathcal{S}-\mathcal{M}$ is given by $\Delta_{i}=\overline{\Delta}$, and the distortion of his long-run prediction is given by $\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
+  2. (ii)
+    For each $i\in\mathcal{M}$, $\partial\overline{\Delta}/\partial\Delta_{i}=\gamma_{i}>0$ and $\partial\overline{\Delta}/\partial\nu_{i}=(\Delta_{i}-\overline{\Delta})/(1+\sum_{j\in\mathcal{M}}\nu_{j})$, which is positive if and only if $\Delta_{i}>\overline{\Delta}$.
+To develop intuitions, suppose that the manager in our leading example begins with an erroneous belief that two familiar sources are both impartial, although in reality one is biased against minorities and the other is in favor of them, that is, $b_{1}^{*}<0<b_{2}^{*}$ and $\tilde{b}_{1}=\tilde{b}_{2}=0$.
+Over time, the manager observes the signals generated by these sources but cannot fully reconcile their differences.
+There is a third, unfamiliar, source, which is impartial in reality.
+To gauge its position, the manager uses familiar sources as yardsticks.
+If source one is used as the yardstick, the third source appears favorable to minorities.
+If, instead, source two is used as the yardstick, then the third source seems hostile.
+One potential approach is to find a middle ground, that is, a position between $b_{1}^{*}$ and $b_{2}^{*}$, and use it as a neutral point to gauge the position of the unfamiliar source.
+The more biased source $i$ is than source $j$ or the more precise it becomes relative to source $j$, the closer the neutral point is to $b_{i}^{*}$.
+These are the exact properties satisfied by our distortionary metric $\overline{\Delta}$.
+In the case where $\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}<0$, the overall distortion of the manager’s worldview constitutes an unconscious negative bias against minorities of magnitude $\overline{\Delta}=|\gamma_{1}b_{1}^{*}+\gamma_{2}b_{2}^{*}|$, and it is mirrored in his beliefs about unfamiliar sources through misspecified learning.
+In the long run, the DM underestimates the bias of any unfamiliar source that is adversarial toward minorities while overestimating the bias of any unfamiliar source that is favorable to them, assuming that $\overline{\Delta}$ is not excessively large.
+He also underpredicts minority employees’ productivity and pays them less than under the correctly specified model by $\overline{\Delta}$.
+Part (ii) of Proposition 1 explores the comparative statics of $\overline{\Delta}$.
+The result $\partial\overline{\Delta}/\partial\Delta_{i}>0$ $\forall i\in\mathcal{M}$ is the most intuitive.
+It simply says that correcting the unconscious bias associated with any familiar source (bringing $\Delta_{i}$ closer to zero) reduces its impact on the overall distortionary effect.
+This change mitigates the overall distortion if $\overline{\Delta}$ and $\Delta_{i}$ have the same sign; otherwise it exacerbates the overall distortion.
+The comparative statics regarding source precision are more nuanced.
+As we increase the absolute precision of one source, we also increase its relative precision while decreasing the relative precision of the other source.
+As a consequence, the DM takes the first source more seriously in the learning process, and misspecifications of its bias distort the DM’s long-run beliefs and behavior more heavily than before.
+The opposite happens to the second source.
+Depending on the exact nature of misspecifications, the outcome is either an exacerbation of the distortionary effect, or a mitigation thereof.
+In the example described above, $\Delta_{1}>\overline{\Delta}>0>\Delta_{2}$.
+Consequently, raising $\nu_{1}$ by a small amount exacerbates the distortionary effect, while raising $\nu_{2}$ by a small amount does the opposite.
+Note that this ambiguity hinges on the multiplicity of misspecified sources.
+If, instead, only source one is misspecified, then $\Delta_{1}>\overline{\Delta}=\nu_{1}\Delta_{1}/(1+\nu_{1})$.
+Raising $\nu_{1}$ by a small amount unambiguously exacerbates the distortion of the DM’s long-run beliefs and prediction in this case.
+Proof of Proposition 1.
+The proof of Proposition 1, like many to come, exploits a simple observation: in our context, the covariance matrix $\Sigma^{*}$ of the signals is the sum of the covariance matrix $\Omega^{*}$ of the errors and a rank-one update $\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ that reflects the common state of the world.
+This model feature allows us to utilize matrix algebra results concerning rank-one updates of matrices — such as the Sherman-Morrison formula — to characterize the long-run outcome.
+Fact 1 (Sherman and Morrison 1950).
 Let $G\in\mathbb{R}^{N\times N}$ be an invertible matrix and $x,y\in\mathbb{R}^{N}$ be column vectors. Then $G+xy^{\top}$ is invertible if and only if $1+y^{\top}G^{-1}x\neq 0$, in which case
 $(G+xy^{\top})^{-1}=G^{-1}-\frac{G^{-1}xy^{\top}G^{-1}}{1+y^{\top}G^{-1}x}.$
-Applying the Sherman-Morrison formula to $A=\text{diag}(v_{1},\cdots,v_{M})+\mathbf{1}_{M}\mathbf{1}_{M}^{\top}$ yields
-$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1}\\ \vdots\\ \nu_{M}\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}$
-and, hence,
-$\mathbf{1}_{M}^{\top}A^{-1}=\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
-Substituting these results, together with $C=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}$, into Lemma 2 yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\overline{\Delta}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=-\overline{\Delta}$ as desired.
-3.4 Discussions
-When distortions of the DM’s long-run perceptions violate DOOM, they still satisfy a weaker form of DOOM (hereafter WDOOM), in that they depend only on misspecified sources and their correlations with unfamiliar sources, but nothing else.
-In the simple environment presented below, the comparative statics of the distortions are analogous to those described in Theorem 1.
-Proposition 1.
-In the case where $X_{i,t}=b_{i}+\alpha_{i}\omega_{t}+\varepsilon_{i,t}$, $\Delta_{i}=\alpha_{i}\overline{\Delta}$ for all $i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i\in\mathcal{M}}\alpha_{i}^{2}\nu_{i}).$
-Proposition 1 shows that when signals are correlated with the true state to varying degrees, the long-run misperceptions held by the DM will vary between sources, albeit only through $\alpha_{i}$.
-In addition to comparative statics with respect to $\Delta_{i}$ and $\nu_{i}$, a new insight emerges from the analysis: In the case where $\alpha_{i}>0$ $\forall i\in\mathcal{S}$ (if not, let $-X_{i}$ be the subject of analysis), increasing $\alpha_{i}$, $i\in\mathcal{S}-\mathcal{M}$ makes source $i$ more informative and therefore more susceptible to the distortionary effect of misspecified learning.
-In the long run, this worsens the DM’s misperception about source $i$ relative to other sources that he learns about.
-The aforementioned environments satisfy a common feature, namely signals depend on each other only through the true state of the world.
-In the presence of additional sources of correlation, the weighting formula prescribed by Lemma 2 becomes complex very quickly, as shown by the example presented below.
-Example 2.
-Suppose that the period-$t$ signal generated by source $i$ is $X_{i,t}=b_{i}+\omega_{t}+\beta_{i}\theta_{t}+\varepsilon_{i,t}$, where $\theta_{t}\sim\mathcal{N}(0,1)$ represents shocks to people’s attitudes toward minorities triggered by social movements such as MeToo, and is assumed to be independent of the true state and the Gaussian errors.
-The presence of $\theta_{t}$ creates an additional source of correlation between the signals.
-The resulting covariance matrix $\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}+\beta\beta^{\top}$ is a rank-two update of a diagonal matrix, so its inverse can be obtained by iteratively applying the Sherman-Morrison formula first to $\widetilde{\Sigma}\coloneqq\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ and then $\Sigma^{*}=\widetilde{\Sigma}+\beta\beta^{\top}$.
-The algebra is brutal in general. In the case where $N=5$ and $M=3$, it can be shown that
-$\begin{bmatrix}\Delta_{4}\\ \Delta_{5}\end{bmatrix}=\begin{bmatrix}\frac{P\beta_{4}+Q}{R}\\ \frac{P\beta_{5}+Q}{R}\end{bmatrix}$
-where
-$P=\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
-$Q=\left(\sum_{i=1}^{3}\nu_{i}\Delta_{i}\right)\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\Delta_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right),$
-and
-$R=\left(1+\sum_{i=1}^{3}\beta_{i}^{2}\nu_{i}\right)\left(1+\sum_{i=1}^{3}\nu_{i}\right)-\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right)\left(\sum_{i=1}^{3}\beta_{i}\nu_{i}\right).$
-$\Delta_{4}$ and $\Delta_{5}$ depend on the correlations between familiar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{M}$) and those between familiar and unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i\in\mathcal{M}$ and $j\in\mathcal{S}-\mathcal{M}$).
-They are independent of the correlations between unfamiliar sources (as captured by $\beta_{i}\beta_{j}$, $i,j\in\mathcal{S}-\mathcal{M}$), thus satisfying WDOOM.
-3.5 Implications
-We have discussed the intuitions and mathematical underpinnings of our results in previous sections.
-We now examine their implications for the various policy interventions used by real-world organizations to address discrimination.
+When errors are jointly independent, their covariance matrix is simply the diagonal matrix of their variances, i.e., $\Omega^{*}=\text{diag}(v)$.
+Applying the Sherman-Morrison formula to the covariance matrix among familiar sources: $A=\text{diag}(v_{1},\cdots,v_{M})+\mathbf{1}_{M}\mathbf{1}_{M}^{\top}$, yields
+$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\nu_{i}}\begin{bmatrix}\nu_{1}\\ \vdots\\ \nu_{M}\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
+Substituting this finding, along with $C=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}$ as the covariance matrix between familiar and unfamiliar sources, into Theorem 1, we obtain
+$\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\frac{\sum_{i=1}^{M}\nu_{i}\Delta_{i}}{1+\sum_{i=1}^{M}\nu_{i}}\coloneqq\overline{\Delta}$
+as the distortion of the DM’s prediction, and
+$CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\overline{\Delta}$
+as his misperceptions about unfamiliar sources. ∎
+4 Implications
+This section examines consequences of our results for various policy interventions aimed to curb discrimination, as well as their broader theoretical implications.
 Minority representation at upper echelons.
-There is a large economic literature that studies the antidiscrimination effect of increasing minority representations at upper echelons.
-According to Bertrand and Duflo (2017), a simple mechanical mechanism may nevertheless prove to be game changing.
-Historically, individuals from underrepresented groups have had limited influence in decision making.
+There is a large economic literature that studies the anti-discrimination effect of increasing minority representations at upper echelons.
+According to Bertrand and Duflo (2017), a simple mechanical mechanism may nevertheless prove to be game changing: Historically, individuals from underrepresented groups have had limited influence in decision making.
 To the extent that majority decision makers are biased against minority subordinates, discrimination tends to persist.
@@
 Our result offers a gloomier prospect, suggesting that the damage caused by existing members of the evaluation panel may not be easily rectified by infusing new perspectives.
-Indeed, DOOM stipulates that the introduction of fresh, lesser-known committee members from the minority group does not alter the metric that captures the distortionary effect of misspecified learning.
-Diversity efforts at the top may work through alternative channels, but not through expanding the set $\mathcal{S}-\mathcal{M}$.
+Indeed, DOOM stipulates that minority employees are underpaid to the same degree, with or without the introduction of fresh, lesser-known committee members from the minority group to the evaluation panel.
+Diversity efforts at the top may work through alternative channels, but not through expanding the set of lesser-known sources.
 Reduce decision errors.
-Contemporaneous rater training programs offered by numerous real-world organizations emphasize an important lesson: When making decisions that can adversely affect others (e.g., performance reviews), stakeholders should “slow down, meditate, deliberate, and follow elaborate structured procedures” so as to base decisions on facts rather than instincts (Eberhardt, 2020).
+Contemporaneous rater training programs offered by numerous real-world organizations to curb discrimination emphasize an important lesson: When making decisions that can adversely affect others (e.g., performance reviews), stakeholders should “slow down, meditate, deliberate, and follow elaborate structured procedures” so as to base decisions on facts rather than instincts (Eberhardt, 2020).
 Recently, Kleinberg et al.
 (2018) demonstrate the potential of using properly trained AI algorithms to overcome human decision errors in the context of bail decision making.
-To the extent that the aforementioned interventions may increase the signal-to-noise ratios of our sources, their impacts on the outcomes of misspecified learning can be quite nuanced.
-For one thing, changing the signal-to-noise ratios of unfamiliar sources is ineffectual.
-For another, scaling the signal-to-noise ratios of all familiar sources by the same factor $\alpha>1$ increases the DM’s reliance on them when learning about the biases of unfamiliar sources, thereby amplifying the distortions in his long-run perception and behavior:
-$\left|\frac{\sum_{i\in\mathcal{M}}\alpha\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\alpha\nu_{i}}\right|=\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{\alpha^{-1}+\sum_{i\in\mathcal{M}}\nu_{i}}\right|>\left|\frac{\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}\right|=|\overline{\Delta}|.$
-When changes occur disproportionately at different sources, the outcomes of learning may be improved due to the ambiguity of the sign of $\partial\overline{\Delta}/\partial\nu_{i}$.
-Yet for this approach to succeed, precise information about the misspecifications associated with each source is needed, and this knowledge is unlikely to be possessed by the DM.
+To the extent that the aforementioned interventions may increase the precision of our sources, their impacts on the outcomes of misspecified learning can be quite nuanced.
+For one thing, changing the signal-to-noise ratios of unfamiliar sources is ineffectual because of DOOM.
+For another, the success of changing the signal-to-noise ratios of familiar sources hinges on the nature of misspecifications.
+In the case of independent errors, perturbing the signal-to-noise ratios of familiar sources by $(d\nu_{i})_{i\in\mathcal{M}}$ changes $\overline{\Delta}$ by
+$\frac{\sum_{i\in\mathcal{M}}d\nu_{i}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}\left(\sum_{i\in\mathcal{M}}\Delta_{i}\kappa_{i}-\overline{\Delta}\right),\text{ where }\kappa_{i}\coloneqq\frac{d\nu_{i}}{\sum_{j\in\mathcal{M}}d\nu_{j}}.$
+The impact depends not only on the total perturbation $\sum_{i\in\mathcal{M}}d\nu_{i}$, but also on the relative perturbations $\kappa_{i}$ of the various sources, weighted by their respective misspecifications $\Delta_{i}$.
+For this approach to succeed, one needs precise information about the misspecifications to make astute choices about $\kappa_{i}$.
+[^8] In particular, naively scaling the signal-to-noise ratios of all misspecified sources by the same factor $\alpha>1$ results in $d\nu_{i}=(\alpha-1)\nu_{i}$ and $\kappa_{i}=\nu_{i}/\sum_{j\in\mathcal{M}}\nu_{j}$.
+The change in $\overline{\Delta}$ equals $\frac{(\alpha-1)\overline{\Delta}}{1+\sum_{i\in\mathcal{M}}\nu_{i}}$, hence the distortionary effect of misspecified learning is amplified by the intervention.
+In practice, this knowledge is unlikely to be possessed by the DM.
 Debiasing programs.
-Certainly, the aforementioned interventions may operate through an alternative channel, that of correcting the misspecifications associated with familiar sources (bring $\Delta_{i}$s closer to zero).
-Indeed, the idea of using one’s own mental resources to debias decisions is backed by years of sociopsychological research into implicit biases, and recent evidence suggests combining it with the use of structured evaluations to prohibit decision making based on automatic, unconscious instincts (Bohnet, 2016; Greenwald and Lai, 2020).888 Outside the corporate realm, the Oakland Policy Department recently adjusted its foot pursuit policy so that officers were instructed to “step back, slow down, call for backup, and think it through” rather than to follow suspects into backyards or blind alleys.
-Nextdoor Neighbor, a social network for residential neighbors to communicate through, recently started asking its users to provide detailed descriptions of the suspicious activities they wish to report, because “adding frictions allows users to act based on information rather than instinct” (Eberhardt, 2020).
+Certainly, the aforementioned interventions may operate through an alternative channel, that of (partially) correcting the misspecifications associated with familiar sources.
+Indeed, the idea of using one’s own mental resources to debias decisions is backed by years of sociopsychological research into automatic, unconscious biases (Greenwald and Lai, 2020).999 Outside the corporate realm, the Oakland Police Department recently adjusted its foot pursuit policy so that officers were instructed to “step back, slow down, call for backup, and think it through,” rather than follow suspects into backyards or blind alleys.
+Nextdoor, a social network for residential neighbors to communicate through, recently started asking its users to provide detailed descriptions of the suspicious activities they wish to report, because “adding frictions allows users to act based on information rather than instinct” (Eberhardt, 2020).
 Meanwhile, Devine et al.
-(2012) provide the first evidence that prejudice, stereotypes, and intergroup hostility are “bad habits” that can be broken through longitudinal interventions such as counterstereotypical thinking, priming shared identities, perspective taking, individulating, and intergroup contact.
+(2012) provide the first evidence that prejudice, stereotypes, and intergroup hostility are “bad habits” that can be broken through longitudinal interventions such as counterstereotypical thinking, priming shared identities, and intergroup contact.
 Finally, the growing literature on algorithmic fairness proposes various remedies to remove biases from algorithm design and training (Mitchell et al., 2021).
-In addition to interventions that can help bring $b_{i}^{*}$s closer to zero, Kalev et al.
-(2006) and Raji et al.
-(2020) stress the importance of accountability and transparency.
-The latter raise the DM’s awareness and consciousness and can potentially bring perceived biases ($\tilde{b}_{i}$s) closer to true biases ($b_{i}^{*}$s).
+While the above interventions aim to bring true source biases $b_{i}^{*}$ closer to zero, another approach is to align perceived biases $\tilde{b}_{i}$ with true biases.
+This can be achieved by enhancing accountability and transparency within the organization (Kalev et al., 2006; Raji et al., 2020), therefore revealing the biases of lower-ranked raters to the DM.
+Alternatively, the DM can consult third parties who, without any preconception, estimate the biases of familiar and unfamiliar sources simultaneously.
+[^10] This is the essence of the famous Durbin-Wu-Hausman test that econometricians use to detect model misspecification (Patrick, 2020).
+These efforts can raise the DM’s awareness of potential misspecification, (hopefully) prompting a revision of perceptions about familiar sources.
 The efficacy of the aforementioned programs in addressing discrimination is a subject of active research and hot debate (more on this in Section 5.1).
-In case they are effective in bringing $\Delta_{i}$s closer to zero, their impacts on the DM’s belief and behavior are straightforward in the context of the baseline model.
-Outside the baseline model, things can become muddier, a subject to which we turn in Section 4.3.
+In case they are effective in bringing misspecifications closer to zero, their impacts on the DM’s beliefs and prediction are straightforward in the context of the baseline model.
+Outside the baseline model, things can become muddier, as we shall demonstrate in Section 5.1.
+Alternative decision situations.
+Consider a common scenario in finance, consulting, and service industries, where a DM informs his client of the true state $\omega$ using a linear model $X=b+\omega\mathbf{1}_{N}+\epsilon$.
+Suppose $\omega$ and $\epsilon$ have mean zero but may not follow normal distributions, and $\omega$ has unit variance.
+When estimating his model, the DM fixes the mean of $X_{i}$, $i\in\mathcal{M}$, to $\tilde{b}_{i}$, consistent with industry standards or assumptions he deems appropriate.
+In case these assumptions deviate from reality, the DM’s estimator — which minimizes the estimation squared-error — becomes the constrained least squares estimator defined in Footnote 7.
+Based on this, the DM makes a linear prediction to minimize the prediction squared-error.
+Formally, a linear predictor for $\omega$ in $X$ has the form $\beta^{\top}(X-b)$.
+The best linear predictor, BLP, is the one that minimizes prediction squared-error $\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$.
+Proposition 2.
+In situations described above, the distortion in BLP is as in Theorem 1 and in particular satisfies DOOM.
+Identify misspecified sources.
+DOOM can be used to identify misspecified sources.
+In the context of the leading example, suppose that we, as researchers, can alter the covariance matrix of the signals through informational interventions, such as rater training, as discussed in Section 3.3.111111Similar types of informational interventions have been proposed by Bohren et al.
+(2023) as means to distinguish discrimination based on erroneous beliefs from traditional taste-based or rational statistical discrimination.
+However, the issue of misspecified learning of the biases of novel sources and its relationship between the persistent wage gap between majorities and minorities has eluded previous researchers.
+DOOM stipulates that the wage gap between majority employees — who are exempted from unconscious biases — and minority employees — who are subject to unconscious biases — should vary only with the covariances between misspecified sources, but not the covariances between novel sources or those between novel and misspecified sources.
+Alternative forms of misspecification.
+We model misspecification as degenerate, wrong, prior beliefs about familiar sources, consistent with the misspecified Bayesian learning literature.
+However, the qualitative nature of our results remains valid if the support of the prior is an unbounded convex polyhedron in $\mathbb{R}^{N}$.
+[^12] The assumption of unboundedness ensures that the limiting posterior beliefs concentrate on the set of KL minimizers.
+An example would be $\Theta=\times_{i=1}^{N}[\underline{b}_{i},\overline{b}_{i}]$, where $[\underline{b}_{i},\overline{b}_{i}]$ is unbounded for some $i$.
+Under this more general assumption, the KL minimization problem remains a convex optimization problem, solvable using the Lagrangian method.
+Let $\mathcal{M}$ denote the dimensions involved in binding linear constraints and call $\mathcal{M}$ the endogenous set of familiar sources.
+Given $\mathcal{M}$, all previous results, including DOOM, remain valid.
 5 Extensions
-This section presents several extensions of the baseline model.
-We first allow the signals to depend on a common latent factor and the true state and the signals to take high-dimensional values.
-We then examine a problem in which both the biases of the signals and their covariance matrix are the subject of misspecified learning.
-4.1 Latent factor
-This section examines a variant of the baseline model in which the means of the signals can depend on a common latent factor $b_{N+1}\in\mathbb{R}$, such as employees’ productivity-enhancing investments.
-Redefine $b\coloneqq[b_{1},\cdots,b_{N+1}]^{\top}$ and call it fundamental.
-The mean vector of the signals $X\coloneqq[X_{1},\cdots,X_{N}]^{\top}$ is $\Gamma b$, where $\Gamma\coloneqq[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$ is the transformation matrix that operates by adding the latent factor $b_{N+1}$ to the mean of each signal.
-Initially, the true value of the latent factor $b_{N+1}$, together with the biases of the sources in $\mathcal{S}-\mathcal{M}$, is unknown to the DM.
-Everything else is as in the baseline model.
-The support of DM’s prior becomes $\Theta=\{b\in\mathbb{R}^{N+1}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.
-[…]
-$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
-Proposition 2.
-Let everything be as above.
-Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\Delta_{N+1}=-\overline{\Delta}$, where $\overline{\Delta}\coloneqq\sum_{i\in\mathcal{M}}\nu_{i}\Delta_{i}/\sum_{i\in\mathcal{M}}\nu_{i}.$
-Proposition 2 shows that the presence of a common latent factor has little qualitative impact on the metric that captures the distortionary effects of misspecifed learning.
-The latter still equals the total misspecification associated with the sources in $\mathcal{M}$, weighted by their relative precision.
-Its comparative statics are the same as in the baseline model.
-In the leading example where the distortion of the manager’s worldview constitutes an unconscious negative bias against minorities, $\overline{\Delta}>0$.
-Proposition 2 predicts that in the long run, the DM undervalues minority workers’ productivity-enhancing investment by $\overline{\Delta}$, in addition to viewing sources favorable to them as too lenient and those adversarial toward them as impartial.
-Such a misperception clearly has impact on minority workers’ incentive to undertake costly investments, a subject we plan to explore in depth in a follow-up paper.
-The proof of Proposition 2 shows that the transformed matrix $\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma$ is a rank-one update of a diagonal matrix, and so the KL minimization problem can be solved using an analogous approach to that developed in Section 3.3.
-Although it is challenging to exhaust all transformations of the DM’s problem that preserve the structural insights of the current analysis, it is certainly possible to consider specific transformations, such as one where the support of the DM’s prior (equivalently, the choice set in the KL minimization problem) is of the form ×$\times_{i\in\mathcal{S}}[\underline{b}_{i},\overline{b}_{i}]$.
-The baseline model is a special case where $\underline{b}_{i}=\overline{b}_{i}=\tilde{b}_{i}$ for $i\in\mathcal{M}$, whereas $\underline{b}_{i}=-\infty$ and $\overline{b}_{i}=+\infty$ for $i\in\mathcal{S}-\mathcal{M}$.
-It is easy to show, using the Lagrangian method, that this model variation has no qualitative impact on the distortionary effect of misspecified learning.
-All it does is endogenizing $\mathcal{M}$ as the set of sources whose interval constraints are binding; given $\mathcal{M}$, everything said so far remains valid.
-4.2 Multidimensional state and signals
-This section generalizes the baseline model to encompass multidimensional states and signals.
-In each period $t=1,2,\cdots,$ Nature draws the true state $\omega_{t}$ from a $K$-dimensional Gaussian distribution with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{0}$.
-Each dimension represents a distinct skill, such as problem solving, communication, teamwork, etc.
-The DM consults a set $\mathcal{S}=\{1,\cdots,N\}$ of information sources before making a prediction about the true state.
-The signal generated by source $i$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\varepsilon_{i,t}$, where $b_{i}\in\mathbb{R}^{K}$ is the $K$-dimensional bias of the source, and $\varepsilon_{i,t}$ is a source-specific, idiosyncratic error with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{i}$.
-The state and errors are jointly independent as in the baseline model, although different dimensions of the state or an error can exhibit correlations.
-The true bias vector is $b^{*}\coloneqq[b_{1}^{*},\cdots,b_{N}^{*}]^{\top}\in\mathbb{R}^{NK}$.
-Initially, the DM misspecifies the biases of the sources in =$\mathcal{M}=\{1,\cdots,M\}$ as $\tilde{b}_{i}\neq b_{i}^{*}$, $i\in\mathcal{M}$, while being uncertain of the biases of the remaining sources.
-The support of his prior is $\{b\in\mathbb{R}^{NK}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.
-Everything else is as in the baseline model.
-In particular, the DM knows the true covariance matrix $\Sigma^{*}$ of the signals, which in the current setting equals $\text{diag}(\Omega_{1},\cdots,\Omega_{N})+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\otimes\Omega_{0}$ ($\otimes$ denotes the Kronecker product).
-Key variables of our interest, including the distortion of DM’s long-run perceptions about source biases and that of his long-run behavior, are defined the same as before.
-The next proposition establishes the analog of Theorem 1 in the current multidimensional setting.
-Proposition 3.
-Redefine $\overline{\Delta}\coloneqq(\sum_{i=0}^{M}\Omega_{i}^{-1})^{-1}\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}.$ Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\emph{BLP}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
-The proof of Proposition 3 replaces the Sherman-Morrison formula with the Woodbury formula for inverting block matrices (Woodbury, 1950).
-Given the close resemblance between $\overline{\Delta}$ and its single-dimensional counterpart, we have no doubt that its comparative statics are generally ambiguous.
-However, we do know that $\overline{\Delta}$ continues to satisfy DOOM and so is independent of $\Omega_{i}$, $i\in\mathcal{S}-\mathcal{M}$.
-We could also gain insight into simple economic problems such as the one presented below, based on the closed-form formula for $\overline{\Delta}$.
-Example 3.
-The DM faces a stream of female employees whose analytical skills and communication skills (hereafter denoted by $A$ and $B$, respectively) are drawn independently from the standard Normal distribution, i.e., $\Omega_{0}=\mathbb{I}_{2\times 2}$.
-The DM consults two sources about employees’ skills but misspecifies the biases of source one.
-Although the latter undervalues female employees’ analytical skills and overvalues their communication skills, the DM perceives it as impartial, i.e., $\Delta^{A}>0$ and $\Delta^{B}<0$.
-What the DM does know is that source one’s assessments of different skills exhibit a halo effect (Bohnet, 2016).
-That is, a positive impression about one’s analytical skill enhances the assessment of her communication skill, and vice versa.
-Let $\Omega_{1}=\small{\begin{bmatrix}v&v\rho\\ v\rho&v\end{bmatrix}}$ be the covariance matrix of the errors associated with source one’s signal.
-We capture the halo effect by letting $\rho>0$.
-We examine how the DM’s misspecification about source one’s biases distorts his learning about source two’s biases, as well as his predictions about employees’ skills. Tedious algebra shows that
-$\begin{bmatrix}\overline{\Delta}^{A}\\ \overline{\Delta}^{B}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{(1+v)\Delta^{A}-\rho v\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{+}\\ \underbrace{\frac{-\rho v\Delta^{A}+(1+v)\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{-}\end{bmatrix}\text{ and }\begin{bmatrix}\frac{\partial\overline{\Delta}^{A}}{\partial\rho}\\ \frac{\partial\overline{\Delta}^{B}}{\partial\rho}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{2\rho v^{2}(1+v)\Delta^{A}-v(1+v(2+v+\rho^{2}v))\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{+}\\ \underbrace{\frac{-v(1+v(2+v+\rho^{2}v))\Delta^{A}+2\rho v^{2}(1+v)\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{-}\end{bmatrix}.$
-In words, the DM ends up perceiving source two as too lenient about women’s analytical skills and too critical about their communication skills.
-This misperception is exacerbated by the halo effect, whose presence requires that the DM subtract the communication score from the analytical score to tease out the fundamentals from the errors.
-It is through this channel that the DM’s unconscious, negative, bias about the communication score assigned by source one creeps into his learning about the analytical score assigned by source two, making the latter appear more positively biased than it actually is.
-In the long run, the DM underestimates female employees’ analytical skills and overestimates their communication skills.
-To the extent that these estimates are key to HR decision making, such as task assignment, hiring, and retention, they can potentially reinforce occupational segregation and stereotypes that are pervasive in the workplace (see Blau and Kahn 2017 and Bertrand 2020 literature surveys).
-It is also natural for employees who anticipate such a misperception to overinvest in communication skills and to underinvest in analytical skills, thus fueling a vicious circle (Arrow, 1971).
-Our model uncovers the channels through which misspecified learning can potentially affect these labor market outcomes.
-A comprehensive exploration of these mechanisms is better left for future research.
-While it is tempting to suggest, based on the simple model presented above, that companies focus rater training on correcting for halo effects, complications may arise in the presence of multiple misspecified sources, as correcting for the halo effect exhibited by one source makes those exhibited by other sources more salient.
-The total effect could be an exacerbation or a mitigation of the distortions, for the same reason that renders the comparative statics of the signal-to-noise ratio ambiguous in Theorem 1(ii).
+This section examines two extensions of the baseline model, one in which the biases of the signals and their covariance matrix are both the subject of misspecified learning, and another in which the true state and the signals can take high-dimensional values.
 5.1 Unknown covariance matrix
@@
 In environments such as large organizations with numerous divisions and hierarchies, a more natural assumption seems to be that the covariance matrix is also a subject of misspecified learning.
-We entertain this possibility in this section, with a focus on whether the additional source of misspecified learning leads to more biased beliefs and behavior on the part of the DM, or a mitigation thereof.
-Formally, fix the covariance matrix of Gaussian errors $\varepsilon_{1,t}\cdots,\varepsilon_{N,t}$ to any $\Omega^{*}\in\mathbb{R}^{N\times N}$ that is symmetric and positive definite (SPD).
-Fix any $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$, [^9] Throughout the paper, we use $a\wedge b$ to denote the minimum of $a$ and $b$ and $a\vee b$ to denote the maximum of $a$ and $b$.
-and define $\Lambda$ as the set of $N\times N$ SPD matrices whose minimum eigenvalues are at least $\underline{\lambda}$.
-Initially, the DM holds a prior belief about the distribution of the signals.
-The prior is correct about signals equal to their biases plus the true state and state-independent errors.
+We consider this possibility in this section, with a focus on whether the additional source of misspecified learning increases the bias in the DM’s beliefs and prediction or ameliorates his bias.
+Formally, fix the covariance matrix of the errors to any $\Omega^{*}\in\mathbb{R}^{N\times N}$ that is symmetric and positive definite (SPD).
+Fix any $\underline{\lambda}\in(0,\min\{1,\lambda_{\min}(\Omega^{*})\})$, and define $\Lambda$ as the set of $N\times N$ SPD matrices whose minimum eigenvalues are at least $\underline{\lambda}$.
+Initially, the DM holds prior beliefs about the distribution of the signals.
+The prior is correct about signals equalling their biases plus the true state and state-independent errors.
 However, it misspecifies the biases of some sources as in Section 2, while allowing the covariance matrix of the errors to be anything in $\Lambda$.
@@
 Importantly, $\Theta$ contains the true covariance matrix of the signals $\Sigma^{*}=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
+This means that while the DM is initially uncertain about the covariance matrix, he never holds any misspecification against it.
 We characterize outcomes of misspecified learning in two steps, beginning with the following lemma.
+[^13] In the Online Appendix, we provide an alternative, frequentist, foundation for these limiting beliefs.
 Lemma 2.
-[…]
-Let $(\hat{b},\widehat{\Sigma})$ denote the KL minimizer.
-Define the DM’s misperception of source biases, as well as the distortion of his long-run behavior, as before.
-The next theorem pinpoints the difference between these quantities and their analogs in the baseline model where the true covariance matrix is known to the DM from the outset.
+In the environment considered in this section, the DM’s long-run beliefs concentrate on the set of distributions that minimize the KL divergence (1), subject to the constraint that $(b,\Sigma)\in\Theta$.
+Let $\mathcal{N}(\hat{b},\widehat{\Sigma})$ denote the KL minimizer, which is shown to be unique.
+Define the DM’s misperception of source biases, as well as the distortion of his long-run prediction about the true state, as before.
+The next theorem pinpoints the difference between these quantities and their analogs in the baseline model.
 Theorem 2.
-Let everything be as above. Then
+In the environment considered in this section,
   1. (i)
-    $\Delta$ is the same as in the case where the DM knows $\Sigma^{*}$ from the outset, whereas $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$;
+    The DM’s misperception about source biases $\Delta$ is the same as in the baseline model, while his limiting belief about the covariance matrix of the signals is now given by $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$.
   2. (ii)
-    $\mathbb{E}_{(b^{*},\Sigma^{*})}[\widehat{\emph{\text{BLP}}}(X)-\emph{\text{BLP}}^{*}(X)]$ equals $\delta$ times its counterpart in the case where $\Sigma^{*}$ is known from the outset, where $\delta\coloneqq(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}\in(0,1)$.
+    The expected distortion of the DM’s prediction, evaluated under the true signal distribution, equals its counterpart in the baseline model, discounted by $\delta\coloneqq(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}\in(0,1)$.
   3. (iii)
-[…]
-Part (i) of Theorem 2, which characterizes the DM’s long-run belief, adapts Theorem A.1 Case (III) of HKS to the current setting where the covariance matrix of the signals takes the form of $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
-The remaining parts of the theorem are new.
-On the one hand, misspecified learning of the covariance matrix does not interfere with that of source biases and leaves the DM’s long-run misperceptions about the latter unaffected.
-On the other hand, misspecified learning of source biases makes the signals appear noisier to the DM than they truly are, in that the perceived covariance matrix equals its true counterpart, plus a rank-one update $\Delta\Delta^{\top}$ that is positive definite.
-[^10] The perceived noisiness of the signals as measured by, e.g., the determinant of their covariance matrix, is larger with compound misspecified learning than without, i.e., $\text{det}(\widehat{\Sigma})=\text{det}(\Sigma^{*})(1+\Delta^{\top}\Sigma^{*-1}\Delta)>\text{det}(\Sigma^{*})$, where the equality follows from Ding and Zhou (2007).
+    The result in Part (ii) satisfies DOOM.
+When true errors are jointly independent, it equals $-\delta\overline{\Delta}$, and its comparative statics with respect to $\nu_{i}$ and $\Delta_{i}$, $i\in\mathcal{M}$ are generally ambiguous.
+Part (i) of Theorem 2 shows that on the one hand, misspecified learning of the covariance matrix does not interfere with that of source biases, leaving the DM’s long-run misperceptions about the latter unaffected (more on this later).
+On the other hand, the inconsistency caused by misspecifications makes signals appear noisier to the DM than they truly are (in the Loewner order), in that the perceived covariance matrix equals its true counterpart, plus a rank-one update $\Delta\Delta^{\top}$ that is positive definite.
 Consequently, when making predictions about the true state, the DM relies less on signal realizations compared to when he knows the true covariance matrix from the outset.
-This is shown to mitigate the impact of misspecified learning on his behavior, resulting in a less biased prediction of the state on average.
-To better understand the bias mitigation effect, we decompose the distortion of the DM’s long-run behavior as follows:
+As shown in Part (ii) of Theorem 2, this mitigates the impact of misspecified learning on his behavior, resulting in a less biased prediction on average.
+Turning to the comparative statics in Part (iii), what is new here is the ambiguity associated with correcting misspecifications, whose impact is unambiguously positive in the baseline model (recall Part (ii) of Proposition 1).
+The ambiguity arises from compound misspecified learning: Correcting the misspecifications associated with familiar sources makes the environment appear less noisy to the DM than before.
+This reduces the bias mitigation effect discussed above and, as confirmed by numerical analysis, could undo the benefit of misspecification corrections.
+Our finding sheds light on the behavioral and welfare consequences of the debiasing programs introduced in Section 4.
+Years of economic and psychological research have yielded mixed insights into the efficacy of these programs.
+For example, while the success of using one’s own mental resources to intercept implicit biases has been documented in prominent case studies (recall Footnote 9), a recent meta-analysis of organizations that have implemented similar programs reveals a mixed pattern.
+This has led Greenwald and Lai (2020) to conclude that “Convincing evidence for the effectiveness of these strategies is not yet available in peer-reviewed publications.” Even worse, efforts to suppress stereotypes or to reduce intergroup animosity may backfire (Kulik et al., 2000; Kalev et al., 2006).
+We take no stand on this debate and instead caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$ closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect.
+Proof sketch.
+We outline the proof of Theorem 2, focusing on why compound misspecified learning does not further distort the DM’s perception about source biases, as well as the bias mitigation effect.
+A closer inspection of the KL minimization problem reveals that with compound misspecified learning, the DM’s perception about source biases is simply the solution to problem (2), given his limiting belief about the covariance matrix $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$.
+What remains to be verified is that the distortion in the covariance matrix, $\Delta\Delta^{\top}$, does not affect the solution to problem (2).
+Define $\widehat{A}$ as the submatrix of $\widehat{\Sigma}$ restricted to familiar sources, and $\widehat{C}$ as the submatrix capturing the covariances between familiar and unfamiliar sources.
+Our claim boils down to showing that $\widehat{C}\widehat{A}^{-1}\widetilde{\Delta}=CA^{-1}\widetilde{\Delta}$.
+By definition, $\widehat{A}=A+\widetilde{\Delta}\widetilde{\Delta}^{\top}$ and $\widehat{C}=C+\widehat{\Delta}\widetilde{\Delta}^{\top}$, where $A$ and $C$ are the true counterparts of $\widehat{A}$ and $\widehat{C}$, $\widetilde{\Delta}$ is the vector of the misspecifications associated with familiar sources, and $\widehat{\Delta}$ is the vector of optimized misperceptions about unfamiliar sources in the baseline model.
+Expanding $\widehat{C}\widehat{A}\widetilde{\Delta}-CA\widetilde{\Delta}$ yields
+$\widehat{C}\widehat{A}^{-1}\widetilde{\Delta}-CA^{-1}\widetilde{\Delta}$ $=(C+\widehat{\Delta}\widetilde{\Delta}^{\top})(A^{-1}-\frac{A^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}})\widetilde{\Delta}-CA^{-1}\widetilde{\Delta}$
+$=\underbrace{-\frac{CA^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}}_{\text{(I)}}+\underbrace{\frac{\widehat{\Delta}\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}}_{\text{(II)}}=0,$
+where the first equality uses the Sherman-Morrison formula to calculate $\widehat{A}^{-1}$, and the last equality uses $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ (recall Theorem 1).
+Intuitively, distortions in the covariance matrix affect the DM’s perception about source biases through two channels.
+The first arises from familiar sources and is captured by $\widehat{A}^{-1}-A^{-1}=-\frac{A^{-1}\widetilde{\Delta}\widetilde{\Delta}^{\top}A^{-1}}{1+\widetilde{\Delta}^{\top}A^{-1}\widetilde{\Delta}}$.
+The second concerns the covariances between familiar and unfamiliar sources and is captured by $\widehat{C}-C=\widehat{\Delta}\widetilde{\Delta}^{\top}$.
+To assess the overall impact, we project the first effect onto unfamiliar dimensions using $C$.
+Since this same projection is used to derive the optimal misperceptions about unfamiliar sources in the baseline model, i.e., $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$, the two effects must cancel each other.
+Turning to the bias mitigation effect, we decompose the distortion in the DM’s long-run prediction as follows:
 $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ $=\mathbf{1}_{N}^{\top}\widehat{\Sigma}^{-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})$
 $=\underbrace{-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta}_{\text{(I)}}+\underbrace{\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*})}_{\text{(II)}}\underbrace{-\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})\Delta}_{\text{(III)}}.$
-Effect (I) in the above expression is due to the misspecified learning of source biases and is shown to equal $-\overline{\Delta}$ in the baseline model.
+Effect (I) in the above expression is due to the misspecified learning of source biases and is the focus of the baseline model.
 Effect (II) arises from the misspecified learning of the covariance matrix and has zero mean under the true signal distribution.
@@
 […]
-Turning to comparative statics: In light of Theorem 1(ii), it is unsurprising that increasing signal-to-noise ratios could have unintended consequences on the outcomes of misspecified learning.
-What is new here is the ambiguity associated with correcting misspecifications, whose comparative statics in the baseline model were straightforward.
-The ambiguity arises from compound misspecified learning: By correcting the misspecifications associated with familiar sources, we make the environment appear less noisy to the DM than before.
-This reduces the bias mitigation effect discussed above, i.e., $\delta=(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ increases, and in some cases could undo the benefit of misspecification corrections.
-Theorem 2 sheds light on the behavioral and welfare consequences of the debiasing programs introduced toward the end of Section 3.2.
-Years of economic and psychological research have yielded mixed insights into the efficacy of these programs.
-For example, while the success of using one’s own mental resources to intercept implicit biases has been documented in prominent case studies (recall Footnote 8), recent meta-analysis of organizations that have implemented programs of similar kinds reveals a mixed pattern, leading Greenwald and Lai (2020) to conclude that “Convincing evidence for the effectiveness of these strategies is not yet available in peer-reviewed publications.” Even worse, efforts to suppress stereotypes or to reduce intergroup animosity may backfire (Kulik et al., 2000; Kalev et al., 2006), leading scholars including Bohnet (2016) to advocate alternative strategies such as evidence-based continuing learning.
-We take no stand on this debate and instead caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect.
+5.2 Multidimensional state and signals
+This section generalizes the baseline model to encompass multidimensional states and signals.
+In each period $t\in\mathbb{N}$, Nature draws the true state $\omega_{t}$ from a $K$-dimensional Gaussian distribution with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{0}$.
+Each dimension represents a distinct skill, such as problem solving, communication, teamwork, etc.
+The DM consults a set $\mathcal{S}=\{1,\cdots,N\}$ of information sources before making a prediction about the true state.
+The signal generated by source $i$ in period $t$ is $X_{i,t}=b_{i}+\omega_{t}+\epsilon_{i,t}$, where $b_{i}\in\mathbb{R}^{K}$ is the $K$-dimensional bias of the source, and $\epsilon_{i,t}$ is a state-independent Gaussian error with mean $\mathbf{0}_{K}$ and covariance matrix $\Omega_{i}$.
+To focus on the correlations between different dimensions, we assume that errors are jointly independent across sources, while allowing different dimensions of an error and the true state to exhibit correlations.
+The true bias vector is $b^{*}\coloneqq[b_{1}^{*},\cdots,b_{N}^{*}]^{\top}\in\mathbb{R}^{NK}$.
+Initially, the DM misspecifies the biases of the sources in =$\mathcal{M}=\{1,\cdots,M\}$ as $\tilde{b}_{i}\neq b_{i}^{*}$, $i\in\mathcal{M}$, while being uncertain of the biases of the remaining sources.
+The support of his prior is $\{b\in\mathbb{R}^{NK}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\}$.
+Everything else is as in the baseline model.
+In particular, the DM knows the true covariance matrix of the signals, which in the current setting equals $\text{diag}(\Omega_{1},\cdots,\Omega_{N})+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\otimes\Omega_{0}$ ($\otimes$ denotes the Kronecker product).
+The next proposition establishes the analog of Proposition 1 in the current multidimensional setting.
+Proposition 3.
+Redefine $\overline{\Delta}\coloneqq(\sum_{i=0}^{M}\Omega_{i}^{-1})^{-1}\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}.$ Then $\Delta_{i}=\overline{\Delta}$ $\forall i\in\mathcal{S}-\mathcal{M}$ and $\widehat{\emph{BLP}}(X)-\emph{\text{BLP}}^{*}(X)=-\overline{\Delta}$.
+The next example utilizes Proposition 3 to investigate the compound impact of misspecified learning and the halo effect in employee assessments.
+Example 3.
+The DM faces a stream of female employees whose analytical and communication skills (hereinafter denoted by $A$ and $B$, respectively) are drawn independently from the standard normal distribution, i.e., $\Omega_{0}=\mathbb{I}_{2\times 2}$.
+The DM consults two sources about employees’ skills but misspecifies the biases of source one: Although the latter undervalues female employees’ analytical skill and overvalues their communication skill, the DM perceives it as impartial, i.e., $\Delta^{A}>0$ and $\Delta^{B}<0$.
+What the DM does know is that source one’s assessments of different skills exhibit a halo effect (Bohnet, 2016).
+That is, a positive assessment about one’s analytical skill enhances the assessment of her communication skill, and vice versa.
+Let $\Omega_{1}=\footnotesize{\begin{bmatrix}v&v\rho\\ v\rho&v\end{bmatrix}}$ be the covariance matrix of the errors associated with source one’s signal.
+We capture the halo effect by assuming that $\rho>0$.
+We examine how the DM’s misspecification about source one distorts his learning about source two, as well as his prediction about employees’ skills, in the presence of the halo effect.
+Solving $\overline{\Delta}$ using Mathematica yields
+$\begin{bmatrix}\overline{\Delta}^{A}\\ \overline{\Delta}^{B}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{(1+v)\Delta^{A}-\rho v\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{+}\\ \underbrace{\frac{-\rho v\Delta^{A}+(1+v)\Delta^{B}}{(1+(1-\rho)v)(1+v+\rho v)}}_{-}\end{bmatrix}\text{ and }\begin{bmatrix}\frac{\partial\overline{\Delta}^{A}}{\partial\rho}\\ \frac{\partial\overline{\Delta}^{B}}{\partial\rho}\end{bmatrix}=\begin{bmatrix}\underbrace{\frac{2\rho v^{2}(1+v)\Delta^{A}-v(1+v(2+v+\rho^{2}v))\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{+}\\ \underbrace{\frac{-v(1+v(2+v+\rho^{2}v))\Delta^{A}+2\rho v^{2}(1+v)\Delta^{B}}{(1+(1-\rho)v)^{2}(1+v+\rho v)^{2}}}_{-}\end{bmatrix}.$
+In words, the DM ends up perceiving source two as too lenient about women’s analytical skill and too critical about their communication skill.
+This misperception is exacerbated by the halo effect, whose presence requires that the DM subtract the communication score from the analytical score to discern the fundamental from the errors.
+It is through this channel that the DM’s unconscious, negative, bias about the communication score assigned by source one creeps into his learning about the analytical score assigned by source two, making the latter appear more positively biased than it actually is.
+In the long run, the DM underestimates female employees’ analytical skill and overestimates their communication skill.
+To the extent that these predictions are key to HR decision making, such as task assignment, hiring, and retention, they can potentially reinforce occupational segregation and stereotypes that are pervasive in the workplace (Bertrand, 2020).
+A comprehensive exploration of these mechanisms is better left for future research.
+While it is tempting to suggest, based on the simple model presented above, that companies focus rater training on correcting for halo effect, complications may arise in the presence of multiple misspecified sources, as correcting for the halo effect exhibited by one source makes those exhibited by other sources more salient.
+The total effect could be an exacerbation or a mitigation of the distortion, for the same reason that renders the comparative statics of signal precision ambiguous in Proposition 1.
 6 Further applications
-In this section, we utilize our theoretical framework to analyze two further applications: misperceived media bias and project investment with misspecified learning.
-5.1 Misperceived media bias
+In the Online Appendix, we utilize our theoretical framework to analyze two additional applications: misperceived media bias and project investment under misspecified learning.
+In this section, we summarize our main findings, highlighting the implications of DOOM for policy interventions and regulations.
+Additional results and their analysis are relegated to the Online Appendix.
+Misperceived media bias.
+Consider a Republican voter learning about the horizontal bias of various news organizations.
+The misspecified sources consist of the voter’s personal sources and established networks.
+The former are viewed by the voter as unbiased due to bias blind spot (Pronin et al., 2002).
+The latter include, for instance, Fox News and CNN, which he incorrectly believes to be neutral, due to familiarity bias, and excessively left-leaning, due to liberal media bias, respectively (Pennycook and Rand, 2021; Hassell et al., 2020).
+These sources form the yardsticks with which he will learn about new entrants, including accounts on social media.
+Our voter seeks to take a position that minimizes the quadratic distance from the truth.
+The optimal position in the long run is the BLP.
+In the Online Appendix, we fully describe the distortionary effects of misspecified learning on the voter’s long-run beliefs, behavior, and objective welfare (i.e., the expected quadratic loss under the true distribution).
+The welfare loss is $-\overline{\Delta}^{2}$ when signal variances are known to the DM, and $-\delta\overline{\Delta}^{2}$ when they are initially unknown.
+Importantly, both quantities satisfy DOOM.
+DOOM implies that fact-checking interventions may have unintended welfare consequences.
+For one, sources with initially unknown biases are irrelevant to the distortionary effect of misspecified learning.
+By increasing the signal-to-noise ratio of these sources, the voter’s objective welfare increases unambiguously under both the correctly specified model and the misspecified model.
+[^14] The exact numbers are $-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}-\text{distortion}$, respectively.
+In contrast, increasing the signal-to-noise ratio of misspecified sources has ambiguous effects on $\overline{\Delta}^{2}$ and $\delta\overline{\Delta}^{2}$, as shown previously.
+Hence policy interventions that combat fake news by improving source accuracy may backfire when they modulate sources about which people already hold dogmatic misperceptions.
+Given that fact-checking is resource intensive, our model suggests prioritizing social media accounts over traditional media outlets, particularly when the specific nature of misspecifications is unclear to the fact-checker.
+Project investment.
+Consider a manager allocating capital across various projects.
+The manager can acquire information about project returns from various sources, ranging from internal analysts to uninvolved consultants and robo-advisors.
+Internal analysts, for instance, exhibit overoptimism bias that is neglected by the manager, due to the latter’s confirmation bias or the tendency to overly trust sources in his inner circle (Bieske et al., 2023; Malmendier, 2018).
+Uninvolved consultants and robo-advisors are novel and their biases are initially unknown.
+We consider a manager with CARA utility and characterize the distortion in his portfolio choice and objective welfare due to misspecified learning.
+Interestingly, this is succinctly captured by terms depending on $\overline{\Delta}$.
+A consequence of DOOM is that the standard practice of seeking advice from “uninvolved consultants” as a means to counteract entrenched managerial biases (Guenzel and Malmendier, 2020) may offer limited benefits.
+Furthermore, technological advances that reduce data collection and processing costs could affect our manager differently, depending on whether they increase the signal-to-noise ratio of familiar or novel sources.
+Appendix A Proofs
+A.1 Preliminaries
+Proof outline.
+To begin, we describe the problem studied by HKS and discuss the relationship between our analyses.
+HKS examine a misspecified Bayesian learning model in which the DM observes a $D$-dimensional Gaussian signal with mean $Mf\in\mathbb{R}^{D}$ and covariance matrix $\Sigma\in\mathbb{R}^{D\times D}$ in each period $t\in\mathbb{N}$.
+Initially, the DM knows the transformation matrix $M\in\mathbb{R}^{D\times N}$ — which captures group memberships — while misspecifying the first entry of the fundamental $f\in\mathbb{R}^{N}$.
+He may or may not know the true covariance matrix $\Sigma^{*}$.
+In Case (I) where $\Sigma^{*}$ is known, the support of the DM’s prior is
+$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1}\text{ and }\Sigma=\Sigma^{*}\}.$
+In Case (III) where $\Sigma^{*}$ is initially unknown, the support becomes
+$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1}\text{ and }\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}}\},$
+where $\underline{\underline{\lambda}}$ can be any positive number that is smaller than $1$ and $\lambda_{\min}(\Sigma^{*})$.
+Over time, the DM updates his beliefs about $(f,\Sigma)$ based on the prior and signal realizations.
+For both cases, HKS demonstrate that the DM’s limiting posterior beliefs concentrate on the KL minimizer and provide characterizations for the latter.
+The main difference between our analysis and HKS is that our DM uses the outcome of misspecified learning to predict a random state.
+To facilitate analysis of the prediction problem, we make slightly different model assumptions from that of HKS, as outlined below:
+  1. D1
+    In our setting, $f=b$, $M=Identity$, and the DM can misspecify multiple entries of $b$.
+  2. D2
+    We assume that signals equal their biases, plus the true state and state-independent errors.
+The covariance matrix of the signals is thus a rank-one update of the covariance matrix of the errors, i.e., $\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
+This structure is not imposed by HKS, who allow $\Sigma$ to be any SPD matrix with a bounded minimum eigenvalue.
+How do these differences impact the analysis?
+For establishing the convergence of posterior beliefs to the KL minimizer, D1 proves to be inconsequential, whereas D2 requires that we bound the eigenvalues of $\Sigma$ based on the assumptions we make about $\Omega$.
+To this end, we apply spectral stability theories to bound the changes in the eigensystem that result from rank-one updates to an existing matrix.
+Details are in Appendix A.2, where we reproduce HKS’ convergence proof line by line.
+When it comes to characterizing the limiting outcome, our proof can handle multi-sourced misspecifications and generalizes HKS’ guess-and-verify approach that is tailored for single-sourced misspecifications.
+When the covariance matrix is also a subject of learning, D2 introduces additional complications, as we need to convert the DM’s limiting belief about $\Sigma$ into a limiting belief about $\Omega$, using again spectral stability theories.
+Details are in Appendix A.3.
+Notations.
+We follow the notational conventions developed in the main text, using $x$ to denote the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$ and $\text{diag}(x)$ to denote the diagonal matrix created from $x$.
+For each $k\in\mathbb{N}$, we use $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ to denote $k$-vectors of ones and zeros, respectively, and $\mathbb{I}_{k\times k}$ to denote the $k\times k$ identity matrix.
+Finally, $\otimes$ denotes the Kronecker product, $a\vee b$ stands for $\max\{a,b\}$, and $a\wedge b$ for $\min\{a,b\}$.
+A.2 Useful lemmas and their proofs
+Lemma 3.
+Fix any symmetric matrix $G\in\mathbb{R}^{N\times N}$ and let $\lambda_{k}(G)$ denote the $k^{th}$ largest eigenvalue of $G$, $k=1,\cdots,N$.
+The following statements are true for any $x\in\mathbb{R}^{N}$:
+$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+x^{\top}x)$
+[…]
+Proof.
+Weyl’s inequality (Weyl, 1912) states that $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-N}(A+B)$ for any Hermitian matrices $A$ and $B$ belonging to an inner product space of dimension $N$.
+Letting $A=G$ and $B=xx^{\top}$ yields
+$\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)+\lambda_{N}(xx^{\top})=\lambda_{1}(G)+0\leq\lambda_{1}(G+xx^{\top})$ (13)
+[…]
+where the equalities exploit the fact that $xx^{\top}$ has rank one and so has two eigenvalues: 00 and $\text{tr}(xx^{\top})$, and that $\text{tr}(xx^{\top})=\text{tr}(x^{\top}x)=x^{\top}x>0$.
+Interchanging the roles of $G+xx^{\top}$ and $G$ in the above derivation yields
+$\lambda_{N}(G)\leq\lambda_{1}(G+xx^{\top})+\lambda_{N}(-xx^{\top})=\lambda_{1}(G+xx^{\top})-x^{\top}x\leq\lambda_{1}(G)$ (15)
+[…]
+Together, (13), (14), and (16) imply that
+$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+xx^{\top}),$
+that is, the first inequality in the statement of the lemma. Additionally, (13), (14), and (15) imply that
+$\lambda_{1}(G)\vee(\lambda_{N}(G)+xx^{\top})\leq\lambda_{1}(G+xx^{\top})\leq\lambda_{1}(G)+x^{\top}x,$
+that is, the second inequality in the statement of the lemma. ∎
+Proof of Lemma 2.
+In the environment considered in Section 5.1, the support of the DM’s prior beliefs is
+$\Theta=\begin{Bmatrix}(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\text{ for some }\Omega\in\Lambda\end{Bmatrix},$
+where $\Lambda$ denotes the set of $N\times N$ SPD matrices whose minimum eigenvalues are bounded below by $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$.
+The proof presented below verifies that this environment satisfies the assumptions of Berk (1966), hence the DM’s limiting posterior beliefs concentrate on the set of KL minimizers.
+Part (i) of Berk’s assumption says that the subjective density of the signals:
+$l(x\mid b,\Sigma)=\frac{1}{\sqrt{(2\pi)^{N}\text{det}\Sigma}}\exp\left(-\frac{1}{2}(x-b)^{\top}\Sigma^{-1}(x-b)\right),$
+is continuous in $(b,\Sigma)$ on $\Theta$. This is clearly true in our setting, since $\text{det}\Sigma$ and $\Sigma^{-1}$ are both continuous in $\Sigma$.
+Part (ii) of Berk’s assumption says that the subjective density equals zero on a set of measure zero under the true signal distribution.
+This is clearly true, since the subjective density is always positive.
+Part (iii) of Berk’s assumption says that for every $(b,\Sigma)\in\Theta$, there exists an open neighborhood $U\subset\Theta$ such that
+$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in U}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]<\infty.$
+Let $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ denote the minimum and maximum eigenvalues of $\Sigma^{\prime}$, respectively. Then
+$|\log l(x\mid b^{\prime},\Sigma^{\prime})|$ $=|\frac{1}{2}\log((2\pi)^{N}\text{det}\Sigma^{\prime})+\frac{1}{2}(x-b^{\prime})^{\top}\Sigma^{{}^{\prime}-1}(x-b^{\prime})|$
+$\leq\frac{1}{2}\left(N|\log(2\pi\lambda_{\max}(\Sigma^{\prime}))|\vee|\log(2\pi\lambda_{\min}(\Sigma^{\prime}))|+\frac{1}{\lambda_{\min}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right),$
+where the inequality uses the following facts: (i) the determinant of a matrix equals the product of its eigenvalues; (ii) the Rayleigh quotient $(x-b^{\prime})^{\top}\Sigma^{{}^{\prime}-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded above by $\lambda_{\max}(\Sigma^{\prime-1})=1/\lambda_{\min}(\Sigma^{\prime})$; (iii) $\Sigma^{\prime}$ is SPD, and so $\lambda_{\min}(\Sigma^{\prime})>0$.
+Since $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ are continuous in the entries of $\Sigma^{\prime}$, the last expression is finite on every finite neighborhood $U$.
+So far, the proof follows the same steps as HKS’ proof for their Theorem A.1 Case (III).
+The last step of the proof differs slightly, due to our restriction of the covariance matrix to those of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
+Part (iv) of Berk’s assumption says that for every $c\in\mathbb{R}$, there exists $D\subset\Theta$ with a compact complement $\Theta-D$ such that
+$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in D}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]\leq c.$
+Fix any $k_{1},k_{2}>0$.
+Let $D$ be the set of $(b^{\prime},\Sigma^{\prime})$ such that $\|b^{\prime}-b^{*}\|>k_{1}$ and the corresponding $\Omega^{\prime}$ satisfies $\lambda_{\max}(\Omega^{\prime})>k_{2}$.
+By Lemma 3, $\lambda_{\max}(\Sigma^{\prime})\geq\lambda_{\max}(\Omega^{\prime})>k_{2}$, and $\lambda_{\min}(\Sigma^{\prime})\geq\lambda_{\min}(\Omega^{\prime})\geq\underline{\lambda}$.
+Substituting these bounds into HKS’ proof and replicating line by line, we obtain that
+$\log l(x\mid b^{\prime},\Sigma^{\prime})$
+$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
+$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee\|x-b^{\prime}\|^{2})\right).$
+In the above derivation, the first inequality uses the following facts: (i) $\text{det}\Sigma^{\prime}\geq\lambda_{\min}^{N-1}(\Sigma^{\prime})\lambda_{\max}(\Sigma^{\prime})$; (ii) $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$; and (iii) the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded below by $\lambda_{\min}(\Sigma^{\prime-1})=1/\lambda_{\max}(\Sigma^{\prime})$.
+The second inequality follows from minimizing $\log\lambda_{\max}(\Sigma^{\prime})+\|x-b^{\prime}\|^{2}/\lambda_{\max}(\Sigma^{\prime})$ over the range $\lambda_{\max}(\Sigma^{\prime})>k_{2}$.
+To obtain a lower bound for $\|x-b^{\prime}\|^{2}$, we note that $x-b^{\prime}=x-b^{*}+(b^{*}-b^{\prime})=(\omega+\epsilon)+(b^{*}-b^{\prime})$, hence $\|x-b^{\prime}\|^{2}\geq\|\omega+\epsilon\|^{2}+\|b^{*}-b^{\prime}\|^{2}-2\|\omega+\epsilon\|\|b^{*}-b^{\prime}\|$ by the Cauchy-Schwarz inequality.
+Since $\|b^{*}-b^{\prime}\|>k_{1}$ by assumption, for any given value of $\|\omega+\epsilon\|<k_{1}/2$, $\|b^{*}-b^{\prime}\|^{2}+\|\omega+\epsilon\|^{2}-2\|\omega+\epsilon\|\|b^{*}-b^{\prime}\|$ is minimized at $\|b^{*}-b^{\prime}\|=k_{1}$ and so is bounded below by $k_{1}^{2}/4$.
+Consequently, the last line of the above inequality is bounded above by
+$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right).$
+Meanwhile, $\log l(x\mid b^{\prime},\Sigma^{\prime})$ is bounded above by
+$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
+$\leq-\frac{1}{2}N\log\lambda_{\min}(\Sigma^{\prime})\leq N|\log\underline{\lambda}|/2,$
+where the first inequality is obvious, and the second one exploits the fact that $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}$ and the assumption that $\underline{\lambda}<1$.
+Taken together, we obtain that
+$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in\Theta}\log l(x\mid b^{\prime},\Sigma^{\prime})\right]\leq\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\mathbf{1}_{\|\omega+\epsilon\|\leq k_{1}/2}\sup_{(b^{\prime},\Sigma^{\prime})\in D}l(x\mid b^{\prime},\Sigma^{\prime})\right]+N|\log\underline{\lambda}|/2$
+$\leq-\frac{1}{2}\left[N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right]\mathbb{P}(\|\omega+\epsilon\|\leq k_{1}/2)+N|\log\underline{\lambda}|/2,$
+where the last line becomes arbitrarily negative as $k_{1},k_{2}\rightarrow+\infty$.
+It remains to verify that $\Theta-D$ is compact.
+By definition, this set consists of vectors $b^{\prime}$ such that $\|b^{\prime}-b^{*}\|\leq k_{1}$, paired with covariance matrices of the form $\Omega^{\prime}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$ such that the eigenvalues of $\Omega^{\prime}$ belong to $[\underline{\lambda},k_{2}]$.
+The last set is compact, and this completes the proof.
+∎
+Proof of Lemma 1.
+In the environment considered in Section 2, the support of the DM’s prior beliefs is
+$\Theta=\{(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\},$
+where $\Omega^{*}$ denotes the true covariance matrix of the errors.
+Since $\Omega^{*}\in\Lambda$, all that is needed is to replace every $\Sigma$ and $\Sigma^{\prime}$ in the proof of Lemma 2 with $\Sigma^{*}\coloneqq\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
+∎
+A.3 Proofs of theorems and propositions
+Proof of Proposition 2.
+We already defined and derived the constrained least squares estimator in Footnote 7. It remains to derive the best linear predictor of the state without the normality assumption.
+Under the specification $X=b+\omega\mathbf{1}_{N}+\epsilon$ where $\omega$ and $X-b$ both have mean zero, the best linear predictor is given by a vector $\beta$ that minimizes the prediction squared-error $\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$.
+Since
+$\mathbb{E}[\left(\omega-\beta^{\top}(X-b)\right)^{2}]$
+$=\mathbb{E}[\omega^{2}]+\mathbb{E}[\beta^{\top}(X-b)(X-b)^{\top}\beta]-2\mathbb{E}[\beta^{\top}(X-b)\omega]$
+$=1+\beta^{\top}\Sigma^{*}\beta-2\beta^{\top}\mathbf{1}_{N},$
+where the last line is minimized at $\beta=\Sigma^{*-1}\mathbf{1}_{N}$, the BLP is $\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b).$
+Proof of Theorem 2.
+We only prove Part (i) of the theorem. Part (ii) of the theorem has already been established in the main text. Part (iii) of the theorem follows from straightforward algebra.
+To begin, recall the KL miminization problem considered in Case (III) of HKS:
+$\min_{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}}D_{KL}(f^{*},\Sigma^{*}\mid f,\Sigma)\text{ s.t.
+}f_{1}=\tilde{f}_{1}\text{ and }\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}},$
+where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$.
+The solution $(\hat{f},\widehat{\Sigma})$ to this problem is shown to satisfy two properties: First, $\Delta f_{i}\coloneqq\hat{f}_{i}-f_{i}^{*}$ as if the DM knows the true covariance matrix $\Sigma^{*}$.
+Second, $\widehat{\Sigma}=\Sigma^{*}+(M\Delta f)(M\Delta f)^{\top}$.
+The assumption that only a single entry of $f$ is misspecified plays no role in this result, verified by checking the proof of HKS line by line.
+Letting $f=b$ and $M=\mathbb{I}_{N\times N}$ in the above argument leads us to conclude that, in the environment considered in Section 5.1, the solution to the KL minimization problem would be $\Delta=\hat{b}_{i}-b_{i}^{*}$ and $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ had our DM been able to choose any SPD matrix whose minimum eigenvalue is bounded below by $\underline{\underline{\lambda}}$.
+The complication arises from the restriction that the covariance matrix must take the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, where $\Omega$ is SPD with its minimum eigenvalue bounded below by $\underline{\lambda}<1\wedge\lambda_{\min}(\Omega^{*})$.
+We are left to verify that the solution to the unconstrained problem satisfies this additional constraint.
+If so, then it must also solve the constrained optimization problem.
+To verify, note that the solution to the unconstrained problem: $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, is attained by $\widehat{\Omega}=\Omega^{*}+\Delta\Delta^{\top}$.
+$\widehat{\Omega}$ is clearly SPD, and $\lambda_{\min}(\widehat{\Omega})\geq\lambda_{\min}(\Omega^{*})>\underline{\lambda}$ by Lemma 3.
+This completes the proof.
+∎
+Proof of Proposition 3.
+With multidimensional skills, the true covariance matrix of the signals becomes
+$\text{diag}\left(\Omega_{1},\cdots,\Omega_{N}\right)+\left(\mathbf{1}_{N}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{N}^{\top}\otimes\mathbb{I}_{K\times K}\right),$
+where $\otimes$ denotes the Kronecker product. Write this matrix as
+$\left[\begin{array}[]{c|c}A_{KM\times KM}&B_{KM\times K(N-M)}\\ \hline\cr C_{K(N-M)\times KM}&D_{K(N-M)\times K(N-M)}\end{array}\right],$
+where
+$A=\underbrace{\text{diag}(\Omega_{1},\cdots,\Omega_{M})}_{G}+\underbrace{\left(\mathbf{1}_{M}\otimes\mathbb{I}_{K\times K}\right)}_{x}\Omega_{0}\underbrace{\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right)}_{y}$
+and
+$C=\left(\mathbf{1}_{N-M}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right).$
+Since $A$ is a rank-$K$ update of an invertible matrix,
+$A^{-1}$ $=G^{-1}-G^{-1}x\left(\Omega_{0}^{-1}+yG^{-1}x\right)^{-1}yG^{-1}$
+$=\text{diag}(\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1})-\begin{bmatrix}\Omega_{1}^{-1}\\ \vdots\\ \Omega_{M}^{-1}\end{bmatrix}\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\begin{bmatrix}\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1}\end{bmatrix},$
+where the first equality uses the Woodbury formula (Woodbury, 1950). Substituting the expressions for $A^{-1}$ and $C$ into Theorem 1 yields
+$\widehat{\Delta}$ $=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\otimes\left(\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}\right)$
+$=\mathbf{1}_{N-M}\otimes\overline{\Delta},$
+where the second line uses the fact that
+$\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=0}^{M}\Omega^{-1}_{i}=\Omega_{0}\Omega_{0}^{-1}=\mathbb{I}_{K\times K},$
+and so
+$\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}=\left(\sum_{i=0}^{M}\Omega_{i}\right)^{-1}.$
+Additionally,
+$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ $=-\left(\mathbf{1}_{M}^{\top}\otimes\Omega_{0}\right)A^{-1}\widetilde{\Delta}$
+$=-\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}$
+$=-\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\sum_{i=1}^{M}\Omega_{i}^{-i}\Delta_{i}=-\overline{\Delta},$
+which completes the proof. ∎
+Online Appendix for
+“Learning Source Biases: Multisource Misspecifications and Their Impact on Predictions”
+by Junnan He, Lin Hu, Matthew Kovach and Anqi Li
+Appendix O.1 Further Applications
+In this appendix, we utilize the theories developed in the main text to analyze two further applications. The assumption of independent errors is made throughout.
+O.1.1 Misperceived media bias
 A Republican voter consults several news sources on an evolving issue.
@@
 Others, such as social media sites operated by unknown people or bots, are unfamiliar.
-Initially, the voter is uncertain about the horizontal biases of the unfamiliar sources, while holding dogmatic, wrong beliefs about the biases of some familiar sources.
+Initially, the voter is uncertain about the horizontal biases of unfamiliar sources, while holding dogmatic, wrong, beliefs about the biases of some familiar sources.
 As someone who grew up watching Fox News, he believes that this source is more neutral than it actually is due to familiarity bias (Pennycook and Rand, 2021).
-He underestimates the bias associated with his own opinions and judgments, in part due to the bias blind spot and third-person effect (Pronin et al., 2002; Davison, 1983), and in part because he is working toward self-enhancement and would like to view himself as middle-of-the-road (Dennis, 1988).
-[^11] Bias blind spot refers to the tendency for people to recognize biases in human judgment except when that bias is their own.
-The third-person effect predicts that people perceive mass media messages to have a greater effect on others than on themselves.
+He underestimates the bias associated with his own opinions and judgments, in part due to bias blind spot and third-person effect (Pronin et al., 2002; Davison, 1983), [^15] Bias blind spot refers to the tendency for people to recognize biases in human judgment except when that bias is their own.
+The third-person effect predicts that people perceive mass media messages to have a greater influence on others than on themselves.
+and in part because he is working toward self-enhancement and would like to view himself as middle-of-the-road (Dennis, 1988).
 Finally, he believes in a liberal media bias like many other Republicans and displays excessive hostility towards CNN (Hassell et al., 2020).
@@
 They can be captured by letting ={personal source,Fox $\mathcal{M}=\{\text{personal source},\text{Fox News},\text{CNN}\}$ and $\tilde{b}_{i}<b_{i}^{*}$ $\forall i\in\mathcal{M}$ in our framework.
-In each subsequent period $t=1,2,\cdots$, the voter takes a position $a_{t}\in\mathbb{R}$ and incurs a loss $-(a_{t}-\omega_{t})^{2}$ equal to the quadratic distance between this position and the true state of the world.
+In each period $t\in\mathbb{N}$, the voter takes a position $a_{t}\in\mathbb{R}$ and incurs a loss $-(a_{t}-\omega_{t})^{2}$ equal to the quadratic distance between this position and the true state of the world.
 The optimal position in the long run equals $\text{BLP}^{*}(X)$ under the correctly specified model and $\widehat{\text{BLP}}(X)$ under the misspecified model.
 The expected utilities generated by these positions, evaluated under the true signal distribution, equal $\text{EU}^{*}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}\left(\omega-\text{BLP}^{*}(X)\right)^{2}$ and $\widehat{\text{EU}}\coloneqq-\mathbb{E}_{(b^{*},\Sigma^{*})}(\omega-\widehat{\text{BLP}}(X))^{2}$, respectively.
-The difference between them captures the welfare loss from misspecified learning.
-Proposition 4.
-$\widehat{\text{EU}}-\text{EU}^{*}$ equals $-\overline{\Delta}^{2}$ in the baseline model and $-\delta\overline{\Delta}^{2}$ in the environment described in Section 4.3.
-Proposition 4, together with Theorems 1 and 2, fully describes the distortionary effects of misspecified learning on the voter’s long-run beliefs, behavior, and welfare.
-Since $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}<0$ $\forall i\in\mathcal{M}$ in the current example, $\overline{\Delta}<0$.
-The distortion of the voter’s worldview due to misspecifications constitutes a liberal media bias of magnitude $|\overline{\Delta}|$.
+Their difference captures the objective welfare loss from misspecified learning.
+Proposition O.1.
+$\widehat{\text{EU}}-\text{EU}^{*}$ equals $-\overline{\Delta}^{2}$ if the covariance matrix of the signals is known to the DM and $-\delta\overline{\Delta}^{2}$ if it is initially unknown.
+Proposition O.1, together with Proposition 1 and Theorem 2 of the main text, fully describes the distortionary effects of misspecified learning on the voter’s long-run beliefs, position, and objective welfare.
+In the current example, $\overline{\Delta}<0$ because $\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}<0$ $\forall i\in\mathcal{M}$.
+The distortion of the voter’s worldview thus constitutes a liberal media bias of magnitude $|\overline{\Delta}|$.
 Our results establish a mechanism through which this distortion can spill into the online news sphere and facilitate analysis of its behavioral and welfare consequences.
@@
 The hostile media effect refers to the phenomenon that opposing partisans perceive news coverage to be more biased against their side than it actually is.
-Evidence supporting the existence of this phenomenon abounds in political science, communication, and psychology (Feldman, 2014), [^12] In a classical experiment conducted by Arpan and Raney (2003), student subjects read a balanced story about their hometown college football team in one of three newspapers: home town, rival town, or neutral town outside the state.
+Ample evidence in political science, communication, and psychology documents the existence of this phenomenon (Feldman, 2014), [^16] In a classic experiment conducted by Arpan and Raney (2003), student subjects read a balanced story about their hometown college football team in one of three newspapers: home town, rival town, or neutral town outside the state.
 Subjects viewed the coverage as favoring the rival university, even if it was believed to be from the neutral town.
@@
 although research on learning and belief formation about new online sources has remained sparse.
-We demonstrate how the hostile media effect in traditional media can spill into the digital news sphere, as the voter in our model overperceives the bias of any liberal source while underperceiving the bias of any conservative source in the long run.
+We demonstrate how the hostile media effect in traditional media can spill into the digital news sphere, as the voter in our model overperceives the bias of any liberal source while underperceiving the bias of any conservative source.
 False polarization is defined as the difference between the perceived distance between the positions of two groups and the actual distance between them.
 Its presence in the American public was rigorously established by Levendusky and Malhotra (2016b).
-[^13] These authors used the survey data gleaned from a nationally representative sample that asked about the subjects’ self-placements on an issue scale, as well as their placements of a typical Democratic and Republican voter on that scale.
-The data revealed that Americans significantly misperceive the public to be more divided along partisan lines than they are in reality.
-To the extent that individual human beings can now express their opinions online at almost no cost, our model predicts a false polarization in the digital news sphere, as well as a causal relationship between misspecified news bias and false polarization.
-The latter is consistent with the literature’s common attribution of false polarization to partisan media exposure (Levendusky and Malhotra, 2016a).
+[^17] These authors used survey data from a nationally representative sample that asked about the subjects’ self-placements on an issue scale, as well as their placements of a typical Democratic and Republican voter on that scale.
+The data revealed that Americans significantly misperceive the public to be more divided along partisan lines than in reality.
+Since individuals can now express their opinions online at almost no cost, our model predicts a false polarization in the digital news sphere.
+Furthermore, it suggests a causal relationship between misspecified news bias and false polarization, consistent with the literature’s common attribution of false polarization to partisan media exposure (Levendusky and Malhotra, 2016a).
 The existing literature is divided on the impact of (mis)perceived polarization and media bias on voters’ actual positions, with some authors such as Levendusky and Malhotra (2016a) documenting a moderation effect caused by false polarization.
@@
 As we increase the signal-to-noise ratios of these sources, the voter’s expected utility under the true model and, hence, the misspecified model, increases unambiguously ($\text{EU}^{*}=-(1+\sum_{i\in\mathcal{S}}\nu_{i})^{-1}$ and $\widehat{\text{EU}}=\text{EU}^{*}-{\overline{\Delta}}^{2}$).
-The same cannot be said about sources with misspecified biases, as increasing their signal-to-noise ratios generally has an ambiguous effect on the welfare distortion $\overline{\Delta}^{2}$ from misspecified learning.
-Our finding suggests that policy interventions that target misinformation and fake news through improving source accuracy (e.g., fact-checking) may entail unintended welfare consequences, especially when they operate by modulating sources that people already hold dogmatic misperceptions about.
+The same cannot be said about sources with misspecified biases, as increasing their signal-to-noise ratios generally has an ambiguous effect on the welfare distortion $\overline{\Delta}^{2}$.
+Our finding suggests that policy interventions that target misinformation and fake news through improving source accuracy (e.g., fact-checking) may entail unintended welfare consequences, especially when they modulate sources that people already hold dogmatic misperceptions about.
 When it comes to lesser-known sources, fact-checking always improves welfare.
-To the extent that professional fact-checking is costly and does not scale easily, the allocation of time and efforts should prioritize lesser-known sources over established ones.
+Given that professional fact-checking is resource intensive and does not scale easily, the allocation of time and efforts should prioritize lesser-known sources over established ones.
 This finding is, to our best knowledge, new to the literature on (political) misinformation (see Nyhan 2020 for a survey).
@@
 A large empirical literature studies the treatment effects of misperception recalibration (Bursztyn and Yang, 2022).
-While early research indicated the possibility of a “backfire effect” whereby recalibration efforts (e.g., provision of corrective information, cross-cutting exposure) end up reinforcing rather than reducing subjects’ misperceptions, an emerging research consensus finds that the backfire effect could be more tenuous than previously thought and that more attention should be paid to alternative issues, such as the durability and scalability of the interventions (Nyhan, 2021).
-We caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$, closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect studied in Section 4.3.
-5.2 Project investment with misspecified learning
+While early research indicated the possibility of a “backfire effect” whereby recalibration efforts (e.g., provision of corrective information, cross-cutting exposure) end up reinforcing rather than reducing subjects’ misperceptions, an emerging research consensus finds that the backfire effect could be more tenuous than previously thought, and that more attention should be paid to alternative issues, such as the durability and scalability of the interventions (Nyhan, 2021; Hartman et al., 2022).
+We caution that even if misspecifications can be partially corrected (bring $\Delta_{i}$, $i\in\mathcal{M}$ closer to zero), their behavioral and welfare consequences may remain ambiguous due to the compound learning effect: When source precision is initially unknown, the opportunity to learn about it can render the comparative statics of $\delta\overline{\Delta}$ and $\delta\overline{\Delta}^{2}$ ambiguous.
+Discussion.
+The economic literature on media bias is vast but is devoted to understanding actual media biases (Anderson et al., eds, 2016).
+A notable exception is made by Gentzkow et al.
+(2018) (GWZ), who develop a misspecified learning model of media bias.
+To model bias, GWZ introduce a payoff-irrelevant confounder, independent of the payoff-relevant state, and call a source biased if it exhibits a nonzero correlation with the confounder.
+The DM in their model misperceives his personal source as unbiased and uses this belief to guide his media consumption.
+The model of GWZ differs from ours in its construct and predictions, in addition to requiring different methodologies to analyze.
+First, GWZ examine an active learning model where the choice of media consumption is endogenous.
+Second, they model media bias as a correlation with a random confounder rather than as a horizontal bias, which necessitates the use of very different analytical tools from ours.
+A consequence of this modeling choice is that (distortions of) the DM’s long-run position has a zero unconditional mean, which contrasts with the prediction of our model.
+Finally, a central focus of GWZ is on the endogenous trust in news sources, defined as the latter’s (mis)perceived correlation with the true state.
+Trust is correctly specified throughout our paper and is known to the DM for the most part.
+When endogenized as a subject of misspecified learning, we find no effect on perceived news bias and only a moderation of the DM’s position due to the compound learning effect.
+O.1.2 Project investment under misspecified learning
 A manager invests an initial wealth of $W>0$ in $K>1$ projects whose returns $\omega_{k}$, $k=1,\cdots,K$ are independently drawn from the standard normal distribution.
 Prior to decision making, he inquires about the return of each project from a mix of familiar sources — his personal source, internal consultants — and unfamiliar sources — external consultants, robo-advisors.
-The signal generated by each source is the same as in the baseline model, with its bias now capturing how optimistic the source is about the return of the project.
+The signal generated by each source is the same as in the baseline model, with its bias now capturing how optimistic the source is about the project return.
 […]
-Over time, the manager observes signal realizations and updates his belief about source biases.
-His long-run belief underestimates the biases of sources that are optimistic about the returns of projects, while overestimating the biases of those that are neutral or pessimistic.
-We take this belief as given and examine its consequences for the manager’s investment decisions and welfare.
-For ease of notation, we assume, in the remainder of this section, that only sources that report on project $k=1$ have their biases misspecified by the manager.
-In addition, the manager has a CARA Bernoulli utility function with an Arrow-Pratt measure $\lambda>0$ of absolute risk aversion.
-Specifically, for each project $k=1,\cdots,K$, we use $\mu_{k}^{*}$ to denote a typical realization of the best linear predictor of its return under the correctly specified model.
+Over time, the manager observes signal realizations and updates his beliefs about source biases.
+His long-run beliefs underestimate the biases of sources that are optimistic about project returns, while overestimating the biases of those that are neutral or pessimistic.
+We examine how such beliefs influence his portfolio choice and objective welfare.
+For simplicity, we assume that only sources that report on project $k=1$ have their biases misspecified by the manager.
+The latter has a CARA Bernoulli utility function with an Arrow-Pratt measure $\lambda>0$ of absolute risk aversion.
+We work directly with the best linear predictors of asset returns.
+For each project $k=1,\cdots,K$, we use $\mu_{k}^{*}$ to denote a typical realization of the best linear predictor of its return under the correctly specified model.
 The counterpart of this quantity under the misspecified model is denoted by $\widehat{\mu}_{k}$.
-The variance $\eta_{k}^{-1}$ of the best linear predictor is known to decrease with the signal-to-noise ratios of the signals that report on project $k$.
-For any given profile of $\mu_{k}$, $k=1,\cdots,K$, the optimal investment decisions, denoted by $\alpha_{k}$, $k=1,\cdots,K$, solve for
+The variance $\eta_{k}^{-1}$ of the best linear predictor is known to decrease with the precision of the signals that report on project $k$.
+For any given profile of $\mu_{k}$, $k=1,\cdots,K$, the manager solves:
 $\max_{\alpha_{k},k=1,\cdots,K}\sum_{k=1}^{K}\alpha_{k}\mu_{k}-\frac{\lambda}{2}\sum_{k=1}^{K}\alpha_{k}^{2}\eta_{k}^{-1}\text{ s.t. }\sum_{k=1}^{K}\alpha_{k}\leq W.$
 Let $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ denote the optimal portfolio choices under the correctly specified model and the misspecified model, respectively.
-The difference $\Delta_{\alpha_{k}}\coloneqq\widehat{\alpha}_{k}-\alpha_{k}^{*}$ captures the distortionary effect of misspecified learning on the investment in project $k$.
-The expected utilities generated by $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$ are denoted by $\text{EU}^{*}$ and $\widehat{\text{EU}}$, respectively.
-Their difference captures the distortionary effect of misspecified learning on the manager’s welfare.
-Proposition 5.
+The difference $\Delta_{\alpha_{k}}\coloneqq\widehat{\alpha}_{k}-\alpha_{k}^{*}$ captures the distortion in the investment in project $k$.
+The expected utilities generated by $(\alpha_{k}^{*})_{k=1}^{K}$ and $(\widehat{\alpha}_{k})_{k=1}^{K}$, evaluated under the truth, are denoted by $\text{EU}^{*}$ and $\widehat{\text{EU}}$, respectively.
+Their difference captures the distortionary effect of misspecified learning on the manager’s objective welfare.
+Proposition O.2.
 $\Delta_{\alpha_{1}}=-\frac{\eta_{1}}{\lambda}\left(1-\frac{\eta_{1}}{\sum_{k=1}^{K}\eta_{k}}\right)\overline{\Delta},\text{ }\Delta_{\alpha_{j}}=\frac{\eta_{1}\eta_{j}}{\lambda\sum_{k=1}^{K}\eta_{k}}\overline{\Delta}\text{ }\forall j\neq 1,$
 […]
-and their absolute values are all decreasing in $\eta_{k}^{-1}$, $k=1,\cdots,K$.
-We perform four comparative static exercises.
-Consider first the effect of expanding the set $\mathcal{S}-\mathcal{M}$ of unfamiliar sources, due to the advent of robo-advisors or the popular advice that projects should be evaluated by “uninvolved managers” to counteract entrenched biases in managerial decision making (Guenzel and Malmendier, 2020).
+with decreasing magnitudes in $\lambda$ and $\eta_{k}^{-1}$, $k=1,\cdots,K$.
+We perform three comparative static exercises.
+Consider first the effect of expanding the set of unfamiliar sources, due to the advent of robo-advisors or the popular advice of using “uninvolved managers” to counteract entrenched managerial biases (Guenzel and Malmendier, 2020).
 This change has no impact on $\overline{\Delta}$ by DOOM while clearly reducing the errors associated with the best linear predictors (as captured by $\eta_{k}^{-1}$, $k=1,\cdots,K$).
-Consequently, it mitigates the distortionary effects of misspecified learning on the manager’s decision making and expected utility.
-Consider next the effect of increasing the signal-to-noise ratios of the sources, which causes $\eta_{k}^{-1}$, $k=1,\cdots,K$ to decrease.
-Although this change unambiguously mitigates the distortions of the manager’s decisions and expected utility, holding $\overline{\Delta}$ fixed, its impact on $\overline{\Delta}$ is generally ambiguous.
-This suggests that technological advances that lower data collection and processing costs (e.g., the AI revolution) could have differential impacts on our manager, depending on whether they operate by increasing the signal-to-noise ratios of familiar sources or those of novel sources.
-Consider next corrective measures aimed at bringing misspecifications closer to zero, which in reality range from debiasing board members who are prone to select and retain like-minded CEOs (Guenzel and Malmendier, 2020) to nudges such as premortem, i.e., imagine all twists and turns (Cohee and Barnhart, 2023).
-To the extent that the comparative statics of $\overline{\Delta}$ against these measures are generally ambiguous in the presence of compound misspecified learning, their impacts on the manager’s decision making and welfare can be rather nuanced.
-Lastly, consider the effect of expanding the portfolio of the projects overseen by the manager.
-Since this change mechanically increases $\sum_{k=1}^{K}\eta_{k}$, it exacerbates the distortions of the manager’s decisions and expected utility when the distributions of the signals that report on the additional projects are correctly specified.
-In the presence of misspecified learning, adding projects with independent returns to the portfolio creates the illusion that the optimal portfolio is well diversified.
-Consequently, more funds are allocated to projects the manager is overconfident about, thereby exacerbating the distortions of his decisions and welfare.
-6 Concluding remarks
-This paper takes the first step toward understanding the outcomes of Bayesian learning in the presence of multiple misspecified information sources.
-We propose a framework that encompasses a broad set of economic phenomena and develop a methodology that delivers tractable results and new insights.
-Several questions remain open.
-The first question concerns the utilization of current findings as pivotal steps in approaching new inquiries.
-For example, how do managerial misperceptions interact with minority employees’ decisions to undertake productivity-enhancing investment?
-How does the polarization of voters’ behavior influence politicians’ choices of platforms and policies?
-How are equilibrium asset prices affected by distortions in individual investors’ portfolio choices?
-Successfully addressing these questions requires that we incorporate misspecified active learning into the current framework and tackle additional theoretical challenges such as the convergence of equilibria.
-As discussed in Section 1.1, our results on labor market discrimination are distinctive from those due to tastes or rational beliefs, prompting the need for experimental validation.
-Some of the experimental designs proposed in Section 1.1, such as testing the comparative statics regarding informational interventions, have already been implemented by Bohren et al.
-(2023), albeit in a simpler decision environment devoid of misspecified learning about unfamiliar sources.
-Testing models of misspecified learning can be challenging, although successful endeavors have been undertaken previously; see, e.g., Bohren et al.
-(2019) for using the dynamics of beliefs to identify the source of discrimination.
-Considering the state of the literature, an empirical evaluation of our theory by means of experiments seems both interesting and feasible.
-Outside the realm of misspecified Bayesian learning, misperceptions are often explained by learning frictions such as non-Bayesian updating or errors in reasoning.
-Ortoleva (2022) provides an up-to-date survey of these models and their economic applications.
-In contrast, our DM has a misspecified prior but is otherwise a standard Bayesian.
-We view the two approaches as complementary.
-Their connection is explored by Bohren and Hauser (2023), who provide conditions when a non-Bayesian updating rule has a misspecified-learning representation.
-Appendix A Proofs
-A.1 Proof outline
-We begin by describing the problem studied by HKS and discuss the relations between our analyses.
-HKS examine a problem of misspecified Bayesian learning in which the DM observes in each period $t=1,2,\cdots$ a $D$-dimensional Gaussian signal with mean $\Gamma f\in\mathbb{R}^{D}$ and covariance matrix $\Sigma\in\mathbb{R}^{D\times D}$.
-Initially, the DM knows the transformation matrix $\Gamma\in\mathbb{R}^{D\times N}$ while misspecifying the first entry of the fundamental $f\in\mathbb{R}^{N}$.
-He may or may not know the true covariance matrix $\Sigma^{*}$, depending on whether he is in Case (I) or Case (III).
-[^14] Case (II) of HKS captures situations in which the DM holds a degenerate prior belief about the fundamental value.
-The support of his prior belief is
-$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma=\Sigma^{*}\}$
-in Case (I). In Case (III), it becomes
-$\{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}:f_{1}=\tilde{f}_{1},\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}}\},$
-where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$.
-Over time, the DM updates his belief about $(f,\Sigma)$ in a Bayesian manner based on the prior and signal realizations.
-For both cases, HKS demonstrate that the DM’s limiting posterior belief concentrates on the KL minimizer and provide characterizations for the latter.
-The current analysis differs from that of HKS in the following aspects.
-  1. D1
-    We allow the DM to misspecify multiple entries of the fundamental, while HKS focus on the case of a single misspecified entry.
-  2. D2
-    We are interested in the exercise of predicting the true state of the world based on signal realizations.
-To this end, we assume that signals equal biases plus the true state, plus state-independent errors with a covariance matrix $\Omega$.
-  3. D3
-    $\Omega$ is taken to be arbitrary when we solve the KL minimizer and prove that it is the limit point of the DM’s posterior belief.
-When applying our results to derive economic insights in more specific applications, we restrict $\Omega$ to be low-rank updates of diagonal matrices.
-How do these differences impact the analysis?
-For establishing the convergence of the limiting posterior belief to the KL minimizer, it turns out that D1 does not matter, whereas the differences due to D2 are quite nuanced.
-We highlight them in Appendix A.2 when we replicate the convergence proof of HKS line by line.
-Our solution to the KL minimization problem nests that of HKS due to D1.
-Our method exploits the first-order conditions associated with KL minimization, together with the block matrix inversion formula.
-Its difference from HKS’ guess-and-verify approach is underscored in Appendix A.3 when we present the proof of Lemma 2.
-The restriction of $\Omega$ to low-rank updates of diagonal matrices is suitable for the economic applications of our interest.
-The proofs presented in Appendix A.4 exploit existing mathematical results on low-rank updates of invertible matrices.
-We follow the notational conventions developed in the main text.
-Specifically, $x$ denotes the $N$-vector $[x_{1},\cdots,x_{N}]^{\top}$, $\text{diag}(\cdot)$ denotes the diagonal matrix created from a vector, $\mathbf{1}_{k}$ and $\mathbf{0}_{k}$ denote $k$-vectors of ones and zeros, respectively, for each $k\in\mathbb{N}$, and $\mathbb{I}_{k\times k}$ denotes the $k\times k$ identity matrix.
-$\otimes$ denotes the Kronecker product.
-$\max\{a,b\}$ is written as $a\vee b$; $\min\{a,b\}$ as $a\wedge b$.
-A.2 Proofs of the concentration of limiting posterior belief on KL minimizer
-This appendix gathers proofs for the concentration of the DM’s limiting belief on the set of KL minimizers. We begin by presenting a useful lemma.
-Lemma 5.
-Let $G\in\mathbb{R}^{N\times N}$ be any symmetric matrix and $\lambda_{k}(G)$ denote the $k^{th}$ largest eigenvalue of $G$, $k=1,\cdots,N$. The following are true for any $x\in\mathbb{R}^{N}$:
-$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+x^{\top}x)$
-[…]
-Proof.
-Weyl’s inequality (Weyl, 1912) states that $\lambda_{i+j-1}(A+B)\leq\lambda_{i}(A)+\lambda_{j}(B)\leq\lambda_{i+j-N}(A+B)$ for any Hermitian matrices $A$ and $B$ that belong to an inner product space of dimension $N$.
-Letting $A=G$ and $B=xx^{\top}$ in Weyl’s inequality yields
-$\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)+\lambda_{N}(xx^{\top})=\lambda_{1}(G)+0\leq\lambda_{1}(G+xx^{\top})$ (4)
-[…]
-where the equalities use the fact that $xx^{\top}$ has rank one and so has two eigenvalues: 00 and $\text{tr}(xx^{\top})$, together with $\text{tr}(xx^{\top})=\text{tr}(x^{\top}x)=x^{\top}x>0$.
-Interchanging the roles of $G+xx^{\top}$ and $G$ in the above derivation gives us the following:
-$\lambda_{N}(G)\leq\lambda_{1}(G+xx^{\top})+\lambda_{N}(-xx^{\top})=\lambda_{1}(G+xx^{\top})-x^{\top}x\leq\lambda_{1}(G)$ (6)
-[…]
-Together, (4), (5), and (7) imply that
-$\lambda_{N}(G)\leq\lambda_{N}(G+xx^{\top})\leq\lambda_{1}(G)\wedge(\lambda_{N}(G)+xx^{\top}),$
-that is the first inequality in the statement of the lemma. (4), (5), and (6) imply that
-$\lambda_{1}(G)\vee(\lambda_{N}(G)+xx^{\top})\leq\lambda_{1}(G+xx^{\top})\leq\lambda_{1}(G)+x^{\top}x,$
-that is the second inequality in the statement of the lemma. ∎
-Proof of Lemma 4.
-In the environment considered in 4.3, the support of the DM’s prior belief is
-$\Theta=\begin{Bmatrix}(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\text{ for some }\Omega\in\Lambda\end{Bmatrix},$
-where $\Lambda$ denotes the set of $N\times N$ SPD matrices whose minimum eigenvalues are bounded below by $\underline{\lambda}\in(0,1\wedge\lambda_{\min}(\Omega^{*}))$.
-The proof presented below verifies that this environment satisfies the assumptions of Berk (1966), hence the DM’s limiting posterior belief concentrates on the set of KL minimizers.
-Part (i) of Berk’s assumption says that the subjective density of the signals:
-$l(x\mid b,\Sigma)=\frac{1}{\sqrt{(2\pi)^{N}\text{det}\Sigma}}\exp\left(-\frac{1}{2}(x-b)^{\top}\Sigma^{-1}(x-b)\right),$
-is continuous in $(b,\Sigma)$ on $\Theta$. This is clearly true, since $\text{det}\Sigma$ and $\Sigma^{-1}$ are both continuous in $\Sigma$.
-Part (ii) of Berk’s assumption says that the subjective density equals zero on a set of measure zero under the true signal distribution.
-This is clearly true, since the subjective density is always positive.
-Part (iii) of Berk’s assumption says for every $(b,\Sigma)\in\Theta$, there exists an open neighborhood $U\subset\Theta$ of $(b,\Sigma)$ such that
-$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in U}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]<\infty.$
-Let $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ denote the minimum eigenvalue and the maximum eigenvalue of $\Sigma^{\prime}$, respectively. Then
-$|\log l(x\mid b^{\prime},\Sigma^{\prime})|$ $=|\frac{1}{2}\log((2\pi)^{N}\text{det}\Sigma^{\prime})+\frac{1}{2}(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})|$
-$\leq\frac{1}{2}\left(N|\log(2\pi\lambda_{\max}(\Sigma^{\prime}))|\vee|\log(2\pi\lambda_{\min}(\Sigma^{\prime}))|+\frac{1}{\lambda_{\min}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right),$
-where the inequality uses the facts that the determinant of a matrix equals the product of its eigenvalues, the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded above by $\lambda_{\max}(\Sigma^{\prime-1})=1/\lambda_{\min}(\Sigma^{\prime})$, and $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$.
-Since $\lambda_{\min}(\Sigma^{\prime})$ and $\lambda_{\max}(\Sigma^{\prime})$ are continuous in the entries of $\Sigma^{\prime}$, the last expression is finite on every finite neighborhood $U$.
-The proof so far is the same as that of HKS’ Theorem A.1 Case (III).
-The last step of the proof differs slightly due to our restriction of the covariance matrix to those of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
-Part (iv) of Berk’s assumption states that for every $c\in\mathbb{R}$, there exists $D\subset\Theta$ with a compact complement $\Theta-D$ such that
-$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in D}|\log l(x\mid b^{\prime},\Sigma^{\prime})|\right]\leq c.$
-Fix any $k_{1},k_{2}>0$, and let $D$ be the set of $(b^{\prime},\Sigma^{\prime})$ such that $\|b^{\prime}-b^{*}\|>k_{1}$ and the corresponding $\Omega^{\prime}$ satisfies $\lambda_{\max}(\Omega^{\prime})>k_{2}$.
-By Lemma 5, $\lambda_{\max}(\Sigma^{\prime})\geq\lambda_{\max}(\Omega^{\prime})(>k_{2})$ and $\lambda_{\min}(\Sigma^{\prime})\geq\lambda_{\min}(\Omega^{\prime})(\geq\underline{\lambda})$.
-Substituting these bounds into HKS’ proof of their Theorem A.1 Case (III) and replicating line by line, we obtain the following:
-$\log l(x\mid b^{\prime},\Sigma^{\prime})$
-$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
-$\leq$ $-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee\|x-b^{\prime}\|^{2})\right),$
-where the first inequality uses the facts that $\text{det}\Sigma^{\prime}\geq\lambda_{\min}^{N-1}(\Sigma^{\prime})\lambda_{\max}(\Sigma^{\prime})$, $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}>0$, and the Rayleigh quotient $(x-b^{\prime})^{\top}(\Sigma^{\prime})^{-1}(x-b^{\prime})/(x-b^{\prime})^{\top}(x-b^{\prime})$ is bounded below by $\lambda_{\min}(\Sigma^{\prime-1})=1/\lambda_{\max}(\Sigma^{\prime})$.
-The second inequality follows from minimizing $\log\lambda_{\max}(\Sigma^{\prime})+\|x-b^{\prime}\|^{2}/\lambda_{\max}(\Sigma^{\prime})$ over the range $\lambda_{\max}(\Sigma^{\prime})>k_{2}$.
-To obtain a lower bound for $\|x-b^{\prime}\|^{2}$, we note that $x-b^{\prime}=x-b^{*}+(b^{*}-b^{\prime})=(\omega+\varepsilon)+(b^{*}-b^{\prime})$ and so $\|x-b^{\prime}\|^{2}\geq\|\omega+\varepsilon\|^{2}+\|b^{*}-b^{\prime}\|^{2}-2\|\omega+\varepsilon\|\|b^{*}-b^{\prime}\|$ by the Cauchy-Schwarz inequality.
-Since $\|b^{*}-b^{\prime}\|>k_{1}$ by assumption, it follows that for any given value of $\|\omega+\varepsilon\|<k_{1}/2$, $\|b^{*}-b^{\prime}\|^{2}+\|\omega+\varepsilon\|^{2}-2\|b^{*}-b^{\prime}\|\|\omega+\varepsilon\|$ is minimized at $\|b^{*}-b^{\prime}\|=k_{1}$ and so is bounded below by $k_{1}^{2}/4$.
-Consequently, the last line of the above inequality is bounded above by
-$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right).$
-Meanwhile, $\log l(x\mid b^{\prime},\Sigma^{\prime})$ is bounded above by
-$-\frac{1}{2}\left(N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log\lambda_{\max}(\Sigma^{\prime})+\frac{1}{\lambda_{\max}(\Sigma^{\prime})}\|x-b^{\prime}\|^{2}\right)$
-$\leq-\frac{1}{2}N\log\lambda_{\min}(\Sigma^{\prime})\leq N|\log\underline{\lambda}|/2$
-when $k_{2}(<\lambda_{\max}(\Sigma^{\prime}))$ is large, where the second inequality uses the bound that $\lambda_{\min}(\Sigma^{\prime})\geq\underline{\lambda}$ and the assumption that $\underline{\lambda}<1$.
-Together, we obtain the following:
-$\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\sup_{(b^{\prime},\Sigma^{\prime})\in\Theta}\log l(x\mid b^{\prime},\Sigma^{\prime})\right]\leq\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\mathbf{1}_{\|\omega+\varepsilon\|\leq k_{1}/2}\sup_{(b^{\prime},\Sigma^{\prime})\in D}l(x\mid b^{\prime},\Sigma^{\prime})\right]+N|\log\underline{\lambda}|/2$
-$\leq-\frac{1}{2}\left[N\log(2\pi)+(N-1)\log\lambda_{\min}(\Sigma^{\prime})+\log(k_{2}\vee k_{1}^{2}/4)\right]\mathbb{P}(\|\omega+\varepsilon\|\leq k_{1}/2)+N|\log\underline{\lambda}|/2,$
-where the last line becomes arbitrarily negative as $k_{1},k_{2}\rightarrow+\infty$.
-It remains to verify that $\Theta-D$ is compact.
-By definition, this set consists of all vectors $b^{\prime}$ such that $\|b^{\prime}-b^{*}\|\leq k_{1}$, together with all covariance matrices of the form $\Omega^{\prime}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, whereby the eigenvalues of $\Omega^{\prime}$ belong to $[\underline{\lambda},k_{2}]$.
-The last set is compact, and this completes the proof.
-∎
-Proof of Lemma 1.
-In the environment considered in the baseline model, the support of the DM’s prior belief is
-$\Theta=\{(b,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{N\times N}:b_{i}=\tilde{b}_{i}\text{ }\forall i\in\mathcal{M}\text{ and }\Sigma=\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}\}$
-where $\Omega^{*}$ denotes the true covariance matrix of Gaussian errors.
-Since $\Omega^{*}\in\Lambda$, all we need is to replace every $\Sigma$ and $\Sigma^{\prime}$ in the proof of Lemma 4 with $\Sigma^{*}\coloneqq\Omega^{*}+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$.
-∎
-A.3 Proof of Lemma 2
-Since $\Sigma^{*}$ is a positive definite matrix, it is invertible. Applying the block matrix inversion formula to $\Sigma^{*}$ yields
-$\Sigma^{*-1}=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right],$
-where $A$, $B$, $C$, and $D$ were defined in Section 3.3. The system (3) of first-order conditions associated with KL minimization, in its matrix form, is
-$-(D-CA^{-1}B)^{-1}CA^{-1}\widetilde{\Delta}+(D-CA^{-1}B)^{-1}\widehat{\Delta}=\mathbf{0}_{N-M}.$
-Multiplying both sides of this equation by $D-CA^{-1}B$ yields $\widehat{\Delta}=CA^{-1}\widetilde{\Delta}$ as desired.
-Using this result to simplify $\Sigma^{*-1}\Delta$, we obtain the following:
-$\Sigma^{*-1}\Delta$ $=\left[\begin{array}[]{c|c}A^{-1}+A^{-1}B(D-CA^{-1}B)^{-1}CA^{-1}&-A^{-1}B(D-CA^{-1}B)^{-1}\\ \hline\cr-(D-CA^{-1}B)^{-1}CA^{-1}&(D-CA^{-1}B)^{-1}\end{array}\right]\left[\begin{array}[]{c}\widetilde{\Delta}\\ \hline\cr\widehat{\Delta}\end{array}\right]$ (12)
-$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}+A^{-1}B(D-CA^{-1}B)^{-1}(CA^{-1}\widetilde{\Delta}-\widehat{\Delta})\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right]$ (14)
-$=\left[\begin{array}[]{c}A^{-1}\widetilde{\Delta}\\ \hline\cr\mathbf{0}_{N-M}\end{array}\right].$ (17)
-Simplifying $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-\hat{b})-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=-\mathbf{1}_{N}^{\top}\Sigma^{*-1}\Delta$ accordingly yields $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\mathbf{1}_{M}^{\top}A^{-1}\widetilde{\Delta}$ as desired.
-∎
-A.4 Proofs of results that characterize KL minimizers
-Proof of Proposition 1.
-The environment described in the proposition features $\Sigma^{*}=\text{diag}(v)+\alpha\alpha^{\top}$, $A=\text{diag}(v_{1},\cdots,v_{M})+[\alpha_{1},\cdots,\alpha_{M}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$ and $C=[\alpha_{M+1},\cdots,\alpha_{N}]^{\top}[\alpha_{1},\cdots,\alpha_{M}]$.
-Since $A$ is a rank-one update of an invertible matrix,
-$A^{-1}=\text{diag}(\nu_{1},\cdots,\nu_{M})-\frac{1}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\begin{bmatrix}\alpha_{1}\nu_{1}\\ \vdots\\ \alpha_{M}\nu_{M}\end{bmatrix}\begin{bmatrix}\alpha_{1}\nu_{1}&\cdots&\alpha_{M}\nu_{M}\end{bmatrix}$
-by the Sherman-Morrison formula. Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
-$\widehat{\Delta}=CA^{-1}\widetilde{\Delta}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\begin{bmatrix}\alpha_{M+1}\\ \vdots\\ \alpha_{N}\end{bmatrix}\overline{\Delta},$
-where $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i})$.
-To calculate the distortion in the best-linear predictor, multiply (17) by $\alpha^{\top}$ to obtain
-$\alpha^{\top}\Sigma^{*-1}\Delta=[\alpha_{1},\cdots,\alpha_{M}]A^{-1}\widetilde{\Delta}=\left(1-\frac{\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}{1+\sum_{i=1}^{M}\alpha_{i}^{2}\nu_{i}}\right)\sum_{i=1}^{M}\alpha_{i}\nu_{i}\Delta_{i}=\overline{\Delta}.$
-Since $\text{BLP}^{*}(X)=\alpha^{\top}\Sigma^{*-1}(X-b^{*})$ and $\widehat{\text{BLP}}(X)=\alpha^{\top}\Sigma^{*-1}(X-\hat{b})$ in the current environment, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\alpha^{\top}\Sigma^{*-1}\Delta=-\overline{\Delta}$ as desired.
-∎
-Proof of Proposition 2.
-With a common latent factor, the KL minimization problem becomes
-$\min_{\Delta\in\mathbb{R}^{N+1}:\Delta_{i}=\tilde{b}_{i}-b_{i}^{*}\forall i\in\mathcal{M}}\frac{1}{2}\Delta^{\top}\Phi\Delta\text{ where }\Phi\coloneqq\Gamma^{\top}\Sigma^{*-1}\Gamma.$
-Since $\Sigma^{*}=\text{diag}(v)+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, $\Sigma^{*-1}=\text{diag}(\nu)-\nu\nu^{\top}/(1+\mathbf{1}_{N}^{\top}\nu)$ by the Sherman-Morrison formula. Thus
-$\Phi=\Gamma^{\top}\text{diag}(\nu)\Gamma-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\Gamma^{\top}\nu\nu^{\top}\Gamma,$
-where the first part of the right-hand side equals (recall that $\Gamma=[\mathbb{I}_{N\times N}\mid\mathbf{1}_{N}]$)
-$\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\text{diag}(\nu)\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right]=\left[\begin{array}[]{c|c}\text{diag}(\nu)&\nu\\ \hline\cr\nu^{\top}&\mathbf{1}_{N}^{\top}\nu\end{array}\right],$
-and the second part equals
-$\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\left[\begin{array}[]{c}\mathbb{I}_{N\times N}\\ \hline\cr\mathbf{1}_{N}^{\top}\end{array}\right]\nu\nu^{\top}\left[\begin{array}[]{c|c}\mathbb{I}_{N\times N}&\mathbf{1}_{N}\end{array}\right].$
-Subtracting the second part from the first yields
-$\Phi=\left[\begin{array}[]{c|c}\text{diag}(\nu)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\nu^{\top}&\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu\\ \hline\cr\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\nu^{\top}&1-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\end{array}\right].$
-The system of first-order conditions associated with KL minimization is
-$E\begin{bmatrix}\Delta_{M+1}\\ \vdots\\ \Delta_{N+1}\end{bmatrix}=-F\begin{bmatrix}\Delta_{1}\\ \vdots\\ \Delta_{M}\end{bmatrix},\text{ where }$
-$E\coloneqq\begin{bmatrix}\Phi_{M+1,M+1}&\cdots&\Phi_{M+1,N+1}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,M+1}&\cdots&\Phi_{N+1,N+1}\end{bmatrix}\text{ and }F\coloneqq\begin{bmatrix}\Phi_{M+1,1}&\cdots&\Phi_{M+1,M}\\ \vdots&\ddots&\vdots\\ \Phi_{N+1,1}&\cdots&\Phi_{N+1,M}\end{bmatrix}.$
-From the expression for $\Phi$, we can read off those for $E$ and $F$:
-$E=\text{diag}(\nu_{M+1}\cdots,\nu_{N},1)-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},&\cdots,&\nu_{N},&-1\end{bmatrix}$
-and
-$F=-\frac{1}{1+\mathbf{1}_{N}^{\top}\nu}\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{1},&\cdots,&\nu_{M}\end{bmatrix}.$
-Since $E$ is a rank-one update of an invertible matrix, applying the Sherman-Morrison formula shows that
-$E^{-1}=\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)$
-$+\frac{\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)/(1+\mathbf{1}_{N}^{\top}\nu)}{1-\begin{bmatrix}\nu_{M+1},\cdots,\nu_{N},-1\end{bmatrix}\text{diag}^{-1}(\nu_{M+1},\cdots,\nu_{N},1)\begin{bmatrix}\nu_{M+1}\\ \vdots\\ \nu_{N}\\ -1\end{bmatrix}/(1+\mathbf{1}_{N}^{\top}\nu)}$
-$=\text{diag}(v_{M+1},\cdots,v_{N},1)+\frac{1}{\sum_{i=1}^{M}\nu_{i}}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right]\left[\begin{array}[]{c|c}\mathbf{1}_{N-M}^{\top}&-1\end{array}\right].$
-Simplifying the system of first-order conditions accordingly yields
-$\left[\begin{array}[]{c}\Delta_{M+1}\\ \vdots\\ \Delta_{N}\\ \hline\cr\Delta_{N+1}\end{array}\right]=\overline{\Delta}\left[\begin{array}[]{c}\mathbf{1}_{N-M}\\ \hline\cr-1\end{array}\right],\text{ where }\overline{\Delta}\coloneqq\frac{\sum_{i=1}^{M}\nu_{i}\Delta_{i}}{\sum_{i=1}^{M}\nu_{i}}.\qed$
-Proof of Proposition 3.
-With multidimensional skills, the true covariance matrix of the signals becomes
-$\Sigma^{*}=\text{diag}\left(\Omega_{1},\cdots,\Omega_{N}\right)+\left(\mathbf{1}_{N}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{N}^{\top}\otimes\mathbb{I}_{K\times K}\right),$
-where $\otimes$ denotes the Kronecker product. Write $\Sigma^{*}$ as
-$\left[\begin{array}[]{c|c}A_{KM\times KM}&B_{KM\times K(N-M}\\ \hline\cr C_{K(N-M)\times KM}&D_{K(N-M)\times K(N-M)}\end{array}\right],$
-where
-$A=\underbrace{\text{diag}(\Omega_{1},\cdots,\Omega_{M})}_{G}+\underbrace{\left(\mathbf{1}_{M}\otimes\mathbb{I}_{K\times K}\right)}_{x}\Omega_{0}\underbrace{\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right)}_{y}$
-and
-$C=\left(\mathbf{1}_{N-M}\otimes\mathbb{I}_{K\times K}\right)\Omega_{0}\left(\mathbf{1}_{M}^{\top}\otimes\mathbb{I}_{K\times K}\right).$
-Since $A$ is a rank-$K$ update of an invertible matrix,
-$A^{-1}$ $=G^{-1}+G^{-1}x\left(\Omega_{0}^{-1}+yG^{-1}x\right)^{-1}yG^{-1}$
-$=\text{diag}(\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1})-\begin{bmatrix}\Omega_{1}^{-1}\\ \vdots\\ \Omega_{M}^{-1}\end{bmatrix}\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\begin{bmatrix}\Omega_{1}^{-1},\cdots,\Omega_{M}^{-1}\end{bmatrix}$
-where the first equality uses the Woodbury formula (Woodbury, 1950). Substituting the expressions for $A^{-1}$ and $C$ into Lemma 2 yields
-$\widehat{\Delta}$ $=CA^{-1}\widetilde{\Delta}=\mathbf{1}_{N-M}\otimes\left(\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}\right)$
-$=\mathbf{1}_{N-M}\otimes\overline{\Delta},$
-where the second line uses the fact that
-$\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=0}^{M}\Omega^{-1}_{i}=\Omega_{0}\Omega_{0}^{-1}=\mathbb{I}_{K\times K}$
-and so
-$\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}=\left(\sum_{i=0}^{M}\Omega_{i}\right)^{-1}.$
-The distortion in the best linear predictor can be obtained likewise, as follows:
-$\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)$ $=-\left(\mathbf{1}_{M}^{\top}\otimes\Omega_{0}\right)A^{-1}\widetilde{\Delta}$
-$=-\left(\Omega_{0}-\Omega_{0}\left(\sum_{i=1}^{M}\Omega_{i}^{-1}\right)\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\right)\sum_{i=1}^{M}\Omega_{i}^{-1}\Delta_{i}$
-$=-\left(\sum_{i=0}^{M}\Omega_{i}^{-1}\right)^{-1}\sum_{i=1}^{M}\Omega_{i}^{-i}\Delta_{i}=-\overline{\Delta}.\qed$
-Proof of Theorem 2.
-Part (i): The KL miminization problem considered in Case (III) of HKS (recall Appendix A.1) is
-$\min_{(f,\Sigma)\in\mathbb{R}^{N}\times\mathbb{R}^{D\times D}}D_{KL}(f^{*},\Sigma^{*}\mid f,\Sigma)\text{ s.t.
-}f_{1}=\tilde{f}_{1}\text{ and }\Sigma\text{ is SPD with }\lambda_{\min}(\Sigma)\geq\underline{\underline{\lambda}},$
-where $\underline{\underline{\lambda}}$ can be any positive number smaller than $1\wedge\lambda_{\min}(\Sigma^{*})$.
-The solution $(\hat{f},\widehat{\Sigma})$ to this problem is shown to satisfy two properties: First, $\Delta f_{i}\coloneqq\hat{f}_{i}-f_{i}^{*}$ is the same as when the DM knows the true covariance matrix $\Sigma^{*}$ from the outset.
-Second, $\widehat{\Sigma}=\Sigma^{*}+(\Gamma\Delta f)(\Gamma\Delta f)^{\top}$.
-The assumption that only a single entry of $f$ is misspecified plays no role in this result.
-This can be verified by replicating the proof of HKS line by line.
-Letting $f=b$ and $\Gamma=\mathbb{I}_{N\times N}$ in the above argument leads us to conclude that in the environment outlined in Section 4.3, the solution to the KL minimization problem would be $\Delta=\hat{b}_{i}-b_{i}^{*}$ and $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, had our DM been able to choose any SPD matrix whose minimum eigenvalue is bounded below by $\underline{\underline{\lambda}}$.
-The additional constraint we impose is that the covariance matrix must be of the form $\Omega+\mathbf{1}_{N}\mathbf{1}_{N}^{\top}$, where $\Omega$ is PSD with its minimum eigenvalue bounded below by $\underline{\lambda}<1\wedge\lambda_{\min}(\Omega^{*})$.
-We are left to verify that the solution to the unconstrained problem satisfies this additional constraint, and so must also solve the constrained optimization problem.
-To verify, note that the solution to the unconstrained problem: $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$, is attained by $\widehat{\Omega}=\Omega^{*}+\Delta\Delta^{\top}$.
-$\widehat{\Omega}$ is clearly PSD, and $\lambda_{\min}(\widehat{\Omega})\geq\lambda_{\min}(\Omega^{*})>\underline{\lambda}$ by Lemma 5.
-This completes the proof of Part (i) of the theorem.
-Part (ii) of the theorem follows from the decomposition exercise conducted in Section 4.3. Part (iii) follows from straightforward algebra. ∎
-Proof of Proposition 4.
-Consider first the setting examined in Section 3. There, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$, and so
+Consequently, it mitigates the distortionary effects of misspecified learning on the manager’s decision and objective welfare.
+Consider next the effect of increasing source precision, which causes $\eta_{k}^{-1}$, $k=1,\cdots,K$ to decrease.
+Although this change unambiguously mitigates the distortions in the DM’s decision and objective welfare, holding $\overline{\Delta}$ fixed, its impact on $\overline{\Delta}$ is generally ambiguous.
+This suggests that technological advances that lower data collection and processing costs could impact our manager differently, depending on whether they increase the precision of familiar or novel sources.
+Consider next corrective measures aimed at bringing misspecifications closer to zero, which in reality range from de-biasing board members who are prone to select and retain biased CEOs (Guenzel and Malmendier, 2020) to nudges such as premortem, i.e., imagine all twists and turns before decision making (Cohee and Barnhart, 2023).
+Given that the comparative statics of $\overline{\Delta}$ against these measures are generally ambiguous in the presence of compound misspecified learning, their impact on the manager’s decision and objective welfare can be rather nuanced.
+In addition to results that closely mirror those in the main text, our analysis delivers two additional insights.
+First, suppose we expand the portfolio overseen by the manager.
+Since this change mechanically increases $K$ and, hence, $\sum_{k=1}^{K}\eta_{k}$, it exacerbates the distortions in the manager’s decision and expected utility (assuming that the signals that report on the newly introduced projects have correctly specified distributions).
+The intuition is that, in the presence of misspecified learning, adding projects with independent returns to the portfolio creates the illusion that the optimal portfolio is well diversified.
+Consequently, more funds are allocated to projects the manager is overconfident about, resulting in a greater distortion in the welfare.
+Second, all distortionary effects are decreasing in the manager’s degree $\lambda$ of risk aversion.
+Appendix O.2 A Frequentist Perspective
+In Footnote 3 of the main text, we provided a frequentist foundation for the DM’s limiting belief when the covariance matrix of the signals is known from the outset.
+When the covariance matrix is also a subject of learning, we claim that the Bayesian asymptotic estimator of $b$ and $\Sigma$ under the normality assumption is equivalent to the joint frequentist estimator without normality assumption that satisfies simultaneously:
+  1. 1.
+    given $\Sigma$, $b$ is the best least squares linear estimator for means of $X$ (under parametric restriction to $\Theta$), assuming that the true covariance matrix is $\Sigma$;
+  2. 2.
+    given $b$, $\Sigma$ is the simple sample second moment estimator for covariances of $X$, assuming that the true means are $b$, i.e. $\mathbb{E}[(X-b)(X-b)^{\top}]$.
+It is easy to show the first claim: By Footnote 3 of the main text, the estimation of $b$ is obtained by minimizing the population level objective function, assuming that the true sample covariance matrix is $\Sigma$.
+To show the second claim, recall the KL divergence formula:
+$D_{KL}(b^{*},\Sigma^{*}\|b,\Sigma)=\frac{1}{2}\left(\text{tr}(\Sigma^{-1}\Sigma^{*})-N+(b-b^{*})^{\top}\Sigma^{-1}(b-b^{*})+\log\frac{\text{det}\Sigma}{\text{det}\Sigma^{*}}\right).$
+Taking derivative with $\Sigma$ gives the following FOC:
+$0=\frac{\partial D_{KL}}{\partial\Sigma}=$ $\frac{1}{2}\left(\frac{\partial}{\partial\Sigma}\text{tr}(\Sigma^{-1}\Sigma^{*})+\frac{\partial}{\partial\Sigma}\text{tr}\left(\Sigma^{-1}(b-b^{*})(b-b^{*})^{\top}\right)+\Sigma^{-1}\right)$
+$=$ $\frac{1}{2}\left(-\Sigma^{-1}\Sigma^{*}\Sigma^{-1}-\Sigma^{-1}(b-b^{*})(b-b^{*})^{\top}\Sigma^{-1}+\Sigma^{-1}\right).$
+This FOC implies the resulting estimator $\Sigma$ satisfies:
+$\Sigma$ $=\Sigma^{*}+(b-b^{*})(b-b^{*})^{\top}$
+$=\mathbb{E}[(X-b^{*})(X-b^{*})^{\top}]+(b-b^{*})(b-b^{*})^{\top}$
+$=\mathbb{E}[(X-b^{*}+b^{*}-b)(X-b^{*}+b^{*}-b)^{\top}]$
+$=\mathbb{E}[(X-b)(X-b)^{\top}],$
+where the penultimate line follows from the fact that $X-b^{*}$ is mean zero in expectation (taken under the true parameters $b^{*}$ and $\Sigma^{*}$).
+Appendix O.3 Proofs
+Proof of Proposition O.1.
+In the context of the baseline model, $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\overline{\Delta}$. Consequently,
 $\widehat{\text{EU}}-\text{EU}^{*}$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[-(\widehat{\text{BLP}}-\omega)^{2}+(\text{BLP}^{*}-\omega)^{2}]$
@@
 $=-\overline{\Delta}^{2}.$
-Consider next the setting examined in Section 4.3. There,
+In the context of Section 5.1 of the main text,
 $\widehat{\text{EU}}-\text{EU}^{*}$ $=-\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\widehat{\text{BLP}})^{2}]+\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})^{2}]$
 $=-\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*}-(\widehat{\text{BLP}}-\text{BLP}^{*}))^{2}]+\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})^{2}]$
-$=-\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\widehat{\text{BLP}}-\text{BLP}^{*})^{2}]}_{G}+2\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]}_{H}.$
-The remainder of the proof shows that $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ and $H=0$.
-Before proceeding, recall that $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$, and note that
-$\Sigma^{*-1}\Delta=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}]^{\top}$ (18)
-obtains from elaborating on (17).
+$=-\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\widehat{\text{BLP}}-\text{BLP}^{*})^{2}]}_{E}+2\underbrace{\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]}_{F}.$
+The proof presented below shows that $E=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$ and $F=0$.
+Before proceeding, recall that $\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$, and that
+$\Sigma^{*-1}\Delta=\begin{bmatrix}\nu_{1}(\Delta_{1}-\overline{\Delta})\\ \vdots\\ \nu_{M}(\Delta_{M}-\overline{\Delta})\\ \mathbf{0}_{N-M}\end{bmatrix}$
+by Equation (3) of the main text.
 Step 1.
-The decomposition exercise in Section 4.3 showed that
+In Section 5.1 of the main text, we demonstrated that
 $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}+\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*}).$
-Since $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ by Theorem 2, $\widehat{\Sigma}^{-1}-\Sigma^{*-1}=-\Sigma^{*-1}\Delta\Delta^{\top}\Sigma^{*-1}/(1+\Delta^{\top}\Sigma^{*-1}\Delta)$ by the Sherman-Morrison formula.
-Thus
+Since $\widehat{\Sigma}=\Sigma^{*}+\Delta\Delta^{\top}$ by Theorem 2, $\widehat{\Sigma}^{-1}=\Sigma^{*-1}-\Sigma^{*-1}\Delta\Delta^{\top}\Sigma^{*-1}/(1+\Delta^{\top}\Sigma^{*-1}\Delta)$ by the Sherman-Morrison formula.
+Consequently,
 $\mathbf{1}_{N}^{\top}(\widehat{\Sigma}^{-1}-\Sigma^{*-1})(X-b^{*})=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\Delta^{\top}\Sigma^{*-1}(X-b^{*}),$
@@
 $\widehat{\text{BLP}}(X)-\text{BLP}^{*}(X)=-\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}(1+\Delta^{\top}\Sigma^{*-1}(X-b^{*}))$
-and, hence,
-$G=\left(\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\right)^{2}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[(\Delta^{\top}\Sigma^{*-1}(X-b^{*}))^{2}+1+2\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right].$
+and, in turn,
+$E=\left(\frac{\overline{\Delta}}{1+\Delta^{\top}\Sigma^{*-1}\Delta}\right)^{2}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[(\Delta^{\top}\Sigma^{*-1}(X-b^{*}))^{2}+1+2\Delta^{\top}\Sigma^{*-1}(X-b^{*})\right].$
 The third term in the expectation operator has zero mean under the true signal distribution. The first term equals
@@
 […]
-$=\left(\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\right)^{2}\mathbb{E}[\omega^{2}]+\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})^{2}\mathbb{E}[\varepsilon_{i}^{2}]$
+$=\left(\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\right)^{2}\mathbb{E}[\omega^{2}]+\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})^{2}\mathbb{E}[\epsilon_{i}^{2}]$
 […]
 $=\Delta^{\top}\Sigma^{*-1}\Delta,$ ($\because\overline{\Delta}\coloneqq\sum_{i=1}^{M}\nu_{i}\Delta_{i}/(1+\sum_{i=1}^{M}\nu_{i})$)
-and inserting this back into the expression for $G$ yields $G=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$.
+which, when inserted back into the expression for $E$, yields $E=\overline{\Delta}^{2}(1+\Delta^{\top}\Sigma^{*-1}\Delta)^{-1}$.
 Step 2.
 Results so far imply that
-$\omega-\text{BLP}^{*}(X)$ $=\omega-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=\omega-\mathbf{1}_{N}^{\top}\left(\text{diag}(\nu)-\frac{\nu\nu^{\top}}{1+\mathbf{1}_{N}^{\top}\nu}\right)(\omega\mathbf{1}_{N}+\varepsilon)$
-$=\frac{\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}}{1+\mathbf{1}_{N}^{\top}\nu},$
+$\omega-\text{BLP}^{*}(X)$ $=\omega-\mathbf{1}_{N}^{\top}\Sigma^{*-1}(X-b^{*})=\omega-\mathbf{1}_{N}^{\top}\left(\text{diag}(\nu)-\frac{\nu\nu^{\top}}{1+\mathbf{1}_{N}^{\top}\nu}\right)(\omega\mathbf{1}_{N}+\epsilon)$
+$=\frac{\omega-\sum_{i=1}^{N}\nu_{i}\epsilon_{i}}{1+\mathbf{1}_{N}^{\top}\nu},$
 and that
-$\Delta^{\top}\Sigma^{*-1}(X-b^{*})$ $=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}](\omega\mathbf{1}_{N}+\varepsilon)$
-$=\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}.$
-Simplifying $H$ accordingly yields
-$H$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]$
+$\Delta^{\top}\Sigma^{*-1}(X-b^{*})$ $=[\nu_{1}(\Delta_{1}-\overline{\Delta}),\cdots,\nu_{M}(\Delta_{M}-\overline{\Delta}),\mathbf{0}_{N-M}^{\top}](\omega\mathbf{1}_{N}+\epsilon)$
+$=\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\epsilon_{i}.$
+Simplifying $F$ accordingly yields
+$F$ $=\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(\widehat{\text{BLP}}-\text{BLP}^{*})]$
 $=\frac{-\overline{\Delta}}{(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}[(\omega-\text{BLP}^{*})(1+\Delta^{\top}\Sigma^{*-1}(X-b^{*}))]$
-$=\frac{-\overline{\Delta}}{\left(1+\mathbf{1}_{N}^{\top}\nu\right)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\left(\omega-\sum_{i=1}^{N}\nu_{i}\varepsilon_{i}\right)\left(1+\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\varepsilon_{i}\right)\right]$
-$=\frac{-\overline{\Delta}}{(1+\mathbf{1}_{N}^{\top}\nu)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\left(\overline{\Delta}\mathbb{E}[\omega^{2}]-\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})\mathbb{E}[\varepsilon_{i}^{2}]\right)$
+$=\frac{-\overline{\Delta}}{\left(1+\mathbf{1}_{N}^{\top}\nu\right)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\mathbb{E}_{(b^{*},\Sigma^{*})}\left[\left(\omega-\sum_{i=1}^{N}\nu_{i}\epsilon_{i}\right)\left(1+\overline{\Delta}\omega+\sum_{i=1}^{M}\nu_{i}(\Delta_{i}-\overline{\Delta})\epsilon_{i}\right)\right]$
+$=\frac{-\overline{\Delta}}{(1+\mathbf{1}_{N}^{\top}\nu)(1+\Delta^{\top}\Sigma^{*-1}\Delta)}\left(\overline{\Delta}\mathbb{E}[\omega^{2}]-\sum_{i=1}^{M}\nu_{i}^{2}(\Delta_{i}-\overline{\Delta})\mathbb{E}[\epsilon_{i}^{2}]\right)$
 […]
@@
 as desired. ∎
-Proof of Proposition 5.
-The solution to the portfolio choice problem presented in Section 5.2 is
+Proof of Proposition O.2.
+The solution to the portfolio choice problem is given by
 $\alpha^{*}_{j}=\frac{\eta_{j}}{\lambda}\left(\mu^{*}_{j}-\frac{\sum_{k=1}^{K}\mu^{*}_{k}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\right)\text{ }\forall j=1,\cdots,K$
-under the correctly specified model, and
+under the correctly specified model, and it is given by
 $\widehat{\alpha}_{j}=\frac{\eta_{j}}{\lambda}\left(\widehat{\mu}_{j}-\frac{\sum_{k=1}^{K}\widehat{\mu}_{k}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\right)\text{ }\forall j=1,\cdots,K$
@@
 The distortion in the DM’s objective expected utility thus equals
-$\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})\mu_{k}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})^{2}/\eta_{k}-\sum_{k=1}^{K}\alpha^{*}_{k}\mu_{\alpha_{k}}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}\left(\alpha_{k}^{*}\right)^{2}/\eta_{k}$
+$\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})\mu_{k}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}(\alpha^{*}_{k}+\Delta_{\alpha_{k}})^{2}/\eta_{k}-\left[\sum_{k=1}^{K}\alpha^{*}_{k}\mu_{\alpha_{k}}^{*}-\frac{\lambda}{2}\sum_{k=1}^{K}\left(\alpha_{k}^{*}\right)^{2}/\eta_{k}\right]$
 $=\sum_{k=1}^{K}\Delta_{\alpha_{k}}(\mu_{k}^{*}-\lambda\alpha_{k}^{*}/\eta_{k})-\frac{\lambda}{2}\sum_{k=1}^{K}(\Delta_{\alpha_{k}})^{2}/\eta_{k}$
@@
 $\mu_{j}^{*}-\lambda\alpha_{j}^{*}/\eta_{j}=\frac{\sum_{k=1}^{K}\mu_{k}^{*}\eta_{k}-\lambda W}{\sum_{k=1}^{K}\eta_{k}}\text{ }\forall j=1,\cdots,K$
-and $\sum_{k=1}^{K}\Delta_{\alpha_{k}}=0$. The third equality follows from straightforward algebra.∎
+and $\sum_{k=1}^{K}\Delta_{\alpha_{k}}=0$ (that is, wealth is fixed), and the third equality follows from straightforward algebra.∎